### v2

# Cohomogeneity one $\mathsf{RCD}$-spaces

Abstract.
We study $\mathsf{RCD}$-spaces $(X,d,\mathfrak{m})$ with group actions by isometries preserving the reference measure $\mathfrak{m}$ and whose orbit space has dimension one, i.e. cohomogeneity one actions. To this end we prove a Slice Theorem asserting that each slice at a point is homeomorphic to a non-negatively curved $\mathsf{RCD}$-space. Under the assumption that $X$ is non-collapsed we further show that the slices are homeomorphic to metric cones over homogeneous spaces with $\operatorname{Ric}\geq 0$. As a consequence we obtain complete topological structural results and a principal orbit representation theorem. Conversely, we show how to construct new $\mathsf{RCD}$-spaces from a cohomogeneity one group diagram, giving a complete description of $\mathsf{RCD}$-spaces of cohomogeneity one. As an application of these results we obtain the classification of cohomogeneity one, non-collapsed $\mathsf{RCD}$-spaces of essential dimension at most $4$.
Key words and phrases:
$\mathsf{RCD}$-space, group action, cohomogeneity one, non-collapsed
2020 Mathematics Subject Classification:
53C20; 53C21; 53C23
Supported by a UKRI Future Leaders Fellowship [grant number MR/W01176X/1; PI: J Harvey], by DFG (281869850, RTG 2229 “Asymptotic Invariants and Limits of Groups and Spaces” and by DFG-Eigenstelle Fellowship CO 2359/1-1
Supported by PAPIIT-UNAM project IN101322 and CONAHCyT-SNII 323477
Supported in part by a Margarita Salas Fellowship of the Universidad Autónoma de Madrid CA1/RSUE/2021–00625, and by research grants MTM2017-‐85934-‐C3-‐2-‐P, PID2021–124195NB–C32 from the Ministerio de Economía y Competitividad de España (MINECO)
1. Introduction
The theory of $\mathsf{RCD}$-spaces has emerged from, and is still largely motivated, by the extensive and deep work of Cheeger and Colding on Gromov-Hausdorff limit spaces of sequences consisting of Riemannian manifolds with Ricci curvature bounded from below [18, 19, 20]. Cheeger and Colding themselves asked whether a synthetic treatment of Ricci curvature bounds was possible, the most natural answer being the now well developed theory of spaces with the curvature-dimension condition, initiated by Lott, Sturm and Villani [56, 74, 75] and refined to exclude Finslerian spaces by Ambrosio, Gigli and Savaré [2, 32]. Thus, the study of $\mathsf{RCD}$-spaces carries intrinsic theoretical value; however it has also produced several non-trivial applications to the study of Riemannian manifolds which are beyond the scope of the classical standpoint, i.e. only using smooth tools (see [33, Section 7] for some examples of these results).
Therefore it is compelling to understand the geometries and topologies of non-smooth spaces with curvature bounded below. As these spaces admit topological and metric singularities, which may have a complicated arrangement, it is necessary to look for systematic ways to study them. In particular, the topological structure of $\mathsf{RCD}$-spaces is complicated and, as of now, there are only a few results describing it (see for example [59, 70, 78] and references therein).
An approach to produce such a systematic method for the study of $\mathsf{RCD}$-spaces is the so called extended symmetry program. Grove originally proposed in [38] that in the search of new examples or obstructions to the existence of Riemannian metrics with positive sectional curvature we should first study the manifolds with a high degree of symmetry. This approach has been extended to the study of Riemannian manifolds with other lower curvature bounds, and has also been successfully extrapolated to the study of Alexandrov spaces, i.e. length spaces with a synthetic lower curvature bound (see [28, 44] and references therein).
Since the group of measure preserving isometries of an $\mathsf{RCD}$-space is a Lie group [40, 72], we can try to follow the symmetry program philosophy for $\mathsf{RCD}$-spaces. In this article we do this by considering cohomogeneity one $\mathsf{RCD}$-spaces, that is $\mathsf{RCD}$-spaces admitting an isometric and measure-preserving action of a compact Lie group in such a way that the orbit space is of Hausdorff dimension one. Our main results are the following.
Theorem A.
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space. Let $G$ be a compact Lie group acting on $X$ by measure preserving isometries and cohomogeneity one. Denote by $X^{\ast}$ the quotient space $\pi\colon X\to X/G$ equipped with the quotient metric $d^{\ast}(\pi(x),\pi(y)):=d(G(x),G(y))$. Then the following hold:
  1. (a)
    When $X^{\ast}$ is homeomorphic to $[0,1]$, then $X$ is equivariantly homeomorphic to the union of two cone bundles.
  2. (b)
    When $X^{\ast}$ is homeomorphic to $[0,\infty)$, then $X$ is equivariantly homeomorphic to a cone bundle.
  3. (c)
    When $X^{\ast}$ is homeomorphic to $\mathbf{S}^{1}$, then $X$ is the total space of a fiber bundle with homogeneous fiber $G/H$ and structure group $N_{G}(H)/H$.
  4. (d)
    When $X^{\ast}$ is homeomorphic to $\mathbb{R}$, then $X$ is equivariantly homeomorphic to ×$\mathbb{R}\times G/H$.
Moreover the cone fibers in items $(a)$ and $(b)$ are cones over homogeneous spaces.
If the tangent cones of $X$ at some points in the singular orbits $G(x_{+})$, $G(x_{-})$ have a metric cone structure, then the cone fibers are cones over $\mathsf{RCD}(N-k_{\pm}-2,N-k_{\pm}-1)$-spaces, where $k_{\pm}=\dim(G(x_{\pm}))$.
This result gives the full description of $\mathsf{RCD}$-spaces of cohomogeneity $1$, completing the rigidity results of Galaz-García, Kell, Mondino and Sosa in [27, Corollary 1.4]. Also, from the proof it is clear that in the case when we have two spaces $X_{1}$, $X_{2}$ with the same orbit space, and same orbit types, then they are equivariantly homeomorphic. We also show that to obtain the conclusions about the homeomorphism type in Theorem A we only need the geodesic space structure (see Section 3 and Theorem 4.12).
The topological description depends on having a geodesic structure, and the geometry of the quotient space, as stated in Theorem 4.12. Moreover, similar conclusions hold when $G$ is non-compact but acts properly by isometries, see Remark 4.13.
Moreover, in the case of Theorem A (a) we get three subgroups $H$, $G_{-}$, $G_{+}$, that up to conjugation are all the possible isotropy groups of the action. By Theorem 3.14 it follows that $H<G_{\pm}$, and thus we have a tuple $(G,H,G_{-},G_{+})$ associated to the action. We show that the converse is true: for each cohomogeneity one group diagram $(G,H,G_{-},G_{+})$, that is a collection of subgroups where $H$ is a Lie subgroup of $K_{\pm}$, and in turn $K_{\pm}$ are Lie subgroups of $G$ and such that $K_{\pm}/H$ are homogeneous spaces with positive Ricci curvature, determines a cohomogeneity one $\mathsf{RCD}$-space.
Theorem B.
Let $G$ be a compact Lie group and $K_{+}$, $K_{-}$, $H$ Lie subgroups of $G$ such that $(G,H,K_{+},K_{-})$ is a cohomogeneity one group diagram. Then, there exists an $\mathsf{RCD}$-space $(X,d,\mathfrak{m})$ admitting a cohomogeneity one action of $G$ by measure preserving isometries such that the associated group diagram is $(G,H,K_{+},K_{-})$.
These two results give a complete description of $\mathsf{RCD}$-spaces of cohomogenity $1$, which is analogous to the descriptions of topological and Riemannian manifolds of cohomogeneity $1$, and Alexandrov spaces of cohomogeneity $1$ (see [61], [60], [29], and [28]).
A crucial ingredient in the proof of the classification Theorem A is the Slice Theorem. This is used in general in the Grove symmetry program as a recognition tool for the local topology around a fixed orbit.
All known proofs of the Slice Theorem, be it in the smooth or Alexandrov setting, require a notion of “space of directions” and exponential map (see [44]). In the $\mathsf{RCD}$-setting, while a notion of space of directions can be defined a.e. (see [42]), the infinitesimal-to-local relation (in the spirit of Perelman’s conical neighborhood theorem) is a priori not known to hold. That is, it is not yet fully understood how the mGH-tangent space and $L^{2}$-tangent module (which are equivalent by work of Gigli and Pasqualetto [37]) determine the space locally (see however the recent work of Honda and Peng [49] in this direction). We instead work in the setting of optimal transport theory, namely using ideas from needle-decomposition [15], and also exploit the low-dimensionality of the orbit space to circumvent this obstacle.
Theorem C.
Let $(X,d)$ be a geodesic space and $G$ a compact Lie group acting on $X$ by isometries such that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then for any $x_{0}\in X$, there exists $\delta_{0}>0$ such that for any $0<\delta<\delta_{0}$ we have $B_{\delta}(x_{0})\subset\mathcal{T}^{nb}_{u}$ and
$S_{x_{0}}:=\left(\Gamma^{-}_{u}(x_{0})\cap B_{\delta}(x_{0})\right)\cup\{x_{0}\}$
is a slice through $x_{0}$.
Here we used the customary definitions as follows. Denoting $u(x):=d(x,G(x_{0}))$, we set $\Gamma^{-}_{u}(x)=\{y\in X\mid(x,y)\in\Gamma^{-}_{u}\}$, where as usual, $\Gamma^{-}_{u}=\left\{(x,y)\mid u(y)-u(x)=d(x,y)\right\}$. We also denote the so-called non-branching transport set by $\mathcal{T}^{nb}_{u}$. It is worth pointing out that we do not assume the $\mathsf{RCD}$-condition (or any other bounded curvature condition) in this result and thus in particular it does not depend on any specific bound of the curvature. It depends only on the dimension of the orbit space and properties of the metric structure.
Once this description of the slices at each point has been established we move to show that in general the slices admit a non-negatively curved $\mathsf{RCD}$-structure (that is, each slice is homeomorphic to an $\mathsf{RCD}(0,M)$-space, for the appropriate dimension bound $M$ as stated below). When specializing to the case where all tangent cones of $X$ are metric cones we further show that slices are homeomorphic to metric cones over homogeneous spaces with adequate $\mathsf{RCD}$-structures. This condition is satisfied for example when $X$ is non-collapsed [23, Proposition 2.8].
Theorem D.
Let $(X,d,\mathfrak{m}_{X})$ be an $\mathsf{RCD}(K,N)$-space with $N>2$, such that all tangent cones are metric cones. Assume that $G$ is a compact Lie group acting effectively by measure preserving isometries on $X$, such that $(X^{\ast},d^{\ast},\mathfrak{m}_{X}^{\ast})$ has essential dimension equal to $1$. Then the following hold:
  1. (i)
    For any $x_{0}\in X$ the slice $S_{x_{0}}$ admits an $\mathsf{RCD}(0,N-k)$ structure, where $k=\dim(G(x_{0}))$.
  2. (ii)
    Moreover, $S_{x_{0}}$ is homeomorphic to a metric cone over a homogeneous smooth Riemannian manifold with Ricci curvature greater or equal to $N-k-2$.
We point out that by the work of Palais [65] the existence of slices is well understood for compact group actions by isometries. Nonetheless, in the Riemannian and Alexandrov setting the metric structure determines a canonical choice of a slice, which inherits geometric properties of the global space. In this way, for the local topological recognition we may apply a reduction on the dimension being considered while still considering spaces in the appropriate setting. This key point is beyond the conclusions of Palais’ Slice Theorem [65], and it is the point being proven in Theorem C and Theorem D. Thus, by Theorem D, Theorem C gives a positive answer to [22, Problem 1] for non-collapsed $\mathsf{RCD}$-spaces of cohomogeneity one.
The slice in Theorem D is the fiber of the cone bundles in Theorem A. In the non-collapsed case, the homogeneous spaces are determined by the slice.
Finally, we use our previous results to obtain a topological classification of non-collapsed $\mathsf{RCD}$-spaces of essential dimensions at most $4$, in the spirit of the corresponding classification in the smooth and Alexandrov cases. In fact, we obtain that all the $\mathsf{RCD}$-spaces considered with these restrictions are homeomorphic to Alexandrov spaces. This solves a particular case (that of spaces admitting cohomogeneity one actions) of a conjecture by Mondino asserting that $\mathsf{RCD}$-spaces with essential dimension $3$ are homeomorphic to orbifolds (see Question 9 in Section 2 of [22]).
Theorem E.
Let $G$ be a compact Lie group acting almost effectively by measure preserving isometries and cohomogeneity one on a closed non-collapsed $\mathsf{RCD}(K,N)$-space $(X,d,\mathfrak{m})$ with $N\leq 4$. Then $X$ is homeomorphic to an Alexandrov space.
Combining the characterization in [6] of homogeneous spaces admitting positive Ricci curvature with the cone construction in [51]. it is easy to construct $\mathsf{RCD}(K,N)$-spaces of cohomogeneity one for some $K\geq 0$. This construction allow us to show that that the dimension bound in Theorem E is optimal (see Theorem 7.1). Namely the following example shows that in dimension $5$ there are non-collapsed $\mathsf{RCD}$-spaces of cohomogeneity one which are not Alexandrov spaces of cohomogeneity one:
Theorem F.
The suspension $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ of $\mathbf{S}^{2}\times\mathbf{S}^{2}$ admits an non-collapsed $\mathsf{RCD}(K,5)$-structure for any $K\geq 0$ such that the suspension of the $\mathrm{SO}(3)\times\mathrm{SO}(3)$-action on $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is by measure preserving isometries and by cohomogeneity one, with group diagram $(\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(2)\times\mathrm{SO}(2),\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(3)\times\mathrm{SO}(3))$. Moreover, $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ with this action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ is not an Alexandrov space of cohomogeneity one, i.e. it does not admit a metric making it an Alexandrov space and such that the action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ is by isometries.
We observe that in the previous example we obtain $\mathsf{RCD}$-spaces of cohomogeneity one that are not non-collapsed, and the same holds true for spherical suspensions of homogeneous spaces with sufficient curvature, see Theorem 7.1. Nonetheless, we observe that since these examples are $\mathsf{RCD}$-spaces of cohomogeneity one, the stronger conclusions of Theorem D hold by construction. Thus we ask the following question:
Question 1.1.
Does there exist an $\mathsf{RCD}$-space $X$ equipped with a measure preserving isometric action of cohomogeneity one by a compact Lie group $G$, which is not non-collapsed, and such that the slice at any orbit does not admit an metric cone structure over a homogeneous space?
In other words, given an $\mathsf{RCD}$-space $X$ equipped with a measure preserving isometric action of cohomogeneity one by a compact Lie group $G$, does it always admit an non-collapsed invariant $\mathsf{RCD}$-structure?
We also observe that the spaces in Theorem A (c), and (d) admit an $\mathsf{RCD}(0,N)$-structure of cohomogeneity one, given by a smooth Riemannian metric on the smooth representatives of the homoeomrphism types. Moreover, due to Gigli’s Splitting Theorem [31] the spaces in Theorem A (d) do not admit an $\mathsf{RCD}(K,N)$-structure with $K>0$. For spaces in Theorem A (c), since they have infinite fundamental group, by Myers Theorem (see [59, Theorem 3.5] toghether with [78, Main Theorem]) it follows that they also do not admit an $\mathsf{RCD}(K,N)$-structure with $K>0$. For spaces in Theorem A (b), we give in Theorem I sufficient conditions to guarantee that they also admit an $\mathsf{RCD}(0,N)$-structure of cohomogeneity one. Thus it natural to ask if this also holds for the remaining cases:
Question 1.2.
Let $X$ be an $\mathsf{RCD}$-space as in Theorem A (a) or (b). Does it admit a (possibly different) new $\mathsf{RCD}(0,N)$-structure of cohomogeneity one?
This question was answered in the positive by Grove and Ziller in [39, Theorem B] when $X$ is smooth Riemannian manifold. Moreover, in [39, Theorem A], they showed that a smooth compact Riemannian manifold with an action of cohomogeneity one admits a Riemannian metric of positive Ricci curvature if and only if it has finite fundamental group. Due to the fact that Myers Theorem holds for $\mathsf{RCD}$-spaces, we ask if the cardinality of the fundamental group is the unique obstruction to “synthetic positive Ricci curvature” in the case of compact $\mathsf{RCD}$-spaces of cohomogeneity one.
Question 1.3.
Let $X$ be an $\mathsf{RCD}$-space as in Theorem A (a). If $X$ has finite fundamental group, does it admit an $\mathsf{RCD}(K,N)$-structure with $K>0$?
We present a simple presentation of the fundamental group of an $\mathsf{RCD}$-space for the form Theorem A (a) in Theorem 4.14.
The organization of the article is the following. In Section 2 we collect the definitions and results from the theory of non-smooth differential geometry and transformation groups that we will need in the sequel. In Section 3 we prove Theorem C. We also prove two results showing that our construction using the transport set is a slice without the cohomogeneity one restriction, albeit in the case that the orbit space happens to be isometric to an Alexandrov space, which are of independent interest. In Section 4 we prove Theorem D obtaining along the way a version of the orbit representation theorem. We also establish the topological rigidity result contained in Theorem A. We continue in Section 5 proving that one can construct $\mathsf{RCD}$-spaces from any given cohomogeneity one group diagram, as expressed in Theorem B. Finally, we address the low-dimensional classification of Theorem E in Section 6.
Acknowledgements The authors wish to thank Fernando Galaz-García, Nicola Gigli, John Harvey, Alexander Lytchak, Marco Radeschi and Masoumeh Zarei for very useful discussions and communications and for commenting on a first draft of this manuscript. We thank Elia Bruè for useful discussions about the infinitesimal action which improved section 4, and for pointing us to the work of Pan and Wei [66].
2. Preliminaries
2.1. Calculus on metric measure spaces
Here we introduce the differential structure that we need in order to define the appropriate curvature notions that we use later. For a more detailed exposition, the reader may consult the book by Gigli and Pasqualetto [36].
By a metric measure space we mean a triple $(X,d,\mathfrak{m})$, where $(X,d)$ is a complete and separable metric space, and the reference measure $\mathfrak{m}$ is a non-negative Borel measure on $X$ which is finite on balls.
Let $f\colon X\rightarrow\mathbb{R}$ be a Lipschitz function. We define its local Lipschitz constant $\mathrm{Lip}f\colon X\rightarrow\mathbb{R}$ as:
$\mathrm{Lip}f(x):=\limsup_{y\rightarrow x}\frac{|f(x)-f(y)|}{d(x,y)}.$
We denote the space of Lipschitz functions by $\mathrm{LIP}(X).$
Definition 2.1 (Cheeger energy).
Let $(X,d,\mathfrak{m})$ be a metric measure space. Given a function $f\in L^{2}(\mathfrak{m})$, we define its Cheeger energy as:
[…]
Definition 2.2 (Sobolev space).
Let $(X,d,\mathfrak{m})$ be a metric measure space. We define the Sobolev space $W^{1,2}(X):=\{f\in L^{2}(\mathfrak{m})\,|\,\mathrm{Ch}(f)<\infty\}$, equipped with the norm $\|f\|^{2}_{W^{1,2}}:=\|f\|^{2}_{2}+\mathrm{Ch}^{2}(f)$. In the case that this norm comes from an inner product, we say that $(X,d,\mathfrak{m})$ is infinitesimally Hilbertian.
[…]
Remark 2.3.
It should be noted that we can define Sobolev spaces, equivalently, via test plans as in Chapter $2$ of [36]. However, for our purpose we find more convenient to make use of approximations by Lipschitz functions.
Definition 2.4 (Pointwise inner product).
Given $f,g\in W^{1,2}(X)$ we define
$\langle\nabla f,\nabla g\rangle:=\frac{1}{2}\left(|\nabla(f+g)|^{2}-|\nabla(f-g)|^{2}\right)\in L^{1}(\mathfrak{m}).$
Let us recall the following characterization of infinitesimal Hilbertianity (see Theorem $4.3.3$ in [36]).
Proposition 2.5.
The space $W^{1,2}(X)$ with the pointwise inner product is a Hilbert space if and only if the parallelogram rule
(1) $2\left(|\nabla f|^{2}+|\nabla g|^{2}\right)=|\nabla(f+g)|^{2}+|\nabla(f-g)|^{2}$
holds $\mathfrak{m}$-a.e. for all $f,g\in W^{1,2}(X)$.
Definition 2.6 (Laplacian).
Let $(X,d,\mathfrak{m})$ be an infinitesimally Hilbertian space. A function $f\in W^{1,2}(X)$ is defined to be in the domain of the Laplacian, $D(\Delta)$, if there exists some $g\in L^{2}(\mathfrak{m})$ such that:
[…]
We denote the function $g$ (which if exists, is unique) by $\Delta f$ and refer to it as the Laplacian of $f$.
In case that the underlying space is infinitesimally Hilbertian we have at our disposal the following good properties of the Laplacian and of the pointwise inner product (see Theorem $4.3.3$ part $v),$ Remark $5.2.2,$ and Proposition $5.2.3$ in [36]).
Proposition 2.7.
Let $(X,d,\mathfrak{m})$ be an infinitesimally Hilbertian space. Then we have the following properties:
  • $D(\Delta)$ is a vector space.
  • The Laplacian $\Delta$ is linear.
  • The operator $\Delta(\cdot)$ is closed.
[…]
2.2. Bakry-Émery condition
Instead of giving the usual definition of an $\mathsf{RCD}-$space, we present the Bakry-Émery condition, originally introduced in [3] by Ambrosio, Gigli and Savaré; the reason being that this formulation enables us to prove that the gluing of cohomogeneity one $\mathsf{RCD}-$spaces is again an $\mathsf{RCD}-$space in Section 5.
Definition 2.8 (Bakry-Émery condition).
Let $(X,d,\mathfrak{m})$ be a metric measure space. We say that it satisfies the $\mathsf{BE}(K,N)$-condition for some $K\in\mathbb{R},$ $N\geq 1$, if for all $f\in D(\Delta)$ with $\Delta f\in W^{1,2}(X)$ and for all $g\in D(\Delta)$ non-negative, bounded with $\Delta g\in L^{\infty}(\mathfrak{m})$ we have
[…]
Satisfying the $\mathsf{BE}(K,N)$ condition is equivalent to the space satisfying the $\mathsf{CD}^{\ast}(K,N)$ for infinitesimally Hilbertian metric measure spaces $(X,d,\mathfrak{m})$, under the under the following assumptions:
  1. (1)
    The Sobolev-to-Lipschitz property is satisfied, that is, $\forall f\in W^{1,2}(X)$ with $|\nabla f|\leq 1$ there exists a Lipschitz representative of $f$.
  2. (2)
    For any $x_{0}\in X$ and $c>0$ we have
[…]
This fact was shown by Ambrosio, Gigli and Savaré for $N=\infty$ (under equivalent technical assumptions) in [3, Theorem 1.1] and for $N<\infty$ by Erbar, Kuwada and Sturm in [24, Theorem 7]. Note also the equivalence between the $\mathsf{CD}^{\ast}(K,N)$- and $\mathsf{CD}(K,N)$- conditions due to Cavalletti and Milman, [17, Corollary 13.7].
Since we only consider compact spaces with a Borel measure, it is clear that condition (2) above is automatically fulfilled.
From these considerations, we point out that we do not need to make use of the optimal transport formulation of the curvature-dimension condition, for example as stated in [40], so instead we settle on the following:
Definition 2.9 ($\mathsf{RCD}$-spaces).
Let $K\in\mathbb{R},N\geq 1$, we say that a metric measure space $(X,d,\mathfrak{m})$ is an $\mathsf{RCD}(K,N)$-space if it satisfies the $\mathsf{BE}(K,N)$ condition, is infinitesimally Hilbertian, satisfies Sobolev-to-Lipschitz (1), and the volume growth condition (2).
One important thing to mention is that in general the Bakry-Émery condition is strictly weaker than the $\mathsf{RCD}$-condition. In [47] Honda gives examples of spaces that satisfy a $\mathsf{BE}(K,N)$ condition but that are not $\mathsf{RCD}$-spaces. The reason being that they do not satisfy the Sobolev-to-Lipschitz property. Moreover, the Sobolev-to-Lipschitz property is essential as the $\mathsf{BE}(K,N)$-condition alone cannot imply anything about the geometry of the space as was pointed out to us by Gigli with the following examples. Any metric space equipped with an atomic measure satisfies $\mathsf{BE}(K,N)$ for any $K$ and $N$ as the Cheeger energy is identically 00. A more geometric example is the following: For a given compact smooth embedded submanifold $M$ in $\mathbb{R}^{n}$ one can consider the induced Riemannian metric $g$ from $\mathbb{R}^{n}$ (and in turn the Riemannian distance $d_{g}$) and Riemannian volume $d\mathrm{vol}_{g}$. Then by compactness, the metric measure space $(M,d_{g},d\mathrm{vol}_{g})$ is an $\mathsf{RCD}(K,N)$-space for adequate $K$ and $N$. Now, $M$ can be endowed with the restriction of the Euclidean distance as well, i.e., $d(x,y)=\left\|x-y\right\|$ (and the same measure $d\mathrm{vol}_{g}$). It is clear that this distance in general does not make $M$ a geodesic space and thus, this metric measure structure cannot satisfy the $\mathsf{RCD}$-condition in general. However, the Cheeger energies associated to both distances agree as can be seen directly from the definition above since the Lipschitz constants of functions almost agree for both distances since they almost coincide at small scales.
Often, if there is no room for confusion, we simply refer to metric measure spaces satisfying Definition 2.9 as $\mathsf{RCD}$-spaces without making any explicit mention of the parameters $K$ and $N$.
Denote by $\mathbf{X}$ the class of metric measure spaces $(X,d,\mathfrak{m})$ satisfying:
  1. (i)
    $(X,d)$ is complete and separable.
  2. (ii)
    $\mathfrak{m}$ is a Borel measure (i.e. defined on the Borel $\sigma$-algebra of open sets of $X$) with $\mathrm{supp}(\mathfrak{m})=X$, satisfying for all $r>0$ and $x\in X$:
    $\mathfrak{m}(B_{r}(x))\leq ce^{Ar^{2}},$
    for appropriate constants $A,c\geq 0$.
  3. (iii)
    $(X,d,\mathfrak{m})$ is infinitesimally Hilbertian.
Then, we have the following global-to-local and local-to-global results.
Theorem 2.10 (Global-to-Local for $\mathsf{RCD}(K,N)$, Proposition 7.7 in [4]).
Let $(X,d,\mathfrak{m})\in\mathbf{X}$ be an $\mathsf{RCD}(K,N)$-space, and let $U\subset X$ be open. If $\mathfrak{m}(\partial U)=0$ and $(\bar{U},d)$ is geodesic, then $(\bar{U},d,\mathfrak{m}\llcorner\bar{U})$ is an $\mathsf{RCD}(K,N)$-space.
Theorem 2.11 (Local-to-Global for $\mathsf{RCD}(K,N)$, Theorem 7.8 in [4]).
Let $(X,d,\mathfrak{m})\in\mathbf{X}$ be a length space and assume that there exists a covering $\{U_{i}\}_{i\in I}$ of $X$ by non-empty open subsets such that $\mathfrak{m}(U_{i})<\infty$ if $K<0$, and $(\overline{U}_{i},d,\mathfrak{m}\llcorner_{\overline{U}_{i}})\in X$ satisfy $\mathsf{RCD}(K,N)$. Then $(X,d,\mathfrak{m})$ is a $\mathsf{RCD}(K,N)$ space.
Let us describe some features on $\mathsf{RCD}$-spaces that will be useful to us. Proofs of them can be found for example in [24]. We remark that the tensorization property in item (2) below was originally shown to hold for the $\mathsf{BE}$-condition in [3].
Proposition 2.12.
Let $(X,d_{X},\mathfrak{m}_{X}),(Y,d_{Y},\mathfrak{m}_{Y})$ be $\mathsf{RCD}$-spaces that satisfy the curvature-dimension bounds $(K_{X},N_{X}),(K_{Y},N_{Y})$ respectively then:
  1. (1)
    $(X,d_{X},\mathfrak{m}_{X})$ is also an $\mathsf{RCD}(L,M)$ for all $L\leq K_{X},$ $M\geq N_{X}.$
  2. (2)
    The product $(X\times Y,d_{X\times Y},\mathfrak{m}_{X}\otimes\mathfrak{m}_{Y})$ is an $\mathsf{RCD}(\min\{K_{X},K_{Y}\},N_{X}+N_{Y})-$space.
  3. (3)
    If $\alpha,\beta>0$ then the metric measure space $(X,\alpha d_{X},\beta\mathfrak{m}_{X})$ is an $\mathsf{RCD}(\alpha^{-2}K_{X},N_{X}).$
One important feature of the class of $\mathsf{RCD}(K,N)$-spaces is that it is compact with respect to measured Gromov-Hausdorff convergence (see [2], [30], [32], [35], [56], [74], [75]).
Definition 2.13.
Let $(X,d,x),(Y,d_{Y},y)$ be complete and separable pointed metric spaces. Given $\epsilon>0$. We say that a map $f_{\epsilon}\colon B_{1/\epsilon}(x)\rightarrow B_{1/\epsilon}(y)$ is an $\epsilon$-Gromov-Hausdorff approximation if:
  1. (1)
    $f_{\epsilon}(x)=y$.
  2. (2)
    For all $u,v\in B_{1/\epsilon}(x)$ $|d(u,v)-d_{Y}(f_{\epsilon}(u),f_{\epsilon}(v))|<\epsilon$.
  3. (3)
    for all $y\in B_{1/\epsilon}(y)$ there exists some $x\in B_{1/\epsilon}(x)$ such that $d_{Y}(f_{\epsilon}(x),y)<\epsilon.$
Definition 2.14.
Let $\{(X_{n},d_{n},x_{n})\}_{n\in\mathbb{N}}$ be a sequence of complete separable pointed metric spaces. We say that the sequence converges in the pointed Gromov-Hausdorff sense to a complete separable pointed metric space $(Y,d_{Y},y)$ if there exists a sequence $\{\epsilon_{n}\}_{n\in\mathbb{N}}\subset\mathbb{R}$ such that $\epsilon_{n}\rightarrow 0$ as $n\rightarrow\infty$, and $\epsilon_{n}$-Gromov-Hausdorff approximations from $B_{1/\epsilon_{n}}(x)\subset(X_{n},d_{n},x_{n})$ to $B_{1/\varepsilon_{n}}(y)\subset(Y,d_{Y},y)$.
If in addition our spaces are equipped with non-negative Borel reference measures which are finite on balls we define:
Definition 2.15.
A sequence $\{(X_{n},d_{n},\mathfrak{m}_{n},x_{n})\}_{n\in\mathbb{N}}$ of complete pointed metric measure spaces converges in the pointed measured Gromov-Hausdorff sense to a complete pointed metric measure space $(Y,d_{Y},\mathfrak{m}_{Y},y)$ if it converges in the pointed Gromov-Hausdorff sense with the extra assumption that the $\epsilon_{n}$-Gromov-Hausdorff approximations are measurable and satisfy:
$f_{\epsilon\#}\mathfrak{m}_{n}\rightharpoonup\mathfrak{m}_{Y}.$
We abbreviate this by:
$(X_{n},d_{n},\mathfrak{m}_{n},x_{n})\xrightarrow{pmGH}(Y,d_{Y},\mathfrak{m}_{Y},y).$
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}$-space, $x\in X$, and some $r\in(0,1)$. Consider now the pointed, rescaled, and normalized metric measure space $(X,r^{-1}d,\mathfrak{m}^{x}_{r},x)$, where
$\mathfrak{m}^{x}_{r}:=\left(\int_{B_{r}(x)}1-\frac{d(x,y)}{r}d\mathfrak{m}(y)\right)^{-1}\mathfrak{m}.$
Using this notation we define:
Definition 2.16 (Tangent space).
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space for $K\in\mathbb{R}$ and $N\in[1,\infty)$. Given a point $x_{0}\in X$, we say that a pointed metric measure space $(Y,d_{Y},\mathfrak{m}_{Y},y)$ is a tangent of $X$ at $x_{0}$ if there exists a sequence $\{r_{i}\}_{i\in\mathbb{N}}\subset(0,1)$ such that $r_{i}\rightarrow 0$, and the sequence
$(X,r^{-1}_{i}d,\mathfrak{m}^{x_{0}}_{r_{i}},x_{0})\xrightarrow{pmGH}(Y,d_{Y},\mathfrak{m}_{Y},y).$
We denote the set of all tangents of $X$ at $x_{0}$ by $\mathrm{Tan}(X,d,\mathfrak{m},x_{0}).$
Remark 2.17.
We recall that if a space $(X,d,\mathfrak{m})$ is an $\mathsf{RCD}(K,N)$-space, then for $r,\lambda>0$ the space $(X,d/r,\lambda\mathfrak{m})$ is an $\mathsf{RCD}(r^{2}K,N)$-space. Since the $\mathsf{RCD}$-condition is stable under Gromov-Hausdorff limits, we have that tangents of $\mathsf{RCD}(K,N)$-spaces are $\mathsf{RCD}(0,N)$-spaces.
Remark 2.18.
Given an $\mathsf{RCD}$-space $(X,d,\mathfrak{m})$ and $x_{0}\in X$, the set $\mathrm{Tan}(X,d,\mathfrak{m},x_{0})$ is non-empty (see Proposition $2.2$ in [58]), but it may contain several spaces which may not even be metric cones (see Examples $8.41,8.80,8.95$ in [18]).
Given $m\in\mathbb{N}\cap[1,N]$, let $\mathcal{R}_{m}$ denote the set of points $x\in X$ such that $\mathrm{Tan}(X,d,\mathfrak{m},x)=\{(\mathbb{R}^{m},d_{\mathbb{E}},\mathcal{L}^{m},0)\}.$ Using regular Lagrangian flows, Brué and Semola in [12] proved that there is precisely a unique $m$ such that the set $\mathcal{R}_{m}$ has full measure:
Theorem 2.19 (Essential dimension).
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)-$space with $K\in\mathbb{R}$ and $N\in[1,\infty).$ Then there exists a unique $m\in\mathbb{N}\cap[1,N]$ such that $\mathfrak{m}(X-\mathcal{R}_{m})=0$. We define such $m$ to be the essential dimension of $(X,d,\mathfrak{m})$, and we denote it by $\mathrm{dim}_{\mathrm{ess}}(X)$. The set $\mathcal{R}_{m}$ is the set of regular points of $X$.
2.3. Non-collapsed spaces
Here we distinguish a particular class of $\mathsf{RCD}$-spaces: the so called non-collapsed spaces. These spaces were studied by De Philippis and Gigli in [23], and were further characterized in terms of properties of the $N$-dimensional Bishop-Gromov density by Brena, Gigli, Honda, and Zhu in [11].
Definition 2.20 (Non-collapsed $\mathsf{RCD}$-space).
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space. We say that it is non-collapsed if $\mathfrak{m}=\mathcal{H}^{N}$.
Thanks to the structure of $\mathsf{RCD}$-spaces, we can check that for non-collapsed spaces we have necessarily that $N\in\mathbb{N}$ (see Theorem $1.12$ in [23]). Non-collapsed $\mathsf{RCD}$-spaces also enjoy good structural properties which are stronger than those of more general $\mathsf{RCD}$-spaces. For example, we have the following result about the tangents (see Proposition $2.8$ in [23]):
Proposition 2.21.
Let $(X,d,\mathcal{H}^{N})$ be a non-collapsed $\mathsf{RCD}(K,N)$-space. Then the tangent spaces at every point are metric cones.
Regarding the essential dimension, we have the following result by Brena, Gigli, Honda, and Zhu [11, Theorem 1.7] which lets us identify non-collapsed spaces.
Theorem 2.22.
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space. Suppose that the essential dimension of some tangent space $(Y,d_{Y},\mathfrak{m}_{Y},y)$ at some $x\in X$ is equal to $N$, then $\mathfrak{m}=c\mathcal{H}^{N}$ for some $c>0$.
Finally, we recall that the essential dimension of a non-collapsed $\mathsf{RCD}(K,N)$-space is $N$ (see the last paragraph in Page 3 of [11]).
2.4. Warped products
Recall that given $(X,d_{X})$ a complete and separable metric space, a curve $\gamma\colon[0,1]\to X$ is absolutely continuous (AC), if there exists a map $f\in L^{1}[0,1]$ such that for every $0\leq s<t\leq 1$ we have
[…]
Given an AC curve $\gamma\colon[0,1]\to X$, for a.e. $t_{0}\in[0,1]$ fixed the following limit exits
$|\dot{\gamma}(t_{0})|:=\lim_{h\to 0}\frac{d_{X}(\gamma(t_{0}+h),\gamma(t_{0})}{|h|}.$
Moreover the function $t\mapsto|\dot{\gamma}(t)|$ called metric speed of $\gamma$ is in $L^{1}(0,1)$, and is the minimal function satisfying (3).
Let $(B,d_{B})$ and $(F,d_{F},\mathfrak{m}_{F})$ be complete, locally compact geodesic metric spaces. Consider $f\colon B\to\mathbb{R}_{\geq 0}$ a locally Lipschitz function. A curve $\gamma=(\alpha,\beta)\colon[0,1]\to B\times F$ is admissible if $\alpha$ and $\beta$ are absolutely continuous in $B$ and $F$ respectively. For an admissible curve $\gamma\colon[0,1]\to B\times F$ we define the length with respect to $f$ as
[…]
For convenience we define the length of a curve that is not as admissible simply as +.$\infty$.+ ∞ . $L_{f}$ is a length-structure on the class of admissible curves. We define a pseudo-metric $\tilde{d}_{f}$ on $B\times F$, between $(p,x)$ and $(q,y)$ by setting
$\tilde{d}_{f}((p,x),(q,y)):=\inf\{L_{f}(\gamma)\mid\gamma\mbox{ admisible and connects }(p,x)\mbox{ to }(q,y)\}.$
Then we have an induced equivalence relation $\sim$, where $(p,x)\sim(q,y)$ if and only if $\tilde{d}_{f}((p,x),(q,y))=0$. We define the warped product of $B$ and $F$ via $f$ as $B\times_{f}F:=B\times F/\sim$ with the distance $d_{f}([p,x],[q,y]):=\tilde{d}_{f}((p,x),(q,y))$. By definition $B\times_{f}F$ is an intrinsic metric space, and since $B$ and $F$ are complete and locally compact, then the warped product is complete and locally compact. As pointed out in [51, Remark 2.4], the warped product $B\times_{f}F$ is geodesic.
Moreover, as pointed out in [1] we have the following description of geodesics
Proposition 2.23 (Propsition 4.5 (2)).
Let $\gamma=(\alpha,\beta)$ be a geodesic in $B\times_{f}F$ such that $f\circ\alpha>0$. Then $\beta$ is a pregodesic in $F$. That is, up to a reparametrization $\beta$ is a geodesic in $F$.
Lemma 2.24.
Let $F$ be a compact Riemannian manifold. Consider two functions
$f,g\colon[0,\infty)\to F$
such that $f(0)=0=g(0)$, and for $t>0$ we have $f(t),g(t)>0$. Consider $0<s_{1}<s_{2}$, and $p,q\in F$ are such that there exists a unique geodesic between them. Then the trace of the geodesic $\gamma_{f}$ from $(s_{1},p)$ to $(s_{2},q)$ in $[0,\infty)\times_{f}F$ is the same as the trace of the geodesic $\gamma_{g}$ from $(s_{1},p)$ to $(s_{2},q)$ in $[0,\infty)\times_{g}F$.
Proof.
We write the geodesic $\gamma_{f}=(\alpha_{f},\beta_{f})$, and observe that by Proposition 2.23 $\beta_{f}$ is a pregeodesic in $F$. The same holds for the geodesic $\gamma_{g}=(\alpha_{g},\beta_{g})$ in $[0,\infty)\times_{g}F$. Thus $\beta_{f}$ and $\beta_{g}$ are two pregeodesics in $F$ joining $p$ to $q$. Since $p$ and $q$ are not conjugate this implies that the trace of $\beta_{f}$ is the same as the trace of $\beta_{g}$. ∎
We consider $(X,d_{X},\mathfrak{m}_{X})$ and $(Y,d_{Y},\mathfrak{m}_{Y})$ two complete separable length spaces, and $\mathfrak{m}_{X}$ and $\mathfrak{m}_{Y}$ Radon measures. We also assume that $\mathfrak{m}_{X}(X)<\infty$. We consider continuous functions $\omega_{d},\omega_{\mathfrak{m}}\colon Y\to[0,\infty)$. For $\omega=(\omega_{d},\omega_{\mathfrak{m}})$ define the $\omega$-warped product $Y\times_{\omega}X$ as the metric space $Y\times_{\omega_{d}}X$, and set the measure
$\mathfrak{m}_{\omega}:=\pi_{\ast}((\omega_{\mathfrak{m}}\mathfrak{m}_{Y})\otimes\mathfrak{m}_{X}),$
where $\pi\colon Y\times X\to(Y\times_{\omega_{d}}X)$ is the quotient map.
When $\omega_{d}=f$, and $\omega_{\mathfrak{m}}=f^{N}$ for $f\colon Y\to[0,\infty)$ a continuous function, we set $Y\times_{f}^{N}X:=Y\times_{(f,f^{N})}X$. The space $Y\times_{f}^{N}X$ is referred to as the $N$-warped product of $Y$, $X$ and $f$ (see [51]).
We also define here the $(K,N)$-cones for $K\in\mathbb{R}$, $N\in(0,\infty)$: Given a metric measure space $(F,d_{F},\mathfrak{m}_{F})$, the $(K,N)$-cone of $F$, denoted by $\mathrm{Con}^{N}_{K}(F)$ is the metric measure space $(\mathrm{Con}_{K}(F),d_{\mathrm{Con}_{K}},\mathfrak{m}^{N}_{\mathrm{Con}_{K}})$ defined by
$\mathrm{Con}_{K}(F)=\begin{cases}F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})&\mbox{if }K>0,\\ F\times[0,\infty)/(F\times\{0\})&\mbox{if }K\leq 0.\end{cases}$
$d_{\mathrm{Con}_{K}}([p,t],[q,s])$
:=$:=\begin{cases}\cos^{-1}_{K}\Big{(}\cos_{K}(t)\cos_{K}(t)+K\sin_{K}(s)\sin_{K}(t)\cos_{K}\big{(}d_{F}(p,q)\wedge\pi\big{)}\Big{)}&\mbox{if }K\neq 0,\\ \sqrt{s^{2}+t^{2}-2st\cos\big{(}d_{F}(p,q)\wedge\pi\big{)}}&\mbox{if }K=0,\end{cases}$
and
$\mathfrak{m}_{\mathrm{Con}_{K}}^{N}:=\sin_{K}^{N}(t)dt\otimes\mathfrak{m}_{F}.$
Here, for $I_{K}=[0,\pi/\sqrt{K}]$ for $K>0$ and $[0,\infty)$ for $K\leq 0$, we have $\cos_{K}\colon I_{K}\to[0,\infty)$ given by
$\cos_{K}(t):=\begin{cases}\cos\left(\sqrt{K}t\right)&\mbox{if }K>0,\\ \cosh\left(\sqrt{-K}t\right)&\mbox{if }K<0.\end{cases}$
and $\sin_{K}\colon I_{K}\to[0,\infty)$ given by
$\sin_{K}(t):=\begin{cases}\frac{1}{\sqrt{K}}\sin\left(\sqrt{K}t\right)&\mbox{if }K>0,\\ t&K=0,\\ \frac{1}{\sqrt{|K|}}\sinh\left(\sqrt{|K|}t\right)&\mbox{if }K<0.\\ \end{cases}$
Remark 2.25.
If $\operatorname{diam}(F)\leq\pi$ and $F$ is a length space, then the $(K,N)$-cone $\mathrm{Con}^{N}_{K}(F)$ coincides with the $N$-warped product $I_{K}\times_{\sin_{K}}^{N}F$.
2.5. Transformation groups
In order to set the notation, let us recall the basic facts of the theory of transformation groups that we use throughout the manuscript.
Let $(X,d,\mathfrak{m})$ be a metric measure space, and $G$ a topological group. A continuous (left) action of $G$ on $X$ is a continuous map $\alpha\colon G\times X\to X$ such that for any $g,h\in G$ and $x\in X$ we have $\alpha(g,\alpha(h,x))=\alpha(gh,x)$, and for $e\in G$ the identity element and any $x\in X$ we have $\alpha(e,x)=x$. The isotropy group at $x\in X$ is the closed subgroup $G_{x}=\left\{g\in G\mid gx=x\right\}$. We say the action $\alpha$ is effective if the only element of $G$ that fixes all the points in $X$ is the identity element, i.e. $\cap_{x\in X}G_{x}=\{e\}$. The action is proper if the map $A\colon G\times X\to X\times X$ given by $A(g,x)=(gx,x)$ is a proper map (that is, the preimage $A^{-1}(K)\subset G\times X$ of a compact subset $K\subset X\times X$ is compact. The orbit of a point $x\in X$ is the subset of $X$ given by $G(x)=\left\{gx\in X\mid g\in G\right\}$. The orbit type of an orbit $G(x)$ is the set $(G_{x}):=\{gG_{x}g^{-1}\mid g\in G\}$ of the conjugation class of the isotropy.
The orbit space is denoted by $X/G$ and $\pi\colon X\to X/G$ is the natural projection. We equip $X/G$ with the quotient topology. We also use the notation $U^{\ast}:=\pi(U)$, for any $U\subset X$. In particular, $X^{\ast}=X/G$ and $\pi(x)=x^{\ast}$ for each $x\in X$.
When we have that $d(\alpha(g,x),\alpha(g,y))=d(x,y)$ for any $g\in G$ and any $x,y\in X$ we say that the action $\alpha$ is by isometries. Given $g\in G$ fixed we have a homeomorpshism $\alpha_{g}\colon X\to X$ given as $\alpha_{g}(x)=\alpha(g,x)$. In the case when the action is by isometries and $(\alpha_{g})_{\#}\mathfrak{m}=\mathfrak{m}$ for any $g\in G$, we say that the action is by measure-preserving isometries, and that the metric $d$ is $G$-invariant. For simplicity, unless otherwise stated, we denote the action as $gx:=\alpha(g,x)$. In the case the orbits are compact, the orbit space is equipped with the orbital metric
$d^{\ast}(x^{\ast},y^{\ast}):=d(G(x),G(y))$
and the push-forward measure $\mathfrak{m}^{\ast}=\pi_{\#}\mathfrak{m}$.
The following theorem shows that a sufficient condition for the orbits to be compact is that the action is proper. Moreover, properness gives us the description of each orbit as a homogeneous space.
Theorem 2.26.
Let $G$ be a Hausdorff topological group acting properly on a metric space $(X,d)$, where $d$ is $G$-invariant. Fix $x\in X$, and consider $\alpha_{x}\colon G\to X$ given by $\alpha_{x}(g)=gx$. Let $\rho\colon G\to G/G_{x}$ be the quotient map. Then there exists a $G$-equivariant homeomorphism $\tilde{\alpha}_{x}\colon G/G_{x}\to G(x)$ onto the orbit through $x$ such that $\tilde{\alpha}_{x}\circ\rho=\alpha_{x}$. Furthermore, the orbit $G(x)$ is a closed subspace of $X$.
Remark 2.27.
In the case when $G$ is a compact Lie group, an action of $G$ on a metric space $X$ is proper (see for example [21, Theorem 2.2]). Thus from now on, we only consider compact Lie groups.
We recall from [27, Theorem 1.3] that given an $\mathsf{RCD}$-space $(X,d,\mathfrak{m})$ with an action by measure preserving isometries by a compact Lie group $G$, there exist (up to conjugation) a unique subgroup $G_{\min}<G$, such that the orbit $G(y)$ of $\mathfrak{m}$-a.e $y\in M$ is homeomorphic to the quotient $G/G_{\min}$. We call points $y\in M$ such that $G_{y}$ is conjugate to $G_{\min}$ in $G$ points with principal orbit type. The orbit of a point of principal orbit type is referred to as a principal orbit.
The following notion of regularity for the action was introduced in [27] and will also be useful to us.
Definition 2.28.
We say $G$ acts Lipschitz and co-Lipschitz continuously on a metric space $(X,d)$, if for a bi-invariant metric on $G$, and every $y\in M$ with principal orbit type we have constants $C,R>0$ such that for all $0<r<R$ it holds
$B_{r/C}(y)\cap G(y)\subset\{gy\mid g\in B_{r}(e)\subset G\}\subset B_{Cr}(y)\cap G(y)$
In other words, the map $\alpha_{y}\colon G\to G(y)$ mapping $g$ to $gy$ is locally Lipschitz and co-Lipschitz continuous.
For an action of $G$ on $X$, given $x\in X$ fixed, a slice through $x$ is a subset $S_{x}\subset X$ that satisfies:
  1. (i)
    $gS_{x}\subset S_{x},\quad\forall g\in G_{x}$.
  2. (ii)
    If $gS_{x}\cap S_{x}\neq\varnothing$, then $g\in G_{x}$.
  3. (iii)
    Setting $H=G_{x}$, there exists an open neighborhood $U\subset G/H$ around $eH$, and a cross-section $\chi\colon U\to G$ such that the function $F\colon U\times S_{x}\to X$ given by
    $F(gH,s)=\chi(gH)s,$
    is an homeomorphism onto its image.
We now state the main recognition tool in the theory of transformation groups, the so called Slice Theorem (see [10, Definition 4.1, Theorem 4.2]).
Theorem 2.29 (Slice Theorem).
Let a compact Lie group $G$ act by isometries on a metric space $(X,d)$. For all $x\in X$ for which a slice $S_{x}$ exists, there is some $r_{0}>0$ such that for all $r<r_{0}$ there is an equivariant homeomorphism
$G\times_{G_{x}}S_{x}\to B_{r}(G(x))$
Every point in a completely regular space has a slice (see [65]). This fact is also referred to sometimes as the Slice Theorem. It follows then that every point on an $\mathsf{RCD}$-space admits a slice. However, given an Alexandrov space $X$ and $G$ a Lie group acting by isometries, for each point $x\in X$ we can find slice containing $x$ which is a metric cone over a positively curved Alexandrov space (namely the space of normal directions at the point in question; see the the Slice Theorem for Alexandrov spaces due to Harvey and Searle, [44, Theorem B]). This implies that for any point in an Alexandrov space equipped with a compact Lie group action by isometries, there is a slice through the given point which in itself admits an Alexandrov structure.
Moreover, for the case when the space is a Riemannian manifold with a compact Lie group action by isometries the same holds: for any point there exists a slice containing the point, such that it admits a Riemannian space structure.
Thus in general, when considering a geometric space and a group action that preserve the geometric structure, we are interested in finding in a well defined way a slice which belongs to the same class of geometric objects as the global space.
2.6. Equivariant Gromov-Hausdorff convergence
We now define the notion of equivariant Gromov-Hausdorff convergence (see [25]). Given some $r>0$ and $G$ a group acting by isometries on a pointed metric space $(X,d,x)$ we define:
$G(r):=\{g\in G\,|\,gx\in B_{r}(x)\}.$
Definition 2.30.
Let $(X,d,x,G),$ $(Y,d_{Y},y,H)$ be two pointed metric spaces with $G$, $H$ groups acting isometrically on $X$ and $Y$ respectively. An equivariant $\epsilon$-Gromov-Hausdorff approximation is a triple $(f_{\epsilon},\varphi,\psi)$ which consists of maps
$f_{\epsilon}\colon B_{1/\epsilon}(x)\rightarrow B_{1/\epsilon}(y),\quad\varphi\colon G(1/\epsilon)\rightarrow H(1/\epsilon),\quad\psi\colon H(1/\epsilon)\rightarrow G(1/\epsilon)$
such that:
  1. (1)
    $f_{\epsilon}$ is an $\epsilon$-Gromov-Hausdorff approximation.
  2. (2)
    If $g\in G(1/\epsilon)$ and both $p,gp\in B_{1/\epsilon}(x)$, then
    $d_{Y}(f_{\epsilon}(gp),\varphi(g)f_{\epsilon}(p)).$
  3. (3)
    If $h\in H(1/\epsilon)$ and both $p,\psi(h)p\in B_{1/\epsilon}(x)$, then
    $d_{Y}(f_{\epsilon}(\psi(h)p),hf_{\epsilon}(p)).$
Definition 2.31.
A sequence $\{(X_{n},d_{n},x_{n},G_{n})\}_{n\in\mathbb{N}}$ is be said to converge in the equivariant pointed Gromov-Hausdorff sense to a pointed metric space $(Y,d_{Y},y,H)$ if there exist equivariant $\epsilon_{n}$-Gromov-Hausdorff approximations such that $\epsilon_{n}\rightarrow 0$ as $n\rightarrow\infty$.
3. Slice Theorem for cohomogeneity one geodesic spaces
In this Section we characterize the slices at each point of a compact, metric space $(X,d)$ admitting an action of a compact Lie group $G$ by isometries, having $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$ as orbit space, in terms of certain sets from the theory of optimal transport.
It is worth pointing out, however that we make no assumption of curvature conditions in this Section. To proceed, we recall a few notions from the theory of optimal transport (see for example the survey [16] for a brief account of the theory).
For each Lipschitz function $u\colon X\to\mathbb{R}$, we consider the subsets of $X\times X$ given by
$\Gamma_{u}$ :=$:=\left\{(x,y)\mid u(x)-u(y)=d(x,y)\right\},$
$\Gamma^{-}_{u}$ :=$:=\left\{(x,y)\mid(y,x)\in\Gamma_{u}\right\}.$
From these, we also define
$\mathcal{T}_{u}=\mathrm{pr}_{1}(\Gamma_{u}\smallsetminus\{x=y\})\cup\mathrm{pr}_{1}(\Gamma^{-}_{u}\smallsetminus\{x=y\}),$
where $\mathrm{pr}_{1}\colon X\times X\to X$ is given by $\mathrm{pr}_{1}(x,y)=x$. Now, for each $x\in X$ we set
$\Gamma_{u}(x)$ =$=\{y\in X\mid(x,y)\in\Gamma_{u}\},$
$\Gamma^{-}_{u}(x)$ =$=\{y\in X\mid(x,y)\in\Gamma^{-}_{u}\},$
and define an equivalence relation $R_{u}=\Gamma_{u}\cup\Gamma_{u}^{-}$.
We now consider the sets
$A_{+}(u)$ =$=\{y\in\mathcal{T}_{u}\mid\exists z,w\in\Gamma_{u}(y),\ (z,w)\not\in R_{u}\},$
$A_{-}(u)$ =$=\{y\in\mathcal{T}_{u}\mid\exists z,w\in\Gamma_{u}^{-}(y),\ (z,w)\not\in R_{u}\}.$
The non-branching transport set of $u$ is defined as $\mathcal{T}^{nb}_{u}=\mathcal{T}_{u}\setminus(A_{+}(u)\cup A_{-}(u))$.
We now state the main result of this Section, containing the definition of our distinguished slices.
Theorem C.
Let $(X,d)$ be a geodesic space and $G$ a compact Lie group acting on $X$ by isometries such that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then for any $x_{0}\in X$ we define $u(y):=d(y,G(x_{0}))$, and there exists $\delta_{0}>0$ such that for any $0<\delta<\delta_{0}$ we have $B_{\delta}(x_{0})\subset\mathcal{T}^{nb}_{u}$ and
$S_{x_{0}}:=\left(\Gamma^{-}_{u}(x_{0})\cap B_{\delta}(x_{0})\right)\cup\{x_{0}\}$
is a slice through $x_{0}$.
Remark 3.1.
We point out that $S_{x_{0}}$ is homeomorphic to a topological cone, since it consists of a union of geodesics, such that the only common point of two geodesics is the vertex.
To show the validity of Theorem C, we need to prove some results first. Assume now that we have a compact Lie group $G$ acting effectively by isometries on $(X,d)$. Let $u\colon X\to\mathbb{R}$ be a Lipschitz function which is $G$-invariant, i.e. for $x\in X$ we have $u(gx)=u(x)$ for any $g\in G$. Observe that we have a well defined Lipschitz function $u^{\ast}\colon(X/G,d^{\ast})\to\mathbb{R}$ given by $u^{\ast}(x^{\ast}):=u(x)$. Observe that $u^{\ast}$ is well defined, since for $g\in G$ we have that $u(gx)=u(x)$ by hypothesis. Thus $u^{\ast}(x^{\ast})$ does not depend on the representative $x$ of $x^{\ast}$.
Lemma 3.2.
Let $x_{0}\in X$ be fixed and $u\colon X\to\mathbb{R}$ be the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$. Then
$\left\{(x^{\ast},y^{\ast})\mid(x,y)\in\Gamma_{u}\right\}\subseteq\Gamma_{u^{\ast}}.$
Moreover, for every $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$, there exists $g\in G$ such that $(gx,y)\in\Gamma_{u}$ and $d^{\ast}(x^{\ast},y^{\ast})=d(gx,y)$. In an analogous fashion, due to the symmetry in the definition of $\Gamma_{u}^{-}$, the same holds for $\Gamma_{u}^{-}$.
Proof.
Consider $(x,y)\in\Gamma_{u}$. Observe that $d(G(x_{0}),x)=d^{\ast}(x_{0}^{\ast},x^{\ast})$, and thus
$d^{\ast}(x^{\ast},y^{\ast})\leq d(x,y)$ $=u(x)-u(y)$
$=u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})$
$=d^{\ast}(x_{0}^{\ast},x^{\ast})-d^{\ast}(x_{0}^{\ast},y^{\ast})$
$\leq d^{\ast}(x_{0}^{\ast},y^{\ast})+d^{\ast}(y^{\ast},x^{\ast})-d^{\ast}(x_{0}^{\ast},y^{\ast})$
$=d^{\ast}(x^{\ast},y^{\ast}).$
Thus we conclude that $d^{\ast}(x^{\ast},y^{\ast})=u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})$, and therefore $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$, proving the first claim.
We now consider $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$. Then we have that for any $\bar{g},\bar{h}\in G$ it holds that
$d^{\ast}(x^{\ast},y^{\ast})=u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})=u(\bar{g}x)-u(\bar{h}y).$
Also observe that there exists $g\in G$ such that $d^{\ast}(x^{\ast},y^{\ast})=d(gx,y)$. Thus by setting $\bar{g}=g$ and $\bar{h}=e$ we have
$d(gx,y)=d^{\ast}(x^{\ast},y^{\ast})=u(gx)-u(y).$
This implies that $(gx,y)\in\Gamma_{u}$. ∎
Lemma 3.3.
Let $x_{0}\in X$ be fixed and $u\colon X\to\mathbb{R}$ the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$. Then the following hold,
  • (i)
    $\mathcal{T}_{u}^{\ast}=\mathcal{T}_{u^{\ast}}$,
  • (ii)
    $\Gamma_{u^{\ast}}(x^{\ast})=\Gamma_{u}(x)^{\ast}$ and $\Gamma^{-}_{u^{\ast}}(x^{\ast})=\Gamma^{-}_{u}(x)^{\ast}$ for every $x\in X$,
  • (iii)
    $A_{+}(u)^{\ast}=A_{+}(u^{\ast})$ and $A_{-}(u)^{\ast}=A_{-}(u^{\ast})$.
  • (iv)
    $(\mathcal{T}^{nb}_{u})^{\ast}=\mathcal{T}^{nb}_{u^{\ast}}$.
Moreover, if $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$, then $\mathcal{T}^{nb}_{u^{\ast}}$ is equal to $X^{\ast}$ in the first three cases or $\mathbf{S}^{1}\setminus\left\{-x_{0}^{\ast}\right\}$ respectively, where $-x_{0}^{\ast}$ denotes the antipodal orbit to $x_{0}^{\ast}$.
Proof.
First we prove that $\mathcal{T}_{u}^{\ast}\subseteq\mathcal{T}_{u^{\ast}}$. Consider $x^{\ast}\in\mathcal{T}^{\ast}_{u}$, for some $x\in\mathcal{T}_{u}$. Then there exists $y\in X$ with $x\neq y$ and $(x,y)\in\Gamma_{u}\cup\Gamma^{-}_{u}$. We then have two cases depending on whether $(x,y)\in\Gamma_{u}$ or $(x,y)\in\Gamma^{-}_{u}$.
Assume first that $(x,y)\in\Gamma_{u}$. Then, by Lemma 3.2, $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$. Now, suppose, by contradiction, that $x^{\ast}=y^{\ast}$, that is, $y=gx$ for some $g\in G$. Then we have that
$d(x,y)=d(x,gx)=d(G(x_{0}),x)-d(G(x_{0}),gx)=0.$
Thus $y=x$, which is a contradiction. Hence $y\not\in G(x)$, and therefore $x^{\ast}\neq y^{\ast}$. Thus we have that $x^{\ast}\in\mathcal{T}_{u^{\ast}}$. The proof in the case that $(x,y)\in\Gamma^{-}_{u}$ is analogous. Thus, we have shown that $\mathcal{T}_{u}^{\ast}\subseteq\mathcal{T}_{u^{\ast}}$.
To prove the opposite inclusion, now consider $x^{\ast}\in\mathcal{T}_{u^{\ast}}$. Then there is $y^{\ast}\in X^{\ast}$ such that $y^{\ast}\neq x^{\ast}$ and $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}\cup\Gamma_{u^{\ast}}^{-}$. Then by Lemma 3.2, for some representatives $\tilde{x}\in G(x)$ of $x^{\ast}$ and $\tilde{y}\in G(y)$ of $y^{\ast}$, we have $(\tilde{x},\tilde{y})\in R_{u}$: We assume first that $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$. Then it follows that for some $g\in G$ we have $(gx,y)\in\Gamma_{u}$. In the case where $(x^{\ast},y^{\ast})\in\Gamma^{-}_{u^{\ast}}$ we have then that there exists $g^{\prime}\in G$ such that $(x,g^{\prime}y)\in\Gamma_{u}^{-}$. Observe that if $\tilde{x}=\tilde{y}$ we have that $G(x)\cap G(y)\neq\varnothing$, which is a contradiction. Thus we conclude that $\tilde{x}\in\mathcal{T}_{u}$, whence $\mathcal{T}_{u}^{\ast}\supseteq\mathcal{T}_{u^{\ast}}$. Thus we have that $\mathcal{T}_{u}^{\ast}=\mathcal{T}_{u^{\ast}}$, as claimed.
We now show that $\Gamma_{u^{\ast}}(x^{\ast})\subseteq\Gamma_{u}(x)^{\ast}$. Consider $y^{\ast}\in\Gamma_{u^{\ast}}(x^{\ast})$. This implies that there exist representatives $\tilde{x}\in G(x)$ of $x^{\ast}$ and $\tilde{y}\in G(y)$ of $y^{\ast}$ such that $(\tilde{x},\tilde{y})\in\Gamma_{u}$. In particular $\tilde{y}\in\Gamma_{u}(\tilde{x})$. Now assume that $\tilde{x}=gx$ for some $g\in G$, then $g^{-1}\tilde{y}\in\Gamma_{u}(x)$. Thus we conclude that $y^{\ast}\in\Gamma_{u}(x)^{\ast}$. This shows that $\Gamma_{u^{\ast}}(x^{\ast})\subseteq\Gamma_{u}(x)^{\ast}$.
For the other inclusion, given $y\in\Gamma_{u}(x)$, we have that $(x,y)\in\Gamma_{u}$. Thus $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$ by Lemma 3.2 and we conclude that $y^{\ast}\in\Gamma_{u^{\ast}}(x^{\ast})$. Thus the claim that $\Gamma_{u^{\ast}}(x^{\ast})=\Gamma_{u}(x)^{\ast}$ for every $x\in X$ is proved. An analogous argument shows that $\Gamma^{-}_{u^{\ast}}(x^{\ast})=\Gamma^{-}_{u}(x)^{\ast}$ for every $x\in X$.
Now we show that $A_{+}(u)^{\ast}=A_{+}(u^{\ast})$ and $A_{-}(u)^{\ast}=A_{-}(u^{\ast})$. We prove this for $A_{+}(u)$ since the argument for $A_{-}(u)$ is analogous.
Consider $y^{\ast}\in A_{+}(u^{\ast})$. Then there exists $z^{\ast},w^{\ast}\in\Gamma_{u^{\ast}}(y^{\ast})$ such that $(z^{\ast},w^{\ast})\not\in R_{u^{\ast}}$. Fix a representative $\tilde{y}\in G(y)$ of $y^{\ast}$. Then there exist representatives $\tilde{z}\in G(z)$ of $z^{\ast}$ and $\tilde{w}\in G(w)$ of $w^{\ast}$ with $\tilde{z},\tilde{w}\in\Gamma_{u}(\tilde{y})$. If we assume that $(\tilde{z},\tilde{w})\in R_{u}$ it would follow from Lemma 3.2 that $(z^{\ast},w^{\ast})\in R_{u^{\ast}}$, which is a contradiction. Thus we have that $\tilde{z},\tilde{w}\in\Gamma_{u}(\tilde{y})$ and $(\tilde{z},\tilde{w})\not\in R_{u}$, i.e. $\tilde{y}\in A_{+}(u)$. This implies that $y^{\ast}\in A_{+}(u)^{\ast}$, whence, $A_{+}(u^{\ast})\subseteq A_{+}(u)^{\ast}$.
To prove the opposite inclusion, given $y^{\ast}\in A_{+}(u)^{\ast}$ for some $y\in A_{+}(u)$, we have that there exists $z,w\in\Gamma_{u}(y)$ with $(z,w)\not\in R_{u}$. Then $z^{\ast},w^{\ast}\in\Gamma_{u^{\ast}}(y^{\ast})$ and by Lemma 3.2 we obtain $(z^{\ast},w^{\ast})\not\in R_{u^{\ast}}$, i.e. $y^{\ast}\in A_{+}(u^{\ast})$. Thus we have shown that $A_{+}(u)^{\ast}\subseteq A_{+}(u^{\ast})$, and in turn that $A_{+}(u)^{\ast}=A_{+}(u^{\ast})$.
Now it readily follows from the definition of the non-branching transport set and the previous paragraphs that
$\mathcal{T}^{nb}_{u^{\ast}}=\mathcal{T}_{u^{\ast}}\setminus\left(A_{+}(u^{\ast})\cup A_{-}(u^{\ast})\right)=\mathcal{T}_{u}^{\ast}\setminus(A_{+}(u)^{\ast}\cup A_{-}(u)^{\ast})=(\mathcal{T}^{nb}_{u})^{\ast},$
obtaining the desired conclusion.
For the final claim of the Lemma we only explain the case when $X^{\ast}$ is isometric to $[-1,1]$ and $x_{0}^{\ast}=-1$ as all other cases are very similar. Let us observe that when $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$ we have that if $x_{0}^{\ast}=-1$, then $u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})=d^{\ast}(x^{\ast},y^{\ast})$ if and only if $y^{\ast}\leq x^{\ast}$. Thus we conclude that,
$\Gamma_{u^{\ast}}$ =$=\{(s,t)\in[-1,1]\times[-1,1]\mid t\leq s\},$
$\Gamma_{u^{\ast}}^{-}$ =$=\{(s,t)\in[-1,1]\times[-1,1]\mid s\leq t\},$
It is straightforward to verify that $\mathcal{T}_{u^{\ast}}=[-1,1]$. From this it follows that $A_{+}(u^{\ast})=\varnothing=A_{-}(u^{\ast})$, and finally that $\mathcal{T}^{nb}_{u^{\ast}}=[-1,1]$.
The case of $X^{\ast}=\mathbf{S}^{1}$ follows similarly and directly applying [16, Lemma 3.4.2]. ∎
We state the following straightforward lemma without proof.
Lemma 3.4.
Let $u\colon X\to\mathbb{R}$ be the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$. The sets $\Gamma_{u}$, $\Gamma_{u}^{-}$, $\mathcal{T}_{u}$, $A_{+}(u)$, $A_{-}(u)$, $\Gamma_{u}(x)$, and $\Gamma_{u}^{-}(x)$ are $G$-invariant. That is if a point is contained in any of these sets, then their orbit is contained in the set.
We now obtain the following corollary from Lemma 3.3.
Corollary 3.5.
Let $G$ be a Lie group acting by isometries on a metric space $(X,d)$ so that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Let $u\colon X\to\mathbb{R}$ the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$. Then the transport set $\mathcal{T}^{nb}_{u}$ is open.
Proof.
By Lemma 3.3 we have that $(\mathcal{T}^{nb}_{u})^{\ast}$ is $[-1,1]$, $[0,\infty)$ or $\mathbb{R}$, if $X^{\ast}$ is isometric to $[-1,1]$ $[0,\infty)$ or $\mathbb{R}$ respectively, and $(\mathcal{T}^{nb}_{u})^{\ast}$ is $\mathbf{S}^{1}\setminus\{-x_{0}^{\ast}\}$ if $X^{\ast}$ is isometric to $\mathbf{S}^{1}$. Moreover, $(\mathcal{T}^{nb}_{u})^{\ast}=\mathcal{T}^{nb}_{u^{\ast}}$.
In all cases, $y\in\mathcal{T}^{nb}_{u}$ we can find $\varepsilon>0$ small enough so that $B_{\varepsilon}(y^{\ast})\subset\mathcal{T}^{nb}_{u^{\ast}}$. We observe that for every $x^{\ast}\in B_{\varepsilon}(y^{\ast})$ there exists $x\in\pi^{-1}(B_{\varepsilon}(y^{\ast}))$ with $x\in\mathcal{T}^{nb}_{u}$. Thus we we have $\pi^{-1}(x^{\ast})=G(x)\subset\mathcal{T}^{nb}_{u}$, and $G(x)=\pi^{-1}(x^{\ast})\subset\pi^{-1}(B_{\varepsilon}(y^{\ast}))$. This implies that $\pi^{-1}(B_{\varepsilon}(y^{\ast}))\subset\mathcal{T}^{nb}_{u}$. But since we have that the topology of the metric $d^{\ast}$ agrees with the quotient topology, we have that $\pi^{-1}(B_{\varepsilon}(y^{\ast}))\subset X$ is open. This implies that there exists $\delta>0$ such that $B_{\delta}(y)\subset\pi^{-1}(B_{\varepsilon}(y^{\ast}))\subset\mathcal{T}^{nb}_{u}$. Thus we conclude that $\mathcal{T}^{nb}_{u}$ is open. ∎
From this we easily conclude the following corollary.
Corollary 3.6.
Let $G$ be a compact Lie group acting by isometries on a metric space $(X,d)$ so that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, or $\mathbb{R}$. For $u\colon X\to\mathbb{R}$ the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$, the transport set $\mathcal{T}^{nb}_{u}$ is $X$.
With these results established we are in position to prove Theorem C.
Proof of Theorem C.
Without loss of generality, we may assume that $X^{\ast}=[-1,1]$. Fix $x_{0}\in X$ and take $\delta>0$, such that $B_{\delta}(x_{0})\subset\mathcal{T}^{nb}_{u}$, and $B_{\delta}(G(x_{0}))\setminus G(x_{0})$ consists of only principal orbits, that is, $\pi(B_{\delta}(G(x_{0}))\setminus G(x_{0}))$ is homeomorphic to $(0,\delta)$ or $(-\delta,0)\cup(0,\delta)$. We claim that
(Slice) $S_{x_{0}}:=(\Gamma^{-}_{u}(x_{0})\cap B_{\delta}(x_{0}))\cup\{x_{0}\}.$
is a slice through $x_{0}$. In order to prove so, we check each item in the definition of Slice as recalled in Section 2.
Point (i). Take $g\in G_{x_{0}}$, and $y\in S_{x_{0}}$. Since $u(x_{0})=0$ we conclude that $(x_{0},y)\in\Gamma_{u}^{-}$ and in turn that $(y,x_{0})\in\Gamma_{u}$. That is, $d(G(x_{0}),y)=u(y)=d(x_{0},y)$. This implies that
$d(x_{0},gy)=d(g^{-1}x_{0},y)=d(x_{0},y)=d(G(x_{0}),y)=d(G(x_{0}),gy).$
Therefore, $(gy,x_{0})\in\Gamma_{u}$ which gives $(x_{0},gy)\in\Gamma^{-}_{u}$, so that $gy\in\Gamma^{-}_{u}(x_{0})$. Moreover, $d(x_{0},gy)=d(x_{0},y)<\delta$. Finally, observe that all the sets $\mathcal{T}_{u}$, $A_{+}$ and $A_{-}$ are $G$-invariant. Thus $\mathcal{T}^{nb}_{u}$ is $G$-invariant, and therefore, $gy\in S_{x_{0}}$.
Point (ii). We begin by observing that under $R_{u}$, the equivalence class of $y\in S_{x_{0}}$ is the unique geodesic ray $\gamma_{y}$ starting at $x_{0}$ realizing the distance $d(G(x_{0}),y)=d(x_{0},y)$, i.e. realizing the distance from $y$ to $G(x_{0})$.
Now assume that for some $g\in G$ we have $gS_{x_{0}}\cap S_{x_{0}}\neq\emptyset$. That is, there exist $x,y\in S_{x_{0}}$ such that $gx=y$. Then we let $\gamma_{x}$ and $\gamma_{y}$ be the unique geodesic segments starting at $x_{0}$, with end points $x$ and $y$, and realizing the distances $d(G(x_{0}),x)=d(x_{0},x)$ and $d(G(x_{0}),y)=d(x_{0},y)$ respectively. Observe that $g\gamma_{x}$ is also a geodesic segment starting at $gx_{0}$ with endpoint $gx=y$ and realizing the distance $d(gx_{0},y)=d(gx_{0},gx)=d(G(x_{0}),gx)=d(G(x_{0}),y)=d(x_{0},y)$. Thus, as the equivalence class of $y$ in $\mathcal{T}^{nb}_{u}$ under $R_{u}$ is the unique geodesic joining $y$ to $G(x_{0})$, realizing the distance $d(G(x_{0}),y)$, this implies that $g\gamma_{x}=\gamma_{y}$ and in particular has to also start at $x_{0}$. Whence $gx_{0}=x_{0}$, and thus $g\in G_{x_{0}}$.
Point (iii). Observe that we have a principal bundle $\rho\colon G\to G/G_{x_{0}}$, where we consider the right action of $G_{x_{0}}$ on $G$. Set $H=G_{x_{0}}$. This implies the existence of an open neighborhood $U\subset G/H$ containing $eH$, such that $\rho^{-1}(U)$ is equivarantly homeomorphic to $H\times U$. This further implies the existence of a continuous cross-section $\chi\colon U\to G$. We now define a map $F\colon U\times S_{x_{0}}\to X$ by $F(gH,s)=\chi(gH)s$. We claim that $F$ is a homeomorphism onto its image.
To prove the injectivity of $F$, consider $(g_{1}H,s_{1})$, $(g_{2}H,s_{2})\in U\times S_{x_{0}}$, such that $\chi(g_{1}H)s_{1}=\chi(g_{2}H)s_{2}$. This implies $\chi(g_{2}H)^{-1}\chi(g_{1}H)s_{1}=s_{2}\in S_{x_{0}}$. By Point (ii), this implies that $\chi(g_{2}H)^{-1}\chi(g_{1}H)\in H$, and thus for some $g\in H$, we have $g\chi(g_{2}H)=\chi(g_{1}H)$. Thus $g_{2}H=\rho(g\chi(g_{2}H))=\rho(\chi(g_{1}H))=g_{1}H$. This last part implies that $s_{1}=s_{2}$. Therefore $F$ is injective.
We now show that $F$ is onto and open. Take $V\subset U$ open, and consider the homeomorphism $\tilde{\alpha}_{x_{0}}\colon G/H\to G(x_{0})$ guaranteed by Theorem 2.26.
Consider $y\in B_{\delta}(\tilde{\alpha}_{x_{0}}(V))\subset X$. Then there exists an $gH\in V$ such that for $gx_{0}\in G(x_{0})$ we get $\delta>d(G(x_{0}),y)=d(gx_{0},y)$. Observe that we have $\chi(gH)x_{0}=\alpha_{x_{0}}(\chi(gH))=\tilde{\alpha}_{x_{0}}(\rho(\chi(gH)))=\tilde{\alpha}_{x_{0}}(gH)=gx_{0}$. Denote by $\gamma_{y}$ the minimizing geodesic realizing the distance between $gx_{0}$ and $y$. Then the geodesic $\chi(gH)^{-1}\gamma_{y}$ joining $\chi(gH)^{-1}y$ to $x_{0}$ has length $d(\chi(gH)^{-1}y,x_{0})=d(y,gx_{0})=d(G(x_{0}),y)=d(G(x_{0}),\chi(gH)^{-1}y)$. Thus $\chi(gH)^{-1}y$ is a point in $\Gamma_{u}(x_{0})$ and in $B_{\delta}(x_{0})$. Therefore $\chi(gH)^{-1}y$ is a point in $S_{x_{0}}$. Thus we have that $(gH,\chi(gH)^{-1}y)\in V\times S_{x_{0}}$ with $F(gH,\chi(gH)^{-1}y)=y$. This implies that $B_{\delta}(\tilde{\alpha}_{x_{0}}(V))\subset F(U\times S_{x_{0}})$. That is $F$ is an open bijective continuous map onto its image, and thus it is an homeomorphism onto its image. ∎
Remark 3.7.
We point out that in the previous proofs we only need to be able to define a metric $d^{\ast}$ on $X^{\ast}$, and that $(X^{\ast},d^{\ast})$ is isometric to $\mathbb{R}$, a closed ray, a closed interval, or $\mathbf{S}^{1}$ with the length metric induced from $\mathbb{R}^{2}$. When $G$ is non-compact, by asking that the action be proper, i.e. the map $\bar{A}\colon G\times X\to X\times X$ is a proper map, is a sufficient condition so that there exists a $G$-invariant metric $d$ on $X$ such that $d^{\ast}$ is well defined. This follows from the fact that the isotropy groups of proper actions are compact, that the orbits are closed, and the orbit space is Hausdorff (the orbits cannot get arbitrarily close to each other).
One can drop the assumption on the dimension of the orbit space in Corollary 3.5 and Theorem C to obtain the same conclusions if the orbit space is isometric to an Alexandrov space, as we now show. Note however that we are still not assuming any curvature bound on $X$.
Theorem 3.8.
Let $G$ act by isometries on $(X,d)$ a metric space, and assume that $(X^{\ast},d^{\ast})$ is isometric to an Alexandrov space. Moreover, assume that there exists $\mathcal{M}\subset X^{\ast}$ an open subset such that for $x_{0}^{\ast}\in\mathcal{M}$ the distance of the cut locus $\mathcal{C}(x_{0}^{\ast})$ to $x_{0}^{\ast}$ is positive. Then the set $\mathcal{T}^{nb}_{u^{\ast}}$ is open, where $u^{\ast}\colon X^{\ast}\to\mathbb{R}$ is defined as $u^{\ast}(x^{\ast})=d(x_{0}^{\ast},x^{\ast})$. Moreover, for any point $x\in\pi^{-1}(x_{0})$ we have a slice.
Proof.
We fix $x_{0}^{\ast}\in\mathcal{M}$, and consider $u^{\ast}\colon X^{\ast}\to\mathbb{R}$ given by $u^{\ast}(y^{\ast})=d^{\ast}(x_{0}^{\ast},y^{\ast})$. We observe that given $y^{\ast}\in A_{\pm}(u^{\ast})$, there exists two distinct geodesics $\gamma_{1}^{\ast}$ and $\gamma_{2}^{\ast}$ of $(X^{\ast},d^{\ast})$ starting at $x_{0}^{\ast}$, passing through $y^{\ast}$, such that there exist points $w^{\ast}\in\gamma_{1}^{\ast}$ and $v^{\ast}\in\gamma_{2}^{\ast}$ for which the shortest geodesic $\gamma_{3}^{\ast}$ between them does not contain $x_{0}^{\ast}$. This implies that $y^{\ast}$ is a critical point of $u^{\ast}$ by [79, p. 378]. Then our hypothesis guarantee that there is a sufficiently small ball $B$ centered at $x_{0}^{\ast}$ such that $B\cap A_{\pm}(u^{\ast})=\varnothing$. From this our conclusion follows.
The existence of the slice is proven in an analogous way to the proof of Theorem C by considering $\delta>0$ small enough so that $\pi(B_{\delta}(G(x_{0})))\subset B^{\ast}$. ∎
Remark 3.9.
The condition asking $x_{0}^{\ast}$ to be isolated from its cut locus $\mathcal{C}(x_{0}^{\ast})$ is a not necessary for the existence of the slice. For example if $D\subset\mathbb{R}^{2}$ is a convex domain, then we consider $X$ to be its double which is an Alexandrov space with curvature $\geq 0$. For $x\in\partial D\subset X$, we have that the cut locus is $\mathcal{C}(x)=\partial D\setminus\{x\}$ (see [71, Section 1, Example 2]). Thus $x$ is not isolated from its cut locus. Considering $G$ any compact Lie group with a left-invariant metric, the space $Y:=G\times X$ with the product metric is an Alexandrov space, whose orbit space is $X$, and for any point $(g,x)\in G\times X$, $S_{(g,x)}:=\{g\}\times X$ is a slice.
3.1. Kleiner’s Lemma
Here we show that a result analogous to Kleiner’s Lemma for Riemannian manifolds [54, Lemma 2.2.2] and Alexandrov spaces [26, Lemma 2.1] holds for cohomogeneity one geodesic spaces:
Lemma 3.10 (Kleiner’s Lemma).
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, $\mathbf{S}^{1}$. Then for a minimal geodesic $\gamma\colon[0,d]\to X$ such that $d(\gamma(0),\gamma(d))=d^{\ast}(\gamma(0)^{\ast},\gamma(d)^{\ast})$, we have that for $t\in(0,d)$ the isotropy group $G_{\gamma(t)}$ is a subgroup of $G_{\gamma(0)}$ and $G_{\gamma(d)}$.
Proof.
We point out that for any $t\in[0,d]$ there exists a subinterval $(t-\varepsilon,t+\varepsilon)\subset[0,d]$ such that the geodesic $\gamma|_{(t-\varepsilon,t+\varepsilon)}$ is contained in the slice $S_{\gamma(t)}$ given by Theorem C. Thus we can apply an analogous argument to the one given in the proof of [26, Lemma 2.1]:
Without loss of generality, assume that there exists $t_{0}\in(0,d)$ such that there exits $g\in G_{\gamma(t_{0})}$ with $g\not\in G_{\gamma(d)}$. The case of assuming that $g\not\in G_{\gamma(0)}$ is analogous. Then, there exists possibly a new $t_{0}^{\prime}$ and a subinterval $[t_{0}^{\prime}-\delta,t_{0}^{\prime}+\delta]$ such that the geodesic $\gamma|_{[t_{0}^{\prime}-\delta,t_{0}^{\prime}+\delta]}$ is in the slice $S_{\gamma(t_{0}^{\prime})}$ and $g\gamma(t_{0}^{\prime}+\delta)\neq\gamma(t_{0}^{\prime}+\delta)$. But this gives a branching geodesic contained in the slice $S_{\gamma(t_{0}^{\prime})}$ which is a contradiction. ∎
With this we prove that the isotropy group is constant in the interior of the orbit space:
Lemma 3.11 (Principal Isotropy Lemma).
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, $\mathbf{S}^{1}$. All the orbits corresponding to the interior of $X^{\ast}$ have constant isotropy subgroup.
Proof.
Consider $x\in X$, such that $x^{\ast}\in\mathrm{int}(X^{\ast})$. Then there exists $y^{\ast}\in\mathrm{int}(X^{\ast})$, such that $x$ is in the slice $S_{y}$ for some $y\in\pi^{-1}(y^{\ast})$ for $\pi\colon X\to X^{\ast}$. In particular $x$ is in the interior of a minimizing geodesic starting at $y$ realizing the distance between the orbits. Then by Lemma 3.10, it follows that $G_{x}\subset G_{y}$. But we can apply the same idea to $y$, now taking the slice through $x$. Thus we get that $G_{y}=G_{x}$, and this is true for all points in $\pi^{-1}(\mathrm{int}(X^{\ast})$. ∎
3.2. Group Diagram
We now show a relation between the isotropy subgroups for an action of a compact Lie group $G$ on a geodesic space $(X,d)$ by isometries. From this relation we obtain a group diagram for actions of cohomogeneity one by isometries on geodesic spaces.
Lemma 3.12.
Let $G$ be a compact Lie group, acting by isometries on a geodesic space $(X,d)$, such that $X^{\ast}$ is isometric to $[-1,1]$. Given $t_{0}\in(-1,1)$ there exists $\varepsilon>0$ and a geodesic $\gamma\colon(-\varepsilon+t_{0},\varepsilon+t_{0})\to X$ with $\gamma(t_{0})\in\pi^{-1}(t_{0})$ realizing the distance between the orbits. That is, for any $t,s\in(-\varepsilon+t_{0},\varepsilon+t_{0})$ we have $d(G(\gamma(s)),G(\gamma(t)))=d(\gamma(s),\gamma(t))$.
Proof.
We fix $x_{0}\in\pi^{-1}(t_{0})=G(x_{0})$, and choose $\varepsilon>0$ small enough so that for the slice $S_{x_{0}}$ through $x_{0}$ we have $(-\varepsilon+t_{0},\varepsilon+t_{0})\subset\pi(S_{x_{0}})$. We consider the minimal geodesic $\gamma^{\ast}\colon(-\varepsilon+t_{0},\varepsilon+t_{0})\to[-1,1]$ with $\gamma^{\ast}(t_{0})=t_{0}$, and fix $t_{0}-\varepsilon<t_{0}-s\leq t_{0}<t_{0}+t<t_{0}+\varepsilon$. Then there exist points $y,z\in S_{x_{0}}$ such that $y\in S_{x_{0}}\cap\pi^{-1}(t_{0}-s)$ and $z\in S_{x_{0}}\cap\pi^{-1}(t_{0}+t)$. Moreover, by construction of $S_{x_{0}}$, there exist minimizing geodesics $\tilde{\gamma}_{1}\colon[-s,0]\to X$ and $\tilde{\gamma}_{1}\colon[0,t]\to X$ such that $\tilde{\gamma}_{i}(r)\in S_{x_{0}}$ for $r\in[-s,0]$ if $i=1$ and $r\in[0,t]$ if $i=2$. This implies that for the concatenation of $\tilde{\gamma}_{2}$ and $\tilde{\gamma}_{1}$ we have by construction $\pi(\tilde{\gamma}_{2}\ast\tilde{\gamma}_{1})=\gamma^{\ast}|_{[t_{0}-s,t_{0}+t]}$, and moreover $d(\tilde{\gamma}_{1}(-r+t_{0}),x_{0})=r=d^{\ast}(\gamma^{\ast}(t_{0}-r),t_{0})$ for $r\in[0,s]$, and $d(\tilde{\gamma}_{2}(\ell+t_{0}),x_{0})=\ell=d^{\ast}(\gamma^{\ast}(t_{0}+\ell),t_{0})$ for $\ell\in[0,t]$. From this we have for $t_{0}-s\leq t_{0}-r\leq t_{0}<t_{0}+\ell\leq t_{0}+t$:
$d(\tilde{\gamma}_{1}(t_{0}-r),\tilde{\gamma}_{2}(t_{0}+\ell))$ $\leq d(\tilde{\gamma}_{1}(t_{0}-r),x_{0})+d(x_{0},\tilde{\gamma}_{2}(t_{0}+\ell))$
$d^{\ast}(\gamma^{\ast}(t_{0}-r),t_{0})+d^{\ast}(t_{0},\gamma^{\ast}_{2}(t_{0}+\ell))$
$=d^{\ast}(\gamma^{\ast}(t_{0}-r),\gamma^{\ast}_{2}(t_{0}+\ell))$
$=d(G(\tilde{\gamma}_{1}(t_{0}-r)),G(\tilde{\gamma}_{2}(t_{0}+\ell)))$
$\leq d(\tilde{\gamma}_{1}(t_{0}-r),\tilde{\gamma}_{2}(t_{0}+\ell)).$
Thus we conclude that for the concatenation $\gamma:=\tilde{\gamma}_{2}\ast\tilde{\gamma}_{1}\colon[t_{0}-s,t_{0}+t]$ we have for all $r,\ell\in[t_{0}-s,t_{0}+t]$ that
$\mathrm{dist}(\gamma(r),\gamma(\ell))=|r-\ell|.$
Therefore, it follows that $\gamma$ is a geodesic realizing the distance between the orbits. ∎
Lemma 3.13.
Let $G$ be a compact Lie group acting by isometries on a geodesic space $(X,d)$, such that the orbit space $X^{\ast}$ is isometric to $[-1,1]$. Then there exists a minimal geodesic $\gamma\colon[-1,1]\to X$ realizing the distance between the orbits. That is, for any $t,s\in[-1,1]$ we have $d(G(\gamma(s)),G(\gamma(t)))=d(\gamma(s),\gamma(t))$.
Proof.
We fix $\gamma^{\ast}\colon(-1,1)\to(-1,1)$ the minimal geodesic with $\gamma^{\ast}(0)=0$ consider a cover $\{U_{i}\}_{i\in\Lambda}$ of $(-1,1)$ by sets satisfying the hypothesis of Lemma 3.12. Now fix $x_{0}\in\pi^{-1}(0)$, and assume that $0\in U_{i_{0}}$, and $U_{j_{0}}\cap U_{i_{0}}\neq\varnothing$, $U_{k_{0}}\cap U_{i_{0}}\neq\varnothing$. We consider the lift $\gamma_{i_{0}}$ of $\gamma^{\ast}|_{U_{i_{0}}}$ through $x_{0}$ given by Lemma 3.12. We observe that we can extend this geodesic to $U_{j_{0}}$ and $U_{k_{0}}$, since the end points of $\gamma_{i_{0}}$ determine points in $\pi^{-1}(U_{j_{0}})$ and $\pi^{-1}(U_{k_{0}})$, and we can apply again Lemma 3.12. Continuing in this fashion we obtain a geodesic $\gamma\colon(-1,1)\to(X,d)$ which is a lift of $\gamma^{\ast}$. Moreover, by construction for $s,t\in(-1,1)$ we have
$d^{\ast}(\gamma^{\ast}(s),\gamma^{\ast}(t))=d(G(\gamma(s)),G(\gamma(t)))=d(\gamma(s),\gamma(t))$
To finish we observe that by [76, Theorem 1] (see also [7, Theorem 1A], [8]) we can extend $\gamma$ to a continuous curve $\gamma\colon[-1,1]\to(X,d)$. Moreover, by continuity of the quotient map $\pi$, and the distances $d^{\ast}$ and $d$ we have
$d(G(\gamma(-1)),G(\gamma(1)))$ $=d^{\ast}(\gamma^{\ast}(-1),\gamma^{\ast}(1))$
$=d^{\ast}(\pi(\gamma)(-1),\pi(\gamma)(1))$
$=\lim_{\begin{subarray}{c}s\to-1\\ t\to 1\end{subarray}}d^{\ast}(\pi(\gamma)(s),\pi(\gamma)(t))$
$=\lim_{\begin{subarray}{c}s\to-1\\ t\to 1\end{subarray}}d(G(\gamma(s)),G(\gamma(t)))$
$=\lim_{\begin{subarray}{c}s\to-1\\ t\to 1\end{subarray}}d(\gamma(s),\gamma(t))$
$=d(\gamma(-1),\gamma(1)).$
This proves our claim. ∎
With this we can relate the isotropy groups as follows.
Theorem 3.14.
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$. Then we have three isotropy subgroups $H$, $G_{-}$, and $G_{+}$, corresponding to orbits in the interior $(-1,1)$, and the end points $\{-1\}$ and $\{+1\}$ respectively. Moreover, $H$ is a subgroup of both $G_{-}$ and $G_{+}$.
Proof.
We fix $\gamma\colon[-1,1]\to X$ a geodesic realizing the distance between the orbits, guaranteed to exist by Lemma 3.13. Then by Lemma 3.10 the conclusions follow, by taking $H:=G_{\gamma(0)}$, and $G_{\pm}=G_{\gamma(\pm)}$. ∎
The tuple $(G,H,G_{+},G_{-})$ obtained in Theorem 3.14 depends a priori on a choice of a curve. In the following lemma we show how two tuples associated to two different actions relate to each other.
Lemma 3.15.
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$. Let $\gamma_{1}$ and $\gamma_{2}$ be two geodesics realizing the distance between the orbits, and $(G,H^{1},K^{1}_{+},K^{1}_{-})$ and $(G,H^{2},K^{2}_{+},K^{2}_{-})$ their associated tuples. Then for some $g_{0}\in G$ and $a_{\pm}\in N(H^{1})_{0}$ we have
$(G,H^{2},K_{+}^{2},K_{-}^{2})=(G,g_{0}H^{1}g_{0}^{-1},g_{0}a_{+}K_{+}^{2}a_{+}^{1}g_{0}^{-1},g_{0}a_{-}K_{-}^{2}a_{-}^{1}g_{0}^{-1}).$
Here $N(H)_{0}$ denotes the connected component of the identity in the normalizer of $H^{1}$
Proof.
We consider $\{x_{0}^{1}\}=\pi^{-1}(0)\cap\gamma_{1}([-1,1])$ and $\{x_{0}^{1}\}=\pi^{-1}(0)\cap\gamma_{2}([-1,1])$. Then there exists $g_{0}\in G$, such that $gx^{1}_{0}=x_{0}^{2}$. Then $G_{x_{0}^{2}}=g_{0}G_{x^{1}_{0}}g_{0}^{-1}$. Moreover, $\tilde{\gamma}_{2}:=g_{0}^{-1}\gamma_{2}$ is also a geodesic with $\tilde{\gamma}_{2}(0)=x_{0}^{1}$. Moreover, there exists $a\colon[-1,1]\to G$ such that $a(t)\gamma_{1}(t)=\tilde{\gamma}_{2}(t)$. Observe that $a(0)=e$, and thus we have that $G_{\tilde{\gamma}_{2}(t)}=a(t)G_{\gamma_{1}(t)}(a(t))^{-1}$. But by Lemma 3.10 we have that $G_{\gamma_{1}(t)}=\Gamma_{\gamma_{1}(0)}=\Gamma_{x_{0}^{1}}=\Gamma_{\tilde{\gamma}_{2}(0)}=G_{\tilde{\gamma}_{2}(t)}$. Writing $H:=\Gamma_{x_{0}^{1}}$ we conclude that $a(t)\in N(H)$ for all $t\in[-1,1]$. Thus by setting $a_{\pm}:=a(\pm 1)$ we have that
$\tilde{K}_{\pm}^{2}=G_{\tilde{\gamma}_{2}(\pm 1)}=a_{\pm}G_{\gamma_{1}(\pm 1)}a_{\pm}^{-1}=a\pm K_{\pm}^{1}a_{\pm}^{-1}.$
Since $\tilde{\gamma}_{2}(\pm)=g_{0}^{-1}\gamma_{2}(\pm)$, we have
$\tilde{K}_{\pm}^{2}=G_{\tilde{\gamma}_{2}(\pm 1)}=g_{0}^{-1}G_{\gamma_{2}(\pm)}g_{0}=g_{0}^{-1}K_{\pm}^{2}g_{0}.$
Thus we conclude that
$(G,H^{2},K_{+}^{2},K_{-}^{2})=(G,g_{0}H^{1}g_{0}^{-1},g_{0}a_{+}K_{+}^{2}a_{+}^{1}g_{0}^{-1},g_{0}a_{-}K_{-}^{2}a_{-}^{1}g_{0}^{-1}).$
Recall that given two continuous actions of $G$ by homeomorphism on two topological spaces $X_{1}$ and $X_{2}$, the spaces $X_{1}$ and $X_{2}$ are equivariantly homeomorphic if there exists $\phi\colon X_{1}\to X_{2}$ an homeomorphism such that for any $x\in X_{1}$, $y\in X_{2}$ and $g\in G$ we have
$\phi(gx)=g\phi(x),\quad\mbox{and}\quad\phi^{-1}(gy)=g\phi^{-1}(y).$
Theorem 3.16.
Let $G$ be a compact Lie group acting on two geodesic spaces $(X_{1},d_{1})$ and $(X_{2},d_{2})$ by isometries and $X_{1}^{\ast}$ and $X_{2}^{\ast}$ are isometric to $[-1,1]$. Moreover, assume that we have associated the tuples $(G,H^{1},K^{1}_{+},K_{-}^{1})$ and $(G,H^{2},K^{2}_{+},K_{-}^{2})$ respectively. If the spaces are equivariantly homeomorphic then there exists $g_{0}\in G$ and $a_{\pm}\in N(H^{1})_{0}$ such that $H^{2}=g_{0}H^{1}g_{0}^{-1}$ and either $K_{\pm}^{2}=g_{0}a_{\pm}K_{\pm}^{1}a_{\pm}^{-1}g_{0}^{-1}$ or $K_{\pm}^{2}=g_{0}a_{\pm}K_{\mp}^{1}a_{\pm}^{-1}g_{0}^{-1}$.
Proof.
[…]
Theorem 3.17.
Let $(X,d)$ be a geodesic space, and assume that a compact Lie group $G$ acts by isometries in such a way that $X^{\ast}$ is isometric to $[-1,1]$. Moreover, assume that we have a tuple $(G,H,G_{+},G_{-})$ associated to the action (see Section 3.2). Then by considering a new tuple $(G,H_{1},K_{+},K_{-})$ with $H_{1}=g_{0}Hg_{0}$, $K_{\pm}=g_{0}a_{\pm}G_{\pm}a_{\pm}^{-1}g_{0}$ or $K_{\pm}=g_{0}a_{\pm}G_{\mp}a_{\pm}^{-1}g_{0}$ for $g_{0}\in G$ and $a_{\pm}\in N(H)_{0}$, we obtain a space $X_{2}$ which is equivariantly homeomorphic to $X$.
Proof.
We fix the end points $x_{\pm}\in X$ of the geodesic associated to the diagram. Thus we have $G_{\pm}=G_{x_{\pm}}$. The fiber bundles $G\times_{G_{\pm}}S_{x_{\pm}}\to G(x_{\pm})\cong G/G_{\pm}$ are associated to the principal bundles $G_{\pm}\to G\to G/G_{\pm}$.
Without loss of generality we assume that $K_{\pm}=g_{0}a_{\pm}G_{\mp}a_{\pm}^{-1}g_{0}$. Then the principal bundles $G_{\pm}\to G\to G/G_{\pm}$ are equivalent to the principal bundles $K_{\pm}\to G\to G/K_{\pm}$. Thus the associated fiber bunldes $G\times_{K_{\pm}}S_{x_{\pm}}\to\cong G/K_{\pm}$ are equivalent to $G\times_{G_{\pm}}S_{x_{\pm}}\to G/K_{\pm}$. Also recall that $X$ is the union of $G\times_{G_{-}}S_{x_{-}}$ and $G\times_{G_{+}}S_{x_{+}}$ via their boundary. Thus we obtain that $X$ is equivariantly homeomorphic to $Y$ the union of $G\times_{K_{-}}S_{x_{-}}$ and $G\times_{K_{+}}S_{x_{+}}$ via the boundary. By pushing the metric we can find a geodesic in $Y$ realizing the distance between the orbits and with associated tuple $(G,H_{1},K_{+},K_{-})$. ∎
4. Geometry of the slice
In this section we show that for a compact $\mathsf{RCD}$-space $(X,d,\mathfrak{m}_{X})$ with an isometric and measure preserving action of cohomogeneity $1$ by a compact Lie group $G$, the slice $S_{x}$ through $x\in X$ we presented in Section 3 admits an $\mathsf{RCD}$-structure. Moreover in the case when $X$ is non-collapsed, we show that this structure on $S_{x}$ is given by taking the cone over a homogeneous space. We end the section by collecting all the results and combining them with the conclusions in [27] to obtain a topological structural result.
4.1. Infinitesimal action
Here we consider the infinitesimal geometry of $X$ together with the information of the group action. We being by showing that for the directions tangent to the orbit we obtain a line in a tangent cone of $X$.
Lemma 4.1.
Let $(X,d,\mathfrak{m}_{X})$ be an $\mathsf{RCD}$-space such that all tangent cones are metric cones, and $G$ a Lie group acting effectively by measure-preserving isometries. For $x_{0}\in X$ not a fixed point of the action, we consider $K=G_{x_{0}}$ the isotropy group of $x_{0}$, and consider $\mathfrak{g}=\mathfrak{l}\oplus\mathfrak{k}$ a splitting of the Lie algebra of $G$, where $\mathfrak{k}$ is the Lie algebra of $K$. Let $(Y,d_{Y},\mathfrak{m}_{Y},0_{Y})\in\mathrm{Tan}(X,d,\mathfrak{m}_{X},x_{0})$. Then for $v\in\mathfrak{l}$, $v\neq 0$, we have a line $l_{v}$ in $(Y,d_{Y})$ passing through $0_{Y}$.
Proof.
Consider $\{r_{n}\}_{n\in\mathbb{N}}\subset\mathbb{R}$ with $r_{n}\to 0$ as $n\to\infty$, and let $(Y,d_{Y},0_{Y})$ be the pointed-Gromov-Hausdorff limit of the sequence $X_{n}=(X,\frac{1}{r_{n}}d,x_{0})$. By [25] we have a closed limit group $H\subset\mathrm{Isom}(Y,d_{Y})$, such that $(X,\frac{1}{r_{n}}d,G,x_{0})$ converges pointed-equivarianlty-Gromov-Hausdorff to $(Y,d_{Y},H,0_{Y})$
Fix $v\in\mathfrak{l}$, $v\neq 0$, and choose a sequence $\{t_{n}\}\subset\mathbb{R}$, such that the points $\exp(t_{n}v)\cdot x_{0}\in(X,(1/r_{n})d,x_{0})$ has a limit in $Y$, and it is different from $0_{Y}$. Then we have a limit isometry $\gamma_{v}\in H\subset\mathrm{Isom}(Y,d_{Y})$ such that $\gamma_{v}\cdot 0_{Y}\neq 0_{Y}$.
Since by hypothesis $Y$ is a metric cone, then there exists an a ray $l\cong[0,b]$ starting at $0_{Y}$, such that $\gamma_{v}\cdot 0_{Y}$ is contained in the interior of $l$. We apply $\gamma_{v}^{-1}$ to the interior of $l$, to obtain an open ray passing through $0_{Y}$.
Since $(Y,d_{Y})$ is a cone, it is invariant under rescalings. by rescaling by a parameter $\lambda$ going to $\infty$, the segment $\gamma^{-1}_{v}\cdot l$ converges to a line $l_{v}$ in $Y$, passing through $0_{Y}$ and $\gamma_{v}0_{Y}$. ∎
Remark 4.2.
Observe that in the previous Lemma, for $v,w\in\mathfrak{l}$ linearly independent, we can choose the sequence $\{t_{n}\}$ in a way so that for the limit isometries $\gamma_{v}\neq\gamma_{w}$, and thus the lines $l_{v}\neq l_{w}$.
Remark 4.3.
Observe that for $X$ a non-collapsed $\mathsf{RCD}$-space, the tangent cones are metric cones [23, Proposition 2.8].
With the previous lemma, we show that we can split a Euclidean space as metric factor from any tangent cone of $X$.
Lemma 4.4.
Let $(X,d,\mathfrak{m})$ be a $\mathsf{RCD}(K,N)$-space such that all tangent cones are metric cones, and $G$ a Lie group of dimension $m$ acting effectively by measure-preserving isometries. For $x_{0}\in X$, we consider $K=G_{x_{0}}$. Let $(Y,d_{Y},\mathfrak{m}_{Y},0_{Y})\in\mathrm{Tan}(X,d,\mathfrak{m}_{X},x_{0})$. Taking $k=\dim(G(x_{0}))=\mathrm{dim}(G)-\mathrm{dim}(K)$, we have $(Y,d_{Y},\mathfrak{m}_{Y},0_{Y})$ is isometric to $\mathbb{R}^{k}\times(Z,d_{Z},\mathfrak{m}_{Z},0_{Z})$, where $Z$ is an $\mathsf{RCD}(0,N-k)$-space.
Proof.
Observe that $(Y,d_{Y},\mathfrak{m}_{Y})$ is an $\mathsf{RCD}(0,N)$-space. We consider $\mathfrak{l}$ a complement of the Lie algebra of $K$, i.e. $\mathfrak{g}=\mathfrak{l}\oplus\mathfrak{k}$. Then for each element in a basis of $\mathfrak{l}$, by Lemma 4.1 we get a different line in $Y$. Thus by Gigli’s Splitting Theorem [31] the conclusion follows. ∎
Remark 4.5.
The condition of $X$ having all tangent cones to be metric cones is a necessary one: In the collapsed case, Pan and Wei [66] have a non-compact example of a collapsed $\mathsf{RCD}$-space (thus it cannot be a metric cone) with an $\mathbb{R}$-action of cohomogeneity one by measure preserving isometries, which is its own tangent cone, and such that no line can be split.
In the following Lemma we extend results in [43], [68] about the existence of a nice group action on the limit of a convergent sequence of non-collapsed $\mathsf{RCD}$-spaces equipped with isometric measure preserving group actions.
Lemma 4.6.
Let $G$ act on $(X_{n},d_{n},\mathfrak{m}_{n})$ effectively by measure preserving isometries, where $(X_{n},d_{n},\mathfrak{m}_{n})$ are $\mathsf{RCD}(K,N)$-spaces. Let $(Y,d_{Y},\mathfrak{m}_{Y})$ be a metric measure space such that
$(X_{n},d_{n},\mathfrak{m}_{n},x_{n})\to(Y,d_{Y},\mathfrak{m}_{Y},0_{Y})$
in the pointed measured Gromov-Hausdorff topology. Then there exists a Lie group $G_{Y}$ acting effectively by measure preserving isometries on $Y$, such that for large enough $n$ there exist Lie group homomorphisms $\phi_{n}\colon G\to G_{Y}$ which demonstrate the equivariant pointed Gromov-Hausdorff convergence. Furthermore, if the $G$-action is both Lipschitz and co-Lipschitz continuous, and the $X_{n}$ are non-collapsed $\mathsf{RCD}$-spaces or the base points $x_{n}$ are regular points then the morphisms $\phi_{n}$ are injective.
Proof.
In [68, Theorem 5.6] we have the proof of this statement for the case of pointed Gromov Hausdorff convergence when $(X,d,\mathfrak{m})$ is an $\mathsf{RCD}$-space, $X_{n}:=X$, $d_{n}:=r_{n}d$ for a sequence $r_{n}\underset{n\to\infty}{\longrightarrow}0$, and $x_{n}:=x$ is a regular point. Hence we only consider the remaining case here.
We recall that since $(X_{n},d_{n},x_{n})$ converges to $(Y,d_{Y},0_{Y})$ with respect to the pointed Gromov-Hausdorff topology then [25, Proposition 3.6] guarantees the existence of $G_{Y}$ such that $(X_{n},d_{n},G,x_{n})$ converge to $(Y,d_{Y},G_{Y},0_{Y})$ with respect to the equivariant pointed Gromov-Hausdorff topology. In this setting, [43, Theorem 3.1] guarantees the existence of the required Lie group homomorphisms $\phi_{n}\colon G\to G_{Y}$.
Now assume that the spaces $X_{n}$ are non-collapsed (and thus that their Hausdorff dimension is $N$). Then by [11, Theorem 2.18] it follows that $Y$ is non-collapsed, and thus has Hausdorff dimension $N$. Let $K_{n}=Ker(\phi_{n})\subset G$. Then [27, Theorem 1.1] implies that $X_{n}/K_{n}$ are $\mathsf{RCD}(K,N)$-spaces. Moreover the hypothesis of [27, Theorem 6.3] are satisfied. Thus we have we have that for $k_{n}=\mathrm{dim}_{\mathrm{ess}}(X_{n}/K_{n})$ and $N=\mathrm{dim}_{\mathrm{ess}}(X_{n})$ the following relation holds
$N=k_{n}+\mathrm{dim}(K_{n}/H_{n}),$
where $H_{n}$ is the principal isotropy of the $K_{n}$-action. Since the sequence $(X_{n},d_{n},K_{n})$ converges to $(Y,d_{Y},\{e\})$ in the equivariant Gromov-Hausdorff topology by construction, we have that $(X_{n}/K_{n},d^{\ast}_{n},(\mathfrak{m}_{n}^{r_{n}})^{\ast},x_{0}^{\ast})$ converges to $(Y,d_{Y},\mathfrak{m}_{Y},0_{Y})$ in the pointed-measured Gromov-Hausdorff topology. Since the essential dimension is lower semicontinuous we have
$N=\mathrm{dim}_{\mathrm{ess}}(Y)\leq\liminf_{n\to\infty}\mathrm{dim}_{\mathrm{ess}}(X_{n}/K_{n})=k_{n}\leq N.$
From this we conclude that for large enough $n$, the $K_{n}$ are discrete and thus $\phi_{n}$ is injective.
With this we can show the existence of an “infinitesimal action” on the tangent cones of a non-collapsed $\mathsf{RCD}$-space
Theorem G.
Let $(X,d,\mathfrak{m}_{X})$ be an $\mathsf{RCD}$-space such that all tangent cones are metric cones, $G$ be a Lie group of dimension $m$ acting effectively by measure preserving isometries, and both Lipschitz and co-Lipschitz continuously. For $x_{0}\in X$ fixed, we consider $K=G_{x_{0}}$ and let $k=\dim(K)$. Then each $(Y,d_{Y},\mathfrak{m}_{Y},0_{Y})\in\mathrm{Tan}(X,d,\mathfrak{m}_{X},x_{0})$ is isometric to $\mathbb{R}^{m-k}\times(Z,d_{Z},\mathfrak{m}_{Z},0_{Z})$, for some space $Z$ which is an $\mathsf{RCD}(0,N-m+k)$-space. Moreover $K$ acts effectively by measure preserving isometries on $(Z,d_{Z},\mathfrak{m}_{Z})$, and fixes $0_{Z}$.
Proof.
Observe that $K$ leaves the orbit $G(x_{0})$ invariant and acts trivially on it by construction. Moreover for the sequence of rescalings $X_{n}$ of $X$ approximating $Y$ we have an action of $K$ by measure preserving isometries. Thus by Lemma 4.6, since $X_{n}\to Y$ in the pm-Gromov-Hausdorff topology, there exists $G_{Y}$ a Lie group acting by measure preserving isometries on $Y$, and for sufficiently large $n$, Lie group homomorphisms $\phi_{n}\colon K\to G_{Y}$ which are injective and demonstrate the convergence.
Let $\mathfrak{l}$ be a complement of $\mathfrak{k}$ the Lie algebra of $K$, i.e. $\mathfrak{g}=\mathfrak{l}\oplus\mathfrak{k}$. Then for $v\in\mathfrak{l}$, by Lemma 4.1, the curve $\exp(tv)\cdot x_{0}$ goes to a line in $Y$. This implies that for each $n$, the image of $\phi_{n}(K)$ leaves invariant the line corresponding to $v$, and acts trivially on it. Also observe that since $K$ fixes $x_{0}$, then $\phi_{n}(K)$ fixes $0_{Y}$.
Since $Y$ isomorphic to $\mathbb{R}^{m-k}\times Z$, we conclude that we have an $\phi_{n}(K)$-action on $\mathbb{R}^{m-k}\times Z$ by measure preserving isometries. By the above observation $\phi_{n}(K)$ acts trivially on $\mathbb{R}^{m-k}$. Thus $\phi_{n}(K)$ acts effectively by measure preserving isometries on $Z$. Since $\phi_{n}(K)$ fixes $0_{Y}$, then $\phi_{n}(K)$ fixes $0_{Z}$.
Last, since $\phi_{n}$ is injective, we have that $\phi_{n}(K)$ is isomorphic as a Lie group to $K$. ∎
Observe that in Lemma 4.6 when $x_{0}$ is regular we obtain actions of $G_{x_{0}}$ on the tangent cone $(Y,d_{Y},0_{y})\cong(\mathbb{R}^{n},d_{E},0)$ of $X$ at $x_{0}$. Then by the same proof of Theorem G we obtain the following theorem.
Theorem H (Principal orbit representation).
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}$-space of essential dimension $n$, $G$ be a Lie group of dimension $m$ acting effectively by measure preserving isometries, and both Lipschitz and co-Lipschitz continuously. For $x_{0}\in X$ a fixed regular point, we consider $K=G_{x_{0}}$. Then the action of $G$ induces a linear representation of $K$ into $\mathrm{O}(n^{\ast})$, where $n^{\ast}$ is the essential dimension of $X^{\ast}$.
Remark 4.7.
In the case when $X$ is either a Riemannian manifold, or an Alexandrov space, the isotropy action of the principal isotropy group in Theorem H on the slice is trivial. In the case of $\mathsf{RCD}$-spaces we are unsure if also the isotropy action of the isotropy group on the slice is trivial.
4.2. Geometry of the slice
In this subsection we show that the slice $S_{x}$ at $x$ can be identified with a subset of any tangent cone, which admits an $\mathsf{RCD}(0,N-k)$ structure, where $k$ is the dimension of the orbit through $x$.
Lemma 4.8.
Let $(X,d,\mathfrak{m}_{X})$ be a $\mathsf{RCD}(K,N)$-space such that all tangent cones are metric cones. Assume that $G$ is a compact Lie group acting effectively by measure preserving isometries on $X$, such that $(X^{\ast},d^{\ast},\mathfrak{m}_{X}^{\ast})$ has essential dimension equal to $1$. For $x_{0}\in X$ fixed we set $K=G_{x_{0}}$, $k=\mathrm{dim}(G(x_{0}))$, and consider $(Y,d_{Y},0_{Y})\cong(\mathbb{R}^{k},d_{\mathbb{E}},0)\times(Z,d_{Z},0_{Z})\in\mathrm{Tan}(X,d,x_{0})$. Then $Z$ is homeomorphic to $S_{x_{0}}$, and both are homeomorphic to an open cone over a homogeneous space.
Proof.
Since $(X^{\ast},d^{\ast},\mathfrak{m}^{\ast})$ is an $\mathsf{RCD}$-space by [27] of essential dimension equal to $1$, then by [53] we have that $(X^{\ast},d^{\ast})$ is isometric to $(\mathbb{R},d_{\mathbb{E}}$, $([0,\infty),d_{\mathbb{E}})$, $([0,1],d_{\mathbb{E}})$, or $(\mathbf{S}^{1}(1),d)$ where $d$ is the length metric induced by $\mathbb{R}^{2}$. Then a slice $S_{x_{0}}$ exists by Theorem C.
We consider $\{r_{n}\}$ the sequence such that $(X,\frac{1}{r_{n}}d,x_{0})$ converges to $(Y,d_{Y},0_{Y})$, and observe that for each $n\in\mathbb{N}$, the closed Lie group $K$ acts by isometries. Thus by [25, Proposition 3.6] implies that there exists a closed Lie group $\overline{K}$ acting by isometries on $(Y,d_{Y})$.
We take $U\subset G(x_{0})$ sufficiently small open neighborhood of $x_{0}$, and $S_{x_{0}}$ the slice at $x_{0}$. Then by construction, we have an homeomorphism $F\colon U\times S_{x_{0}}\to X$ onto an open neighborhood $W$ of $x_{0}$ in $X$. We observe that for all $n\in\mathbb{N}$ sufficiently large we can assume that the open balls $B_{r_{n}}(x_{0})$ are contained in $W$. From this it follows that $(Y,d_{Y},0_{Y})$ is also a tangent cone of $U\times S_{x_{0}}$ with the restricted metric.
Now we consider $\mathfrak{k}\subset\mathfrak{g}$ the Lie subalgebra of $K$, and $\mathfrak{l}$ a Lie subalgebra such that $\mathfrak{g}=\mathfrak{l}\oplus\mathfrak{k}$. Then, as in the proof of Theorem 4.1, we can see that for each direction in $\mathfrak{l}$ we obtain a line in $(Y,d_{Y})$.
Moreover, the action of $K$ on $U$ is trivial by construction. Thus we have that $(U\times S_{x_{0}})/K=U\times(S_{x_{0}}/K)$. Since $(X,\frac{1}{r_{n}}d,K,x_{0})$ converge with respect to the pointed-equivariant Gromov-Hausdorff topology to $(Y,d_{Y},\overline{K},0_{Y})$, we have that $\Big{(}(U\times S_{x_{0}})/K,\frac{1}{r_{n}}d^{\ast},x_{0}^{\ast})\Big{)}$ converges with respect to the pointed Gromov-Hausdorff topology to $(Y/\overline{K},d_{Y}^{\ast},0_{Y}^{\ast})$. Thus, $(Y/\overline{K},d_{Y}^{\ast},0_{Y}^{\ast})$ is a tangent cone of $\Big{(}(U\times S_{x_{0}})/K,d^{\ast},x_{0}^{\ast})\Big{)}$. But we have that for each tangent direction to $U$, in the limit of the blow up we obtain a line. Thus as in Lemma 4.4, we conclude that $(Y/\overline{K},d_{Y}^{\ast})$ is isometric to $(\mathbb{R}^{k},d_{\mathbb{E}},0)\times(Z,d_{Z},0_{Z})$ for some space $Z$ that is an $\mathsf{RCD}(0,N-k)$-space.
Now we observe that by construction, setting $\delta>0$ small enough and considering the open tubular neighborhood $\operatorname{Tub}^{\delta}(G(x_{0}))\subset X$ of the orbit $G(x_{0})$ we have that $(\operatorname{Tub}^{\delta}(G(x_{0}))/K,d^{\ast})$ is isometric to $([0,\delta),d_{\mathbb{E}})$. But also $((S_{x_{0}}\cap B_{\delta}(x_{0}))/K,d^{\ast})$ is isometric to $([0,\delta),d_{\mathbb{E}})$, since $S_{x_{0}}$ consists of the union of geodesics realizing the distance between orbits to the orbit $G(x_{0})$. Thus we conclude that $(Y/\overline{K},d_{Y}^{\ast})$ is isometric to $(\mathbb{R}^{k},d_{\mathbb{E}})\times([0,\infty),d_{\mathbb{E}})$. Thus $Z/\overline{K}$ is homeomorphic to a ray $[0,\infty)$, i.e the action of $\overline{K}$ on $Z$ is of cohomogeneity one.
Now we consider $\gamma\colon[0,1]\to S_{x_{0}}\subset X$ a minimizing geodesic of $X$, starting at $x_{0}$, and realizing the distance between $G(x_{0})$ and $G(\gamma(1))$. Then the ray $(\gamma([0,1)),\frac{1}{r_{n}}d)$ converges to a minimizing ray $\widetilde{\gamma}\colon[0,\infty)\to(Y,d_{Y})$. Observe that under the submetry $\pi_{X}\colon(X,d)\to(X/K,d^{\ast})$, we have that $\pi_{X}(\gamma)$ is a minimizing geodesic between any of its points. Thus we conclude that for the submetry $\pi_{Y}\colon(Y,d_{Y})\to(Y/\overline{K},d^{\ast}_{Y})$, the curve $\pi_{Y}(\widetilde{\gamma})$ is a minimizing geodesic between its endpoints. Moreover, $\pi_{Y}(\widetilde{\gamma})$ is the limit of the curves $\pi_{X}(\gamma)$ under the blow up. Thus we conclude that $\pi_{Y}(\widetilde{\gamma})$ is contained in $Z/\overline{K}$.
From this it follows that the $\widetilde{\gamma}$ is a geodesic in $Z$, and realizes the distance between the $\overline{K}$-orbit of the vertex $0_{Z}$ and the $\overline{K}$-orbits. Thus for each radial geodesic $\gamma$ in $S_{x_{0}}$ starting at $x_{0}$, we obtain a limit radial geodesic $\widetilde{\gamma}$ in $Z$ starting at $0_{Z}$. Since we can uniquely identify each radial geodesic in $S_{x_{0}}$ and $Z$, starting at $x_{0}$ and $0_{Z}$ respectively, with their unique intersection point with the boundary of a closed ball $\overline{B}_{\varepsilon}(x_{0})$ and $\overline{B}_{\varepsilon}(0_{Z})$, then we have an homeomorphism between $\overline{B}_{\varepsilon}(x_{0})$ and $\overline{B}_{\varepsilon}(0_{Z})$.
Last we observe that $K$ acts transitively on $\partial\overline{B}_{\varepsilon}(x_{0})$, and $\overline{K}$ acts transitively on $\partial\overline{B}_{\varepsilon}(0_{Z})$. This is due to the fact that $\pi_{X}(\overline{B}_{\varepsilon}(x_{0}))$ is the boundary of a closed ball around $x_{0}^{\ast}$ in $(S_{x_{0}},d^{\ast})$; but such a ball is homeomorphic to $[0,\varepsilon]$. Thus $\pi_{X}(\overline{B}_{\varepsilon}(x_{0}))$ consists of only one point, and this implies that the a $K$-orbit is equal to $\partial\overline{B}_{\varepsilon}(x_{0})$, i.e. the action of $K$ on $\partial\overline{B}_{\varepsilon}(x_{0})$ is transitive. An analogous reasoning holds for $Z$, $\overline{K}$, and $\partial\overline{B}_{\varepsilon}(0_{Z})$. Moreover, $S_{x_{0}}$ is homeomorphic to the topological cone over $\partial\overline{B}_{\varepsilon}(x_{0})$, and analogously $Z$ is homeomorphic to the topological cone over $\partial\overline{B}_{\varepsilon}(0_{Z})$. From this the claim follows. ∎
Remark 4.9.
Consider $(X,d,\mathfrak{m}_{X})$ a $\mathsf{RCD}(K,N)$-space such that all tangent cones are metric cones and $G$ a compact Lie group acting effectively by measure preserving isometries on $X$, such that $(X^{\ast},d^{\ast},\mathfrak{m}_{X}^{\ast})$ has essential dimension equal to $1$, as in Theorem 4.8. Then for $x_{0}\in X$ fixed and $\delta>0$ small enough, if we take $K=G_{x_{0}}$ and $H=G_{y}$ for $y\in\operatorname{Tub}^{\delta}(G_{x_{0}})$, then we have that $S_{x_{0}}$ is homeomorphic to the cone over $K/H$. This is due to the conclusions of Theorem 4.8 the fact that $H$ is the largest subgroup in $K$ that fixes the boundary of $\operatorname{Tub}^{\delta}(G(x_{0}))$, and thus it fixes $\partial\overline{B}_{\delta}(x_{0})\subset S_{x_{0}}$.
Theorem D.
Let $(X,d,\mathfrak{m}_{X})$ be an $\mathsf{RCD}(K,N)$-space with $N>2$, and such that all tangent cones are metric cones. Assume that $G$ is a compact Lie group acting effectively by measure preserving isometries on $X$, such that $(X^{\ast},d^{\ast},\mathfrak{m}_{X}^{\ast})$ has essential dimension equal to $1$. Then the following hold:
  1. (i)
    For any $x_{0}\in X$ the slice $S_{x_{0}}$ admits an $\mathsf{RCD}(0,N-k)$ structure, where $k=\dim(G(x_{0}))$.
  2. (ii)
    Moreover, $S_{x_{0}}$ is homeomorphic to a metric cone over a homogeneous smooth Riemannian manifold with Ricci curvature greater or equal to $N-k-2$.
Proof.
From our hypothesis it follows from Lemma 4.4, that given $x_{0}\in X$ and a fixed tangent cone $(Y,d_{Y},\mathfrak{m}_{Y})\in\mathrm{Tan}(X,d,\mathfrak{m}_{X},x_{0})$ then $Y$ is isometric to $(\mathbb{R}^{k},d_{\mathbb{E}},\mathcal{L}^{k})\times(Z,d_{Z},\mathfrak{m}_{Z})$, where $(Z,d_{Z},\mathfrak{m}_{Z})$ is an $\mathsf{RCD}(0,N-k)$-space. From Lemma 4.8 it follows that $S_{x_{0}}$ exists and is homeomorphic to $Z$, and the first conclusion follows.
Since the tangent cones are metric cones, it follows that $Z$ has to be locally isometric to a metric cone with vertex $0_{Z}$ over an $\mathsf{RCD}(N-k-2,N-k-1)$-space $(F,d_{F},\mathfrak{m}_{F})$ by [52, Theorem 1.4].
Moreover, there is a closed Lie group $\overline{K}$ acting on $(Z,d_{Z},\mathfrak{m}_{Z})$ by measure preserving isometries. As seen in the proof of Lemma 4.8, the $\overline{K}$-orbits are the boundary of closed balls centered at $0_{Z}$. This implies that the boundary of closed balls centered at $0_{Z}$ are homogeneous spaces $\overline{K}/\overline{H}$. This fact, together with the local isometry with the cone over $(F,d_{F},\mathfrak{m}_{F})$, imply that $F$ is homeomorphic to the homogeneous manifold $\overline{K}/\overline{H}$, and that $\overline{K}$ acts by measure preserving isometries over $(F,d_{F},\mathfrak{m}_{F})$. By [69, Proposition 5.14] it follows that $(F,d_{F},\mathfrak{m}_{F})$ is homeomorphic to a homogeneous smooth Riemannian manifold $(M,g_{M})$. Moreover, we have that $(M,d_{g_{M}},\mathfrak{m}_{g_{M}})$ equipped with the Riemannian distance and volume is also an $\mathsf{RCD}(N-k-2,N-k-1)$-space. By [73, Theorem 2 (i)] it follows that $\operatorname{Ric}(g_{M})\geq N-k-2$ as claimed. ∎
Remark 4.10.
For a non collapsed space $(X,d,\mathfrak{m})$ satisfying the hypothesis of Theorem D, by Remark 4.9 we have an action of $G_{x_{0}}$ on the base of the cone $F:=\overline{K}/\overline{H}$ the $\mathsf{RCD}(N-k-2,N-k-1)$-space by homeomorphisms, and for $x_{0}$ not in an orbit of principal type, the space $F$ is homeomorphic to $G_{x_{0}}/H$. However, it is not known a priory if the the transitive action of $G_{x_{0}}$ is by measure preserving isometries with respect to the $\mathsf{RCD}(N-k-2,N-k-1)$-structure.
4.3. Topological Rigidity
Combining Theorem C with Theorem 2.29 and Theorem D we obtain the following corollary:
Theorem A.
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space. Let $G$ be a compact Lie group acting on $X$ by measure preserving isometries and cohomogeneity one. Then the following hold:
  1. (a)
    When $X^{\ast}$ is homeomorphic to $[0,1]$, then $X$ is homeomorphic to the union of two cones bundles.
  2. (b)
    When $X^{\ast}$ is homeomorphic to $[0,\infty)$, then $X$ is homeomorphic to a cone bundle.
  3. (c)
    When $X^{\ast}$ is homeomorphic to $\mathbf{S}^{1}$, then $X$ is the total space of a fiber bundle with homogeneous fiber $G/H$ and structure group $N_{G}(H)/H$.
  4. (d)
    When $X^{\ast}$ is homeomorphic to $\mathbb{R}$, then $X$ is homeomorphic to ×$\mathbb{R}\times G/H$.
Moreover the cone fibers in items $(a)$ and $(b)$ are cones over homogeneous spaces.
If the tangent cones of $X$ at some points in the singular orbits $G(x_{+})$, $G(x_{-})$ have a metric cone structure, then the cone fibers are cones over $\mathsf{RCD}(N-k_{\pm}-2,N-k_{\pm}-1)$-spaces, where $k_{\pm}=\dim(G(x_{\pm}))$.
This completes the structural topological results in [27, Corollary 1.4]
Remark 4.11.
Assume we are in the case that $X^{\ast}$ is closed interval. Let us denote by $G_{\pm}$ the isotropy groups of the singular orbits corresponding to the end point of the interval, and by $H$ the principal isotropy corresponding to the orbits in the interior of the interval. We point out that in the case of Theorem A (a) together with Sections 3.1 and 3.2, the space $X$ and the group action are determined by a group diagram:
[…]
together with two $\mathsf{RCD}$-spaces $S_{\pm}$, that have a cohomogeneity one action by $G_{\pm}$ by measure preserving isometries. Moreover, in the case when $X$ is non-collapsed, $S_{\pm}$ is the cone over a homogeneous space $Z_{\pm}=\bar{G}_{\pm}/\bar{H}_{\pm}$ with effective an transitive action of $G_{\pm}$, respectively. This is in contrast to context of Riemannian manifolds and Alexandrov spaces, where the homogeneous spaces $Z_{\pm}$ are determined by the subgroups $G_{\pm}$ and $H$. Namely for Riemannian manifolds and Alexandrov spaces we have $S_{\pm}=G_{\pm}/H$.
We point out that the topological description in Theorem A does not depend on the measure, only on the metric structure:
Theorem 4.12.
Let $(X,d)$ be a geodesic space, and $G$ compact acting by isometries on $X$. Assume that the orbit space $(X^{\ast},d^{\ast})$ is to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then:
  1. (a)
    When $X^{\ast}$ is homeomorphic to $[0,1]$, then $X$ is homeomorphic to the union of two cones bundles.
  2. (b)
    When $X^{\ast}$ is homeomorphic to $[0,\infty)$, then $X$ is homeomorphic to a cone bundle
  3. (c)
    When $X^{\ast}$ is homeomorphic to $\mathbf{S}^{1}$, then $X$ is the total space of a fiber bundle with homogeneous fiber $G/H$, and structure group $N_{G}(H)/H$.
  4. (d)
    When $X^{\ast}$ is homeomorphic to $\mathbb{R}$, then $X$ is homeomorphic to ×$\mathbb{R}\times G/H$.
Proof.
We begin by proving (c), and (d). By Lemma 3.11 it follows that for any point $x\in X$, the isotropy $G_{x}$ equals a fix subgroup $H<G$. Moreover, by Theorem 2.26 we conclude that for any $x\in X$, the orbit through $x$ is equivariantly homeomorphic to $G/H$, i.e. $G(x)\cong G/H$. Then, as pointed out in the proof of [27, Corollary 4.9], any lift $\tilde{\gamma^{\ast}}\colon I\to X$ of any short enough geodesic $\gamma^{\ast}\colon I\to\mathbf{S}^{1}=X^{\ast}$ gives a trivalization. In the case (c), since any two points in $X$ have conjugated isotropy subgroups, we get the conclusion on the structure group of the bundle. The case (d) follows from the fact that $\mathbb{R}$ is contractible.
The cases (a) and (b) follow from Remark 3.1, Theorem 2.29, and the observation that the spaces $\pi^{-1}([-1,0])$ and $\pi^{-1}([0,1])$ are equivariantly homeomorphic to the cone bundles $G\times_{G_{-}}\operatorname{Cl}(S_{-})$ and $G\times_{G_{+}}\operatorname{Cl}(S_{+})$ respectively. Here $\pi\colon X\to X^{\ast}$ is the orbit quotient map, $G_{\pm}$ are the isotropy groups of the orbits $\pi^{-1}(\{\pm 1\})$, and $S_{\pm}$ are slices through points in $\pi^{-1}(\{\pm 1\})$. ∎
Remark 4.13.
Observe that Theorem 4.12 is valid when $G$ is a non-compact group acting properly, and such that the orbit space is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$.
With Theorem 4.12 we can compute the fundamental group of the geodesic spaces of cohomogeneity one, following [45].
Theorem 4.14.
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $X^{\ast}$ is isometric to $[-1,1]$. We also consider a tuple $(G,H,G_{+},G_{-})$ given by Theorem 3.14 and assume that the orbits $K_{\pm}/H$ are connected. Then the fundamental group of $X$ is isomorphic to
$\big{(}(\pi_{1}(G/H)/N_{+})\ast(\pi_{1}(G/H)/N_{-})\big{)}/N(\mathrm{span}\{(\rho_{+})_{\ast}(\omega)(\rho_{+})_{\ast}(\omega)^{-1}\mid\omega\in\pi_{1}(G/H)\}),$
where $N_{\pm}=\ker\{(\rho_{\pm})_{\ast}\colon\pi_{1}(G/H)\to\pi_{1}(G/H)\}$ for the orbit projection maps $\rho_{\pm}\colon G/H\to G/G_{\pm}$ give by $\rho_{\pm}(gH):=gK_{\pm}$.
Proof.
We recall that the group tuple $(G,H,G_{+},G_{-})$ is associated to a geodesic $\gamma\colon[-1,1]\to X$ realizing the distance between the orbits. Set $x_{\pm}:=\gamma(\pm 1)$ and $x_{0}=\gamma(0)$. Then $x_{\pm}\in\pi^{-1}(\pm 1)$ and $x_{0}\in\pi^{-1}(0)$.
We consider $U_{+}^{\ast},U_{-}^{\ast}\subset[-1,1]$ with $U_{+}^{\ast}\cap U_{-}^{\ast}=(-\varepsilon,\varepsilon)$, and apply van Kampen’s Theorem to the decomposition of $X$ by the open subsets $U_{-}:=\pi^{-1}(U_{-}^{\ast})$, $U_{+}:=\pi^{-1}(U_{+}^{\ast})$, and base point $x_{0}$.
We point out that the slices $S_{x_{\pm}}$ are contractible, and thus the tubular neighborhoods $\mathrm{Tub}(\pi^{-1}(\pm 1)\cong G\times_{G_{\pm}}S_{x_{\pm}}$ onto the orbits $G(x_{\pm})$. Moreover, the point $x_{0}$ goes to the points $x_{\pm}$ under this deformation retract. Thus we have that $U_{\pm}\simeq\mathrm{Tub}(\pi^{-1}(\pm 1)\simeq\pi^{-1}(\pm 1)\cong G/K_{\pm}$, and $U_{-}\cap U_{+}\simeq\pi^{-1}(0)\cong G/H$. Under this homotopy equivalences we have that $x_{0}$ is mapped to the classes $K_{\pm}\in G/K_{\pm}$ and $H\in G/H$. Thus we have the following commutative diagram:
[…]
where $j_{\pm}\colon(U_{+}\cap U_{-},x_{0})\hookrightarrow(U_{\pm},x_{0})$ is the inclusion map, and $(G/H,H)\to(G/K_{\pm},K_{\pm})$ is the map given by $gH\mapsto gK_{\pm}$. Thus, instead of the induced maps $(j_{\pm})_{\pm}\colon\pi_{1}(U_{+}\cap U_{-},x_{0})\to(U_{\pm},x_{0})$ we can consider the maps $(\rho_{\pm})_{\ast}\colon\pi_{1}(G/H,H)\to\pi_{1}(G/K_{\pm},K_{\pm})$
Using the fibration long exact sequence of homotopy groups:
$\cdots\pi_{1}(G/H,H)\overset{(\rho_{\pm})_{\ast}}{\longrightarrow}\pi_{1}(G/K_{\pm},K_{\pm})\to\pi_{0}(K_{\pm}/H,H)\to\cdots$
we conclude that $\pi_{1}(G/K_{\pm},K_{\pm})\cong\pi_{1}(G/H,H)/\ker\pi_{1}(G/H,H)$.
Putting all together, we obtain by van Kampen’s Theorem that
$\pi_{1}(X,x_{0})$ $\cong(\pi_{1}(U_{+},x_{0})\ast\pi_{1}(U_{-},x_{0}))/N(\mathrm{span}\{(j_{+})_{\ast}(\omega)(j_{-})_{\ast}(\omega)^{-1}\mid\omega_{\in}\pi_{1}(U_{+}\cap U_{-},x_{0})\})$
$\cong\big{(}(\pi_{1}(G/H)/N_{+})\ast(\pi_{1}(G/H)/N_{-})\big{)}/N_{0},$
where $N_{0}=N(\mathrm{span}\{(\rho_{+})_{\ast}(\omega)(\rho_{+})_{\ast}(\omega)^{-1}\mid\omega\in\pi_{1}(G/H)\})$, i.e. the normalizer of the subgroup generated by $(\rho_{+})_{\ast}(\omega)(\rho_{+})_{\ast}(\omega)^{-1}$. ∎
From Theorem 4.14 we obtain the following sufficient conditions for a compact geodesic cohomogeneity one space to be simply connected.
Corollary 4.15.
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $X^{\ast}$ is isometric to $[-1,1]$. We also consider a tuple $(G,H,G_{+},G_{-})$ given by Theorem 3.14 and assume that the orbits and $K_{\pm}/H$ are connected. If the inclusions
$(i_{\pm})_{\ast}\colon\pi_{1}(K_{\pm}/H)\to\pi_{1}(G/H)$
are surjective then $X$ is simply-connected.
Proof.
By the fibration long exact sequence of homotopy groups:
$\cdots\to\pi_{1}(K_{\pm}/H,H)\overset{(i_{\pm})_{\ast}}{\longleftrightarrow}\pi_{1}(G/H,H)\overset{(\rho_{\pm})_{\ast}}{\longrightarrow}\pi_{1}(G/K_{\pm},K_{\pm})\to\pi_{0}(K_{\pm}/H,H)\to\cdots$
we have that $\ker(\rho_{\pm})_{\ast}=\mathrm{img}(i_{\pm})_{\ast}$. Then the conclusion follows from Theorem 4.14. ∎
5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
In this Section we show how to construct a cohomogeneity one RCD space from a given cohomogeneity one group diagram, which we define next:
Definition 5.1.
[Cohomogeneity one diagram] Given a compact Lie group $G$, and Lie subgroups $H$, $G_{\pm}$ of $G$ satisfying the following diagram
[…]
we say that it is a cohomogeneity one group diagram if the homogeneous spaces $G_{\pm}/H$ admit $G_{\pm}$-invariant Riemannian metrics with positive Ricci curvature.
Remark 5.2.
We recall that by the characterization in [6], a homogeneous space $G/H$ admits a $G$-invariant metric of positive Ricci curvature if and only if it has finite fundamental group.
We now state our construction theorem.
Theorem B.
Let $G$ be a compact Lie group and $G_{+}$, $G_{-}$, $H$ Lie subgroups of $G$ such that $(G,H,G_{+},G_{-})$ is a cohomogeneity one group diagram. Then, there exists an $\mathsf{RCD}(K,N)$-space $(X,dist,\mathfrak{m})$ admitting a cohomogeneity one action of $G$ by measure preserving isometries, such that the associated group diagram is $(G,H,G_{+},G_{-})$. Moreover, we have $K\leq 0$, and $N=\max\{n_{-}+1+D,n_{+}+1+D,n_{0}+1\}$, where $n_{\pm}=\dim(G_{\pm}/H)$, $n_{0}=\dim(G/H)$ and $D=\dim(G)$.
5.1. Cone-like warped products.
In this section we will work with certain warped products over compact Riemannian manifolds. The distances of the warped products that we will be considering will always be defined on $[0,\infty)\times X/\sim$, where $(t,x)\sim(s,y)$ if and only if $t=s=0$. The warping function $f$ that we consider only vanish at $t=0$.
The next result tells us that under appropriate assumptions on the warping function we can obtain convexity of the closed balls around a particular point, namely, the “vertex” of the warped product.
Proposition 5.3.
Let $(X,d)$ be a compact geodesic space and take $f\colon[0,\infty)\rightarrow[0,\infty)$ a continuous non-decreasing function such that $f(0)=0.$ Then for every $R>0$ we have that $\bar{B}_{R}(0)=[0,R]\times_{f}X\subset[0,\infty]\times_{f}X$ is convex.
Proof.
Let $R>0,$ and consider $\gamma\colon[0,1]\rightarrow[0,\infty)\times_{f}X$ an admissible curve such that $\gamma_{0},\gamma_{1}\in[0,R]\times_{f}X$ but that $\gamma([0,1])\not\subset[0,R]\times_{f}X$. The goal is to find a curve $\tilde{\gamma}$ contained in $[0,R]\times_{f}X$ with length less than or equal to that of $\gamma$. We have $\gamma=(\alpha,\beta)$, where $\alpha\in AC([0,1],[0,\infty])$. Recall that the absolute value of an absolutely continuous function is still an absolutely continuous function. We define the admissible curve $\tilde{\gamma}=(\tilde{\alpha},\beta)$ where $\tilde{\alpha}(t):=\max\{0,R-|\alpha(t)-R|\}.$ Observe that $f^{2}\circ\tilde{\alpha}(t)\leq f^{2}\circ\alpha(t)$ for all $t\in[0,1]$ and that $|\dot{\tilde{\alpha}}_{t}|\leq|\dot{\alpha}_{t}|$ on a set of full measure. Therefore
[…]
and so we have the admissible curve we were looking for in $[0,R]\times X$. Hence we have the result. ∎
For our purposes we will consider the warping function $f_{A}\colon[0,\infty)\rightarrow[0,\infty)$ given by $f_{A}(x):=\min\{1,Ax\}$ for $A>0$. Observe that $f_{A}$ is continuous but not smooth at $x=1/A$. Then, given $\epsilon>0$ we define smooth functions $f_{A,\epsilon}$ to approximate $f_{A}$, by setting:
$f_{A,\epsilon}(x):=\frac{1}{2}\left((Ax+1)-\sqrt{(Ax-1)^{2}+\epsilon^{2}}-(1-\sqrt{1+\epsilon^{2}})\right).$
The following lemma is straightforward to prove.
Lemma 5.4.
We have for all $x\in[0,\infty)$ that $f_{A,\epsilon}(x)\geq 0$, $f_{A,\epsilon}(x)\leq(1+\sqrt{1+\epsilon^{2}})/2$. It also holds that for $x\in[0,2/A]$, $f_{A,\epsilon}(x)\leq f_{A}(x)$. Moreover
$|f_{A,\epsilon}(x)-f_{A}(x)|\leq\epsilon-(1-\sqrt{1-\epsilon^{2}})/2.$
Thus $f_{A,\epsilon}$ converges uniformly to $f_{A}$ as $\epsilon\to 0$.
We consider now warped products using the function $f_{A,\epsilon}$ as warping function, with the objective of showing pointed measured Gromov-Hausdorff convergence of these warped spaces to the warped space with respect to $f_{A}$. In view of the convexity granted by Proposition 5.3, and because of the specific properties of the $f_{A,\epsilon}$, henceforward we restrict ourselves to $[0,2/A]\times_{f_{A,\epsilon}}X$ and $[0,2/A]\times_{f_{A}}X$. Let $[t,x],[s,y]\in[0,2/A]\times X/\sim$ and consider the set
$\mathcal{A}:=\big{\{}\gamma\colon[0,1]\to[0,2/A]\times X/\sim\mid$ $\,\gamma=(\alpha,\beta),\>\alpha\in AC([0,1],[0,2/A]),$
$\beta\in AC([0,1],X),\>\gamma_{0}=[t,x],\gamma_{1}=[s,y]\big{\}}$
We set $\bar{\mathcal{A}}$ to be the closure of $\mathcal{A}$ with respect to the compact-open topology. Clearly $\bar{\mathcal{A}}=\bar{\mathcal{A}}([t,x],[s,y])$ (that is, $\bar{\mathcal{A}}$ depends on the points in question), but unless necessary we omit this in the notation. For simplicity, in the following we drop the dependence on the parameter $A$ from the notation, and simply write $f$ and $f_{\epsilon}$ instead of $f_{A}$ and $f_{A,\epsilon}$.
Similarly to what happens with $f_{\epsilon}$, $f$ we have the following result.
Lemma 5.5.
The length functionals $\ell_{\epsilon}$ and $\ell$ of the warped products $[0,2/A]\times_{f_{\epsilon}}X$ and $[0,2/A]\times_{f}X$, respectively, satisfy the following for all curves $\gamma\in\bar{\mathcal{A}}$:
  1. (1)
    $\ell_{\epsilon}(\gamma)\leq\ell(\gamma)$, for all $\epsilon>0$;
  2. (2)
    $\ell_{\epsilon_{1}}(\gamma)\leq\ell_{\epsilon_{2}}(\gamma)$, for all $0<\epsilon_{2}<\epsilon_{1}$.
Proof.
We first prove $(1)$. Fix $\epsilon>0$ and $\gamma=(\alpha,\beta)\in\mathcal{A}$. We know from the fact that $f_{\epsilon}\leq f$ for all $x\in[0,2/A]$ that
$\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{\epsilon}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}\leq\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}$
on the set of full measure where both $|\dot{\alpha}_{t}|$ and $|\dot{\beta}_{t}|$ are defined. On the remaining null set the functions are arbitrarily defined so they pose no problem. Now, integrating we get
[…]
Which implies that $\ell_{\epsilon}\leq\ell$. The proof of $(2)$ is carried out by a similar argument. ∎
We now establish pointwise convergence of the length functionals $\ell_{\epsilon}$ to $\ell$ as $\epsilon\to 0$.
Proposition 5.6.
The length functionals $\ell_{\epsilon}$ converge pointwise to $\ell$ when $\epsilon\to 0$.
Proof.
Take $\gamma\in\mathcal{A}$ and notice that for all $\epsilon>0$ the functions
$t\mapsto\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{\epsilon}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}$
are non-negative and measurable. Furthermore we have that for all $\epsilon_{1}<\epsilon_{2}$
$0\leq\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{\epsilon_{2}}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}\leq\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{\epsilon_{1}}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}\leq\infty$ And that due to the uniform convergence of the $f_{\epsilon}$, the pointwise limit is
$\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}=\lim_{\epsilon\rightarrow 0}\sqrt{|\dot{\alpha}_{t}|^{2}+f^{2}_{\epsilon}\circ\alpha(t)|\dot{\beta}_{t}|^{2}}.$
Then, by Beppo Levi’s Lemma (exchanging integration with supremum) we obtain:
[…]
Which is just what we wanted. ∎
In order to prove the pointed measured Gromov-Hasudorff convergence, we show that the induced distances converge pointwise; thus we need to check the convergence of the infima of the lengths $\ell_{\epsilon}$ on $\bar{\mathcal{A}}$ to the infimum of the length $\ell$ on $\bar{\mathcal{A}}$. To do so we make use of the notion of $\Gamma$-convergence. The interested reader may consult [9] for further details.
Proposition 5.7.
The length functionals $\ell_{\epsilon}$ $\Gamma$-converge to $\ell$ as $\epsilon\to 0$.
Proof.
Recall that the length functionals are lower semi-continuous. Now, from Lemma 5.5 we know that $\ell_{\epsilon}$ are non-decreasing with respect to $\epsilon$. Then using [9, Remark 2.12 ii)] we get that
$\Gamma\text{-}\lim_{\epsilon\rightarrow 0}\ell_{\epsilon}=\lim_{\epsilon\rightarrow 0}\ell_{\epsilon}=\ell.$
The next ingredient we need is the equicoercivity of the family $\{\ell_{\epsilon}\}$ (see [9, Definition 2.9]), which we now prove.
Proposition 5.8.
The sequence $\{\ell_{\epsilon}\}_{0<\epsilon\leq 1}$ is equicoercive.
Proof.
Let $\lambda\geq 0$ and consider the set $K_{\lambda}:=\{\gamma\mid\ell_{1}(\gamma)\leq 2\lambda\}$. As we are dealing only with curves in compact spaces, [13, Theorem 2.5.14] tells us that $K_{\lambda}$ is compact. We will only work with $0<\epsilon\leq 1$. Take $\gamma=(\alpha,\beta)\in\{\gamma\mid\ell_{\epsilon}(\gamma)\leq\lambda\}$. It is immediate to see that
[…]
Since $f_{1}\leq f_{\epsilon}$ for all $x\in[0,2/A]$, it follows that
[…]
So then
[…]
Hence we have equicoercivity for $0<\epsilon\leq 1$. ∎
We now have all the ingredients needed to prove the convergence of the distances. Before moving on to the next result, let us point out that so far we have only considered a fixed topology on $\bar{\mathcal{A}}$. However in order to apply the Fundamental Theorem of $\Gamma$-convergence we need to fix a metric on $\bar{\mathcal{A}}$ that induces the open-compact topology.
Note that for all $\epsilon>0$ we have that $f\leq 2f_{\epsilon}$. This is useful because it implies that the $f$-length $\ell$ is bounded from above by the $2f_{\epsilon}$-length, that is, $\ell\leq\ell_{2f_{\epsilon}}$. This will in turn yield that $d\leq d_{2f_{\epsilon}}$. Now, if $\gamma$ is an admissible curve then it has small $f_{\epsilon}$-length if and only if its $2f_{\epsilon}$-length is also small. Then, the induced distances $d_{\epsilon},d$ are equivalent. All this allows us to put on $\bar{\mathcal{A}}$ the sup norm induced by $d$ without any problem.
Proposition 5.9.
The distances $d_{\epsilon}$ induced by the lengths $\ell_{\epsilon}$ converge uniformly to the distance $d$ induced by $\ell$.
Proof.
Let $[t,x],[s,y]\in[0,1]\times X/\sim$. Then from Proposition 5.7 we have that the functionals $\ell_{\epsilon}\colon\bar{\mathcal{A}}\rightarrow[0,\infty]$ $\Gamma$-converge to the functional $\ell\colon\bar{\mathcal{A}}\rightarrow[0,\infty]$. Now Proposition 5.8 tells us that for $0<\epsilon\leq 1$ we have equicoercivity, so then we can apply the Fundamental Theorem of $\Gamma$-convergence (see [9, Theorem 2.10]) and obtain that:
$\inf_{\bar{\mathcal{A}}}\ell=\lim_{\epsilon\rightarrow 0}\inf_{\bar{\mathcal{A}}}\ell_{\epsilon},$
or equivalently that $d_{\epsilon}([t,x],[s,y])\rightarrow d([t,x],[s,y])$ as $\epsilon\rightarrow 0$. Since the points in $[0,1]\times X/\sim$ were chosen arbitrarily we have pointwise convergence of the distance functions. Now, it is easy to convince oneself that the following properties hold:
  • $d_{\epsilon}\leq d$ for all $\epsilon>0$.
  • $d_{\epsilon_{2}}\leq d_{\epsilon_{1}}$ for all $0<\epsilon_{1}<\epsilon_{2}$.
To conclude just observe that $d_{\epsilon}$, $d$ are continuous functions on a compact space so then by Dini’s Theorem we have that the convergence must be uniform. ∎
We are now able to show the measured Gromov-Hausdorff convergence of the warped product spaces.
Proposition 5.10.
The warped products $([0,2/A]\times_{f_{\epsilon}}X,d_{\epsilon})$ equipped with the reference measures $\mathfrak{m}_{\epsilon}:=f^{N}_{\epsilon}dt\otimes\mathfrak{m}_{X}$ for $N>1$ converge in the measured Gromov-Hausdorff topology to the warped product $([0,2/A]\times_{f}X,d)$ equipped with the measure $\mathfrak{m}:=f^{N}dt\otimes d\mathfrak{m}$.
Proof.
We show that the identity map $\mathrm{Id}\colon[0,2/A]\times_{f_{\epsilon}}X\rightarrow[0,2/A]\times_{f}X$ is a Gromov-Hausdorff-approximation. Define
$C_{\epsilon}=\dfrac{\int_{X}\!\int_{0}^{2/A}f^{N}\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}}{\int_{X}\!\int_{0}^{2/A}f^{N}_{\epsilon}\mathrm{d}t\,\mathrm{d}\mathfrak{m}}\geq 1.$
By the Dominated Convergence Theorem it is clear that as $\epsilon\rightarrow 0$, then $C_{\epsilon}\rightarrow 1$. With these constants we have that the measures $C_{\epsilon}f^{N}_{\epsilon}dt\otimes d\mathfrak{m}$ and $f^{N}dt\otimes d\mathfrak{m}$ have the same mass.
Let $\delta>0,$ then for all $\epsilon$ sufficiently small we have that
$\left|C_{\epsilon}-1\right|<\delta,\quad\left|f^{N}(t)-f^{N}_{\epsilon}(t)\right|<\delta,\quad\mbox{and}\>\left|d([t,x],[s,y])-d_{\epsilon}([t,x],[s,y])\right|<\delta$
hold for all $[t,x],[s,y]\in[0,1]\times X/\sim.$ Observe that this last condition tells us that $\mathrm{Id}$ is a $\delta$-Gromov-Hausdorff approximation for sufficiently small $\epsilon$.
We now show the weak convergence of measures. Take $\varphi\in C_{b}([0,2/A]\times_{f}X)$ and note that, since over $[0,2/A]$ we have $0\leq f_{\epsilon}(t)\leq 1$ then
$\left|\int(f^{N}(t)-C_{\epsilon}f^{N}_{\epsilon}(t))\varphi(t,x)\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}\right|\leq$ $\int|f^{N}(t)-C_{\epsilon}f^{N}_{\epsilon}(t)||\varphi|(t,x)\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}$
$\leq$ $\int|f^{N}(t)-f^{N}_{\epsilon}(t)||\varphi|(t,x)\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}$
$+\int|f_{\epsilon}^{N}(t)-C_{\epsilon}f_{\epsilon}^{N}(t)||\varphi|\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}$
$\leq$ $\int|f^{N}(t)-f^{N}_{\epsilon}(t)||\varphi|(t,x)\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}$
$+\int|f_{\epsilon}^{N}(t)||1-C_{\epsilon}||\varphi|\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}$
$\leq$ $\int|f^{N}(t)-f^{N}_{\epsilon}(t)||\varphi|(t,x)\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}$
$+\int|C_{\epsilon}-1||\varphi|\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}$
$\leq$ $2\delta\int|\varphi|(t,x)\,\mathrm{d}t\,\mathrm{d}\mathfrak{m}.$
So we also have the convergence of the measures. Hence we have the sought measured Gromov-Hausdorff convergence. ∎
We now prove that when we require $X$ to have small diameter, then near the boundary the metric on $[0,2/A]\times_{f}X$ is the product metric.
Proposition 5.11.
Let $(X,d_{X})$ be a compact geodesic metric space with $\operatorname{diam}(X)<1/A$. Then there exists $\delta>0$ such that the metric of the warped product $[0,2/A]\times_{f}X$ restricted to $[2/A-\delta,2/A]\times_{f}X$ is the product metric.
Proof.
First we prove that for $t$ sufficiently close to $2/A$ the geodesics joining points in the set $\{t\}\times X$ are always contained in $[1/A,2/A]\times X$. Since the warping function $f$ equals $1$ for all $t\in[1/A,2/A]$ this will yield that $\{t\}\times X$ is actually totally geodesic.
Proceeding by contradiction we assume that for all $\delta>0$ there exists $t\in[2/A-\delta,2/A]$ such that there are points $[t,x],[t,y]\in\{t\}\times X$ such that a minimal geodesic $\gamma$ joining them is not contained in $[1/A,2/A]\times X$. First observe that this means that $d([t,x],[t,y])<\operatorname{diam}(X)<1/A$ with the diameter in terms of the distance in $X$. Define the points
$t_{0}:=\min\{t\,|\,\gamma_{t}\in\{1/A\}\times X\}<t_{1}:=\max\{t\,|\,\gamma_{t}\in\{1/A\}\times X\}.$
It is clear that
$d([t,x],[t,y])=d([t,x],\gamma_{t_{0}})+d(\gamma_{t_{0}},\gamma_{t_{1}})+d(\gamma_{t_{1}},[t,y]).$
Now, since $\gamma_{t_{0}}\in\{1/A\}\times X$ we have that
$d([t,x],\gamma_{t_{0}})\geq d([t,x],\{1/A\}\times X)=d([t,x],[1/A,x])=|1/A-t|\geq 1/A-\delta,$
where the last inequality follows from recalling that $t\in[2/A-\delta,2/A]$. Analogously, we obtain the same lower bound for $d(\gamma_{t_{1}},[t,y])$. Thus for all $\delta>0$ we have that
$1/A>d([t,x],[t,y])\geq 2/A-2\delta+d(\gamma_{t_{0}},\gamma_{t_{1}}).$
This implies that $1/A>2/A+d(\gamma_{t_{0}},\gamma_{t_{1}})$, which is a contradiction. Hence we must have that there exists some $\delta>0$ such that for all $t\in[2/A-\delta,2/A]$ the set $\{t\}\times X$ is totally geodesic.
We will now prove that geodesics joining points in $[2/A-\delta,2/A]\times X$ must lie there. Again, we proceed by contradiction: take two points $[r,x],[s,y]$ with $r<s$, and a minimal geodesic $\gamma$ with $\gamma_{0}=[r,x],\gamma_{1}=[s,y]$ such that it is not contained in $[2/A-\delta,2/A]\times X$. Now consider $t_{0}:=\max\{t\,|\,\gamma_{t}\in\{r\}\times X\}.$ Notice that then $\gamma([t_{0},1])\subset[2/A-\delta,2/A]\times X$. The restriction of $\gamma$ to the interval $[0,t_{0}]$ can be reparametrized to obtain a geodesic between $[r,x]$ and $\gamma_{t_{0}}.$ However, the previous argument implies then that $\gamma([0,t_{0}])\subset[2/A-\delta,2/A]\times X$, giving the contradiction.
Finally, as we have proved that geodesics between points in $[2/A-\delta,2/A]\times X$ must lie in there, it follows that the metric must be the product one. ∎
We are now tasked with looking at the structure of the Sobolev space $W^{1,2}$ of the warped product we have considered. We use the tools developed in [34] where Gigli and Han give conditions to ensure that the minimal weak upper gradient of a Sobolev function $g$ in the warped product $[0,\infty)\times_{f}X$ can be described in terms on the minimal weak upper gradients of the functions $g^{(t)}:=g(t,\cdot)$ and $g^{(x)}:=g(\cdot,x)$.
Proposition 5.12.
Let $(X,d,\mathfrak{m})$ be an infinitessimally Hilbertian space, and consider $f\colon[0,\infty)\rightarrow[0,\infty)$ to be a warping function such that $\{f=0\}$ is discrete and $f$ decays at least near its zeroes (see equations 3.14, 3.15 in [34]). Then the warped product $[0,\infty)\times_{f}X$ is infinitessimally Hilbertian.
Proof.
Take $h,g\in W^{1,2}([0,1]\times_{f}X)$. Then by [34, Propositions 3.10 and 3.12], we have that the minimal weak upper gradient of $h$ is given by:
[…]
And similarly for $g$. [36, Theorem 4.3.3] states that $W^{1,2}([0,1]\times_{f}X)$ being a Hilbert space is equivalent to a pointwise parallelogram law. So we just need to check that this parallelogram law is satisfied from the hypothesis that both $[0,1]$ and $X$ satisfy such a parallelogram law:
$2\left(|\nabla h|^{2}(t,x)+|\nabla g|^{2}(t,x)\right)=$ $2\big{(}f^{-2}(t)|\nabla h^{(t)}|^{2}(x)+|\nabla h^{(x)}|^{2}(t)$
$+f^{-2}(t)|\nabla g^{(t)}|^{2}(x)+|\nabla g^{(x)}|^{2}(t)\big{)}$
$=$ $f^{-2}(t)2\left(|\nabla h^{(t)}|^{2}(x)+|\nabla g^{(t)}|^{2}(x)\right)$
$+2\left(|\nabla h^{(x)}|^{2}(t)+|\nabla g^{(x)}|^{2}(t)\right)$
$=$ $f^{-2}(t)\left(|\nabla(h+g)^{(t)}|^{2}(x)+|\nabla(h-g)^{(t)}|^{2}(x)\right)$
$+\left(|\nabla(h+g)^{(x)}|^{2}(t)+|\nabla(h-g)^{(x)}|^{2}(t)\right)$
$=$ $|\nabla(h+g)|^{2}(t,x)+|\nabla(h-g)|^{2}(t,x)$
We now turn our attention to the problem of finding a curvature-dimension bound for the warped products $([0,2/A]\times_{f}X,d)$, when we assume that $X$ satisfies a curvature dimension bound.
Recall that given $K\in\mathbb{R}$ and $B$ a geodesic space, we say that a function $f\colon B\to[0,\infty)$ is $\mathcal{F}K$-concave if for every geodesic $\gamma$ we have:
(4) $(f\circ\gamma)^{\prime\prime}(t)+K(f\circ\gamma)(t)\leq 0.$
With this definition in hand we recall the following result of Ketterer (building upon the work of Alexander and Bishop), regarding curvature bounds and $\mathcal{F}K$-concavity.
Theorem 5.13 (Theorem A in [52]).
Let $B$ be a complete $d$-dimensional (local) Alexandrov space of $\mathrm{curv}\geq K$, such that $B\setminus\partial B$ is a smooth Riemannian manifold. Let $f\colon B\to[0,\infty)$ be $\mathcal{F}K$-concave and smooth over $B\setminus\partial B$. Assume that $\partial B\subseteq f^{-1}(0)$. Let $(F,\mathfrak{m}_{F})$ be a weighted complete Finsler manifold. Let $N\geq 1$, and $K_{F}\in\mathbb{R}$. If $N=1$ and $K_{F}>0$, we assume that $\mathrm{diam}(F)\leq\pi/\sqrt{K_{F}}$. In any case, $F$ satisfies the $\mathsf{CD}^{\ast}((N-1)K_{F},N)$-condition. Moreover, we assume that
  1. (1)
    if $\partial B=\varnothing$, suppose $K_{F}\geq Kf^{2}$;
  2. (2)
    if $\partial B\neq\varnothing$, suppose $K_{F}\geq 0$ and $|\nabla f|_{p}\leq\sqrt{K_{F}}$ for all $p\in\partial B$.
Then the $N$-warped product $B\times_{f}^{N}F$ satisfies the $\mathsf{CD}^{\ast}((N+d-1)K,N+d)$-condition.
As we apply this result in our context, we first show the $\mathcal{F}K$-concavity of the approximating functions $f_{A,\epsilon}$ for $K=0$.
Lemma 5.14.
The functions $f_{A,\epsilon}$ are $\mathcal{F}K$-concave for $K=0$.
Proof.
We compute the second derivatives of $f_{A,\epsilon}$:
$f_{A,\epsilon}^{\prime\prime}(x)=$ $\frac{d}{dx}\left(\frac{1}{2}\left(A-A(Ax-1)((Ax-1)^{2}+\epsilon^{2})^{-\frac{1}{2}}\right)\right)$
$=$ $-\frac{A}{2}\Big{(}A((Ax-1)^{2}+\epsilon^{2})^{-\frac{1}{2}}-(Ax-1)\frac{1}{2}((Ax-1)^{2}+\epsilon^{2})^{-\frac{3}{2}}2(Ax-1)A\Big{)}$
$=$ $-\frac{A^{2}}{2}\Big{(}((Ax-1)^{2}+\epsilon^{2})^{-\frac{1}{2}}-(Ax-1)^{2}((Ax-1)^{2}+\epsilon^{2})^{-\frac{3}{2}}\Big{)}$
$=$ $-\frac{A^{2}}{2}\Big{(}((Ax-1)^{2}+\epsilon^{2})^{-\frac{3}{2}}((Ax-1)^{2}+\epsilon^{2})-(Ax-1)^{2}((Ax-1)^{2}+\epsilon^{2})^{-\frac{3}{2}}\Big{)}$
$=$ $-\frac{A((Ax-1)^{2}+\epsilon^{2})^{-\frac{3}{2}}\epsilon^{2}}{2}$
$=$ $-\frac{A\epsilon^{2}}{\sqrt{((Ax-1)^{2}+\epsilon^{2})^{3}}}\leq 0$
Now we consider $(F,d_{F},\mathfrak{m}_{F})$ to be a weighted Finsler (Riemannian) manifold satisfying the $\mathsf{CD}^{\ast}((N+1)K_{F},N)$-condition (respectively, the $\mathsf{RCD}((N-1)K_{F},N)$-condition). We also consider $B=[0,\infty)$.
Theorem 5.15.
Let $(F,\mathfrak{m}_{F})$ be a weighted Finsler manifold which satisfies the $\mathsf{CD}((N-1)K_{F},N)$-condition for $K_{F}\geq 0$ such that, if $N=1$ then $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$, and $\sqrt{K_{F}}\geq(A/2)(1+(1/\sqrt{1+\epsilon^{2}}))$. Then the $N$-warped product $[0,\infty)\times^{N}_{f_{A,\epsilon}}F$ is an $\mathsf{CD}^{\ast}(NK,N+1)$-space for $K=0$. Moreover, if $(F,\mathfrak{m}_{F})$ is infitesimally Hilbertian, then $[0,\infty)\times^{N}_{f_{A,\epsilon}}F$ is an $\mathsf{RCD}(NK,N+1)$-space for $K=0$.
Proof.
The interval $[0,\infty)$ with the standard Euclidean metric is an Alexandrov space of $\operatorname{curv}\geq 0$. By Lemma 5.14 a we have that $f_{A,\epsilon}$ is $\mathcal{F}K$-concave for $K=0$. Moreover we have $|f_{A,\epsilon}^{\prime}|\leq\sqrt{K_{F}}$. Thus the hypothesis of Theorem 5.13 are satisfied yielding the first claim. Moreover, when $(F,\mathfrak{m}_{F})$ is infinitesimally Hilbertian, we obtain the second conclusion from Proposition 5.12. ∎
With this result, we are able to show that spaces in Theorem A (b) admit an $\mathsf{RCD}(0,N)$-structure on which $G$ acts by measure preserving isometrics of cohomogeneity one.
Theorem I.
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space, and let $G$ be a compact Lie group acting on $X$ by measure preserving isometries, such that the orbit space $X/G$ is homeomorphic to $[0,\infty)$. Let $G_{0}$ be the unique singular isotropy group, and $H$ the principal isotropy group. Under the assumption that $G_{0}/H$ is simply-connected, there exists a new metric $\tilde{d}$ and measure $\tilde{\mathfrak{m}}$ on $X$ making into an $\mathsf{RCD}(0,N^{\prime})$-space, where $N^{\prime}=\dim(G_{0}/H)+1+\dim(G)$. Moreover, $G$ acts on $(X,\tilde{d},\tilde{\mathfrak{m}})$ by measure preserving isometries with cohomogeneity one.
Proof.
We consider on $G$ a bi-invariant Riemannian metric. By Lemma 4.8 and Theorem D, we have that cone fibers which are homeomorphic to the space $F=G_{0}/H$. As proven in [6], the restriction of the $G$-bi-inviariant Riemannian metric to $G_{0}$ induces a $G_{0}$ Riemannian metric of positive Ricci curvature on $F$. Thus the space $(F,d_{0},\mathfrak{m}_{0})$ with the induced metric and measure is an $\mathsf{RCD}(K_{0},N_{0})$-space with $K_{0}>0$, and $N_{0}=\dim(F)$. Thus there exists $K_{F}>0$ such that $K_{0}=(N_{0}-1)K_{F}$.
Observe that in the case when $N_{0}=1$, then by taking $\lambda\geq\operatorname{diam}(F)/(\pi/\sqrt{K_{F}})$, we have that by rescaling $F$ by $\lambda$, the rescaled space $(F,\lambda d_{0},(\lambda)^{N_{0}}\mathfrak{m}_{0})$ is an $\mathsf{RCD}((N_{0}-1)K_{F}/\lambda,N_{0})$-space which satisfies $\operatorname{diam}(F,\lambda d_{0})\leq\pi\lambda^{2}/\sqrt{K_{F}}$. Thus without loss of generality, we may assume that when $N_{0}=1$, we have $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$.
We consider now $A>0$ and $\varepsilon>0$ appropriately such that $\sqrt{K_{F}}\geq(A/2)(1+(1+\sqrt{1+\epsilon^{2}}))$. Then by Theorem 5.15 we have that the $N_{0}$-warped product $[0,\infty)\times_{f_{A,\epsilon}}F$ is an $\mathsf{RCD}(0,N_{0}+1)$-space. Thus the product metric measure space $G\times([0,\infty)\times_{f_{A,\epsilon}}F)$ is an $\mathsf{RCD}(0,N_{0}+1+D)$-space, where $D=\dim(G)$, and $G_{0}$ acts by measure preserving isometries via the diagonal action $\Delta\colon G\times([0,\infty)\times_{f_{A,\epsilon}}F)$ given by
$\Delta(g_{0},(g,[t,p]):=(gg_{0}^{-1},[t,g_{0}p]).$
Thus by [27, Theorem 1.1] the quotient space $(G\times([0,\infty)\times_{f_{A,\epsilon}}F))/\Delta$ is an $\mathsf{RCD}(0,N_{0}+1+D)$ space. But by Theorem A this space is homeomorphic to $X$. ∎
We now consider $N$-warped products of the form $[0,2/A]\times_{f_{A}}^{N}F$:
Theorem 5.16.
Let $(F,d_{F},\mathfrak{m}_{F})$ be a weighted Finsler manifold which satisfies the $\mathsf{CD}((N-1)K_{F},N)$-condition for $K_{F}\geq 0$ such that, if $N=1$ then $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$. Assume that $K_{F}\geq A^{2}$. Then the $N$-warped product $[0,2/A]\times_{f_{A}}^{N}F$ is a $\mathsf{CD}^{\ast}(0,N+1)$-space, and an $\mathsf{RCD}(0,N+1)$-space if $(F,d_{F},\mathfrak{m}_{F})$ satisfies the $\mathsf{RCD}((N-1)K_{F},N)$-condition.
Proof.
By the previous theorem, the $N$-warped products $[0,\infty)\times_{f_{A,\epsilon}}^{N}F$ are $\mathsf{CD}^{\ast}(0,N+1)$-spaces (respectively $\mathsf{RCD}(0,N+1)$-spaces if $F$ is infinitesimally Hilbertian).
We now consider the open subset $[0,2/A)\times F/(0,p)\sim(0,q)$. This is the open ball of radius $2/A$ centered at the vertex $\ast$ of the cones, with respect to the metrics $d_{A}$ or $d_{\epsilon}$. Moreover as we showed, it is totally geodesic, and by construction we have that $\mathfrak{m}_{\epsilon}(\partial[0,2/A)\times F/(0,p)\sim(0,q))=0=\mathfrak{m}_{A}(\partial[0,2/A)\times F/(0,p)\sim(0,q))$. Thus we conclude that $[0,2/A]\times_{f_{A,\epsilon}}F$ are $\mathsf{CD}^{\ast}(0,N+1)$-spaces (respectively, $\mathsf{RCD}(0,N+1)$-spaces when $F$ is infinitesimally Hilbertian).
Since the $N$-warped product $[0,2/A]\times_{f_{A}}^{N}F$ is the pointed measured Gromov-Hausdorff limit of the spaces $[0,2/A]\times_{f_{A,\epsilon}}^{N}F$ as $\epsilon\to 0$, and the $\mathsf{CD}^{\ast}$, $\mathsf{RCD}$-conditions are stable under this type of convergence, we obtain the desired result. ∎
We observe that for spaces with positive Ricci curvature and dimension at least $2$ we can also consider the the function $f_{2}$ given by $A=2$, due to the following result:
Theorem 5.17 ((9) in [58]).
If $(X,d,\mathfrak{m})$ is $\mathsf{CD}^{\ast}(K,N)$, then for $\lambda>0$ the space $(X,\lambda d,\mathfrak{m})$ is $\mathsf{CD}^{\ast}(\lambda^{-2}K,N)$.
We consider an $n$-dimensional closed Riemannian manifold $F$ such that for $0<\kappa_{F}$ we have $\kappa_{F}\leq\mathrm{Ric}(F)$, where $\mathrm{Ric}(F)=\min\{\mathrm{Ric}(x)\mid x\in TM\>\|x\|=1\}$. Thus we have that $(F,d_{F},\mathfrak{m}_{F})$ is a $\mathsf{CD}^{\ast}(\kappa_{F},n)$-space. By taking $\lambda=\sqrt{\kappa_{F}/4(n-1)}$ we have that $(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{CD}^{\ast}((n-1)4,n)$-space. Given that $F$ is a Riemannian manifold, we obtain that $(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n)$-space, and also $(F,\mathfrak{m}_{F})$ is a weighted Finsler manifold. Thus we have that $|f_{2,\epsilon}^{\prime}(0)|\leq 2$, and by Theorem 5.15 that $[0,\infty)\times_{f_{2,\epsilon}}(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n+1)$-space. Since the balls of radius $1$ are geodesic, then by taking the limit as $\epsilon\to 0$ we obtain that that $[0,1]\times_{f_{2}}(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n+1)$-space.
5.2. Gluing of $\mathsf{RCD}$-cone-bundles.
We now consider gluings along the boundary of (certain quotients of) the $N$-warped products we have considered so far using the functions $f_{A}\colon[0,2/A]\to[0,\infty)$. Consider the cone-like warped product $C=[0,2/A]\times F/(0,p)\sim(0,q)$ with the distance $d_{C}=d_{A}$ and measure $\mathfrak{m}_{C}=f^{n-1}_{A}d\mathfrak{m}_{F}\otimes d\mathcal{L}$, where $(F,d_{F},\mathfrak{m}_{F})$ is an (infinitesimally Hilbertian) weighted Finsler manifold satisfying the $\mathsf{CD}^{\ast}((N-1)K_{F},N)$-condition for some $K_{F}>0$. We consider $G$ equipped with a bi-invariant metric $d_{G}$ and the Haar measure $\mathfrak{m}_{G}$. We take $G\times C$ equipped with the product distance $d_{G}\oplus d_{C}$ and measure $\mathfrak{m}_{G}\otimes\mathfrak{m}_{C}$. Thus, taking $D:=\dim(G)$, by Theorem 5.16 the space $G\times C$ is an $\mathsf{RCD}(0,N+1+D)$-space.
We further assume that a Lie subgroup $K\subset G$ acts by measure preserving isometries on $F$, the fiber of the cone $C$. Thus $K$ also acts by measure preserving isometries on the metric product $G\times C$ by a diagonal action. That is, each $h\in K$ acts by
$h(g,[t,x]):=(gh^{-1},[t,hx]),$
where, $hx$ represents the $K$-action on $F$. Thus the quotient space $(G\times C)/\Delta$ equipped with the quotient distance $d^{\ast}$ and quotient measure $\mathfrak{m}^{\ast}$ is an $\mathsf{RCD}(0,N+1+D)$-space by [27].
We collect this conclusion in the following lemma:
Lemma 5.18.
Consider $G$ an $D$-dimensional compact Lie group, $K\subset G$ a Lie subgroup, and $(F,d_{F},\mathfrak{m}_{F})$ an infinitesimally Hilbertian weitghted Finsler manifold, which satisfies the $\mathsf{RCD}((N-1)K_{F},N)$-condition on which $K$ acts transitively by measure preserving isometries for some $K_{F}>0$. Assume that $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$, if $N=1$. Consider $A>0$ such that $K_{F}>A^{2}$, and the function $f_{A}(x):=\min\{1,Ax\}$. Consider the warped product $C:=[0,2/A]\times_{f_{A}}^{N}F$, and denote by $\Delta$ the diagonal action of $K$ on $G\times C$. Then the space $G\times C/\Delta$ is an $\mathsf{RCD}(0,N+1+D)$-space.
Moreover when we consider a homogeneous space $X=G/H$, equipped with $G$-invariant distance $d$ and measure $\mathfrak{m}$ making it an $\mathsf{RCD}$-space, the measure $\mathfrak{m}$ has to be (a multiple of) the Hausdorff measure as we now show (this has also been independently proved by [48, Theorem 2]).
Proposition 5.19.
Let $G$ be a compact Lie group acting transitively by measure preserving isometries on an $\mathsf{RCD}(K,N)-$space $(X,d,\mathfrak{m})$. Then $\mathfrak{m}$ equals a multiple of the $n$-dimensional Hausdorff measure of $X$, where $n$ is the topological dimension of $X$.
Proof.
By [68, Proposition 5.14] we have that $X$ must be isometric to a Riemannian manifold. So, in particular we have that all the tangent spaces must be isometric to $\mathbb{R}^{n}$ for some natural number $n$. Then, by [12, Theorem 1.12] we have then that $\mathfrak{m}=\rho\mathcal{H}^{n}$, where $\mathcal{H}^{n}$ denotes the $n$-dimensional Hausdorff measure. We claim now that for all $g\in G$. $\rho\circ g=\rho$. Let $A\in\mathcal{B}(X)$, $g\in G$, and recall that $g_{\#}\mathcal{H}^{n}=\mathcal{H}^{n}$. As the action is by measure preserving isometries $\mathfrak{m}(A)=g_{\#}\mathfrak{m}(A)$. This implies
[…]
So the density $\rho$ is $G$-invariant. By the Lebesgue differentiation theorem we have that for $\mathcal{H}^{n}$-a.e. $x\in X$
$\rho(x)=\lim_{r\rightarrow 0}\frac{\mathfrak{m}(B_{r}(x))}{\mathcal{H}^{n}(B_{r}(x))}.$
But notice that the limit on the RHS is the same if we put $gx$ instead of $x$. So then $\rho(x)=\rho(gx)$ $\mathcal{H}^{n}$-a.e. This implies then that $\rho$ must be constant. ∎
Theorem 5.20.
Consider $G_{\pm}$ compact Lie groups, $(P_{\pm},d_{\pm}^{P},\mathfrak{m}_{\pm}^{P})$ be $N_{\pm}^{P}$-dimensional Riemannian manifolds with $\operatorname{Ric}(P_{\pm})\geq K^{P}_{\pm}$, and $(F_{\pm},d_{\pm}^{F},\mathfrak{m}_{\pm}^{F})$ be weighted Riemannian manifolds which are $\mathsf{RCD}(K_{\pm}^{F},N_{\pm}^{F})$ with $K_{\pm}^{F}>0$. Assume that $G_{\pm}$ acts effectively on $P_{\pm}$ and $F_{\pm}$ by measure preserving isometrties. Furthemore assume that the balanced product $P_{-}\times_{G_{-}}F_{-}$ is homeomorphic to $P_{+}\times_{G_{+}}F_{+}$. Then the glued space
$X=(P_{-}\times_{G_{-}}\mathrm{Cone}(F_{-}))\cup_{\partial}(P_{-}\times_{G_{+}}\mathrm{Cone}(F_{+}))$
admits an $\mathsf{RCD}(K,N)$-structure. Here $\mathrm{Cone}(F_{\pm})$ denotes the topological cone over $F_{\pm}$.
Proof.
We choose $K_{\pm}>0$ so that $(F_{\pm},d^{F}_{\pm},\mathfrak{m}_{\pm}^{F})$ are $\mathsf{RCD}((N_{\pm}^{F}-1)K_{\pm},N^{F}_{\pm})$, for example by taking $K_{\pm}\leq K_{\pm}^{F}/(N^{F}_{\pm}-1)$. Observe that in the case when either $N^{F}_{+}=1$ or $N^{F}_{-}=1$, then by taking $\lambda_{\pm}\geq\operatorname{diam}(F_{\pm})/(\pi/\sqrt{K_{\pm}})$, we have that by rescaling $F_{\pm}$ by $\lambda_{\pm}$, the rescaled space $(F_{\pm},\lambda_{\pm}d^{F}_{\pm},(\lambda_{\pm})^{N^{F}_{\pm}}\mathfrak{m}^{F}_{\pm})$ is an $\mathsf{RCD}((N^{F}_{\pm}-1)K_{\pm}/\lambda_{\pm},N^{F}_{\pm})$-space which satisfies $\operatorname{diam}(F_{\pm},\lambda_{\pm}d^{F}_{\pm})\leq\pi\lambda_{\pm}^{2}/\sqrt{K_{\pm}}$. Thus without loss of generality, we may assume that if either $N^{F}_{+}=1$ or $N^{F}_{-}=1$, we have $\operatorname{diam}(F_{\pm})\leq\pi/\sqrt{K_{\pm}}$.
We now consider $A_{\pm}>0$ such that $K_{\pm}>A_{\pm}^{2}$. We now consider the following two $\mathsf{RCD}(0,N^{F}_{\pm}+1+N^{P}_{\pm})$-spaces:
$\Big{(}(P_{\pm}\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/G_{\pm},(d^{P}_{\pm}\otimes d_{A_{\pm}})^{\ast},(\mathfrak{m}^{P}_{\pm}\otimes d\mathfrak{m}_{A_{\pm}})_{\ast}\Big{)},$
[…]
We observe that the boundaries of the cone bundles $P_{\pm}\times_{G_{\pm}}([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}\to P_{\pm}/G_{\pm}$ are homeomorphic to $P_{+}\times_{G_{+}}(F_{+})=(P_{+}\times F_{+})/G_{+}=\{2/A_{+}\}\times((P_{+}\times F_{+})/G_{+})$ and $P_{-}\times_{G_{-}}(F_{-})=(P_{-}\times F_{-})/G_{-}=\{2/A_{-}\}\times((P_{-}\times F_{-})/G_{-})$ respectively, and by hypothesis both are homeomorphic to a space $B$.
We also recall that by construction $B$ with the induced metric is geodesic in $(P_{\pm}\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/K_{\pm}$ with respect to the the two $\mathsf{RCD}(0,N^{F}_{\pm}+1+N^{P}_{\pm})$-structures $(d_{\pm},\mathfrak{m}_{\pm})$. Moreover, over the open collars $(P_{\pm}\times((1/A_{\pm},2/A_{\pm}]\times F_{\pm}))/G_{\pm}$ the metric structures $(d_{\pm},\mathfrak{m}_{\pm})$ are the product structures $(d_{\mathbb{R}}\oplus d^{P}_{\pm}\oplus d^{F}_{\pm},f^{N^{F}_{\pm}}_{\pm}d\mathcal{L}\otimes d\mathfrak{m}^{P}_{\pm}\otimes d\mathfrak{m}^{F}_{\pm})$. From this we observe that induced metric structures on $B$ are of the form $((d^{P}_{\pm}\otimes d^{F}_{\pm})^{\ast},(\mathfrak{m}^{P}\pm\otimes\mathfrak{m}^{F}_{\pm})_{\ast})$ and are given by a Riemannian metric $g_{\pm}$ on $B$: since $P_{\pm}\times F_{\pm}$ with $g_{P_{\pm}}\oplus g_{F_{\pm}}$ is a Riemannian manifold inducing $(d^{P}_{\pm}\oplus d^{F}_{\pm},\mathfrak{m}^{P}_{\pm}\oplus\mathfrak{m}^{F}_{\pm})$, and the action of $G_{\pm}$ over $P_{\pm}\times F_{\pm}$ is free, we obtain quotient Riemannian metrics $g_{\pm}$ on $B$ inducing $((d^{P}_{\pm}\oplus d^{F}_{\pm})^{\ast},(\mathfrak{m}^{P}_{\pm}\otimes\mathfrak{m}^{F}_{\pm})_{\ast})$.
We consider $[-1,1]\times B$, and consider $g(t)$ a path in the set of Riemannian metrics on $B$ with $g(t)=g_{-}$ over $[-1,-1+\varepsilon]$ and $g(t)=g_{+}$ over $[1-\varepsilon,1]$. We set $C_{0}:=[-1,1]\times B$ and equip it with the metric $d_{0}$ and measure $\mathfrak{m}_{0}$ induced by the Riemannian metric $h=dt^{2}\oplus g(t)$. We observe that since $([-1,1]\times B,h)$ is a compact Riemannian manifold, then there exists $K_{0}\in\mathbb{R}$, $|K_{0}|\leq\infty$ such that $\operatorname{Ric}(h)\geq K_{0}$. This implies that $(C_{0},d_{0},\mathfrak{m}_{0})$ is an $\mathsf{RCD}(K_{0},\dim(B)+1)$-space.
We identify $(P_{-}\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ with $C_{-}:=(P_{-}\times([-1-2/A_{-},-1]\times F/(-1-2/A_{-},p)\sim(-1-2/A_{-},q))/G_{-}$ and push the $\mathsf{RCD}(0,N^{F}_{-}+1+N^{P}_{-})$-structure $(d_{-},\mathfrak{m}_{-})$ of $(P_{-}\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ to get an $\mathsf{RCD}(0,N^{F}_{-}+1+N^{P}_{-})$-structure on $C_{-}$, which we also denote by $(d_{-},\mathfrak{m}_{-})$ abusing notation. We also identify $(P_{+}\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ with $C_{+}:=(P_{+}\times([1,1+2/A_{+}]\times F/(1+2/A_{+},p)\sim(1+2/A_{+},q))/G_{+}$ and push the $\mathsf{RCD}(0,N^{F}_{+}+1+N^{P}_{+})$-structure $(d_{+},\mathfrak{m}_{+})$ of $(P_{+}\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ to get an $\mathsf{RCD}(0,n_{+}1+D)$-structure on $C_{+}$, which we also denote by $(d_{+},\mathfrak{m}_{+})$.
We consider the glued space
$X_{1}:=C_{-}\bigcup_{\mathrm{Id}_{B}}C_{0}.$
On $X_{1}$ we define a metric $d_{1}$ given by
$d_{1}(x,y)=\begin{cases}d_{-}(x,y),&x,y\in C_{-}\setminus\partial C_{-}=C_{-}\setminus B,\\ d_{0}(x,y),&x,y\in C_{0}\setminus(\{-1\}\times B),\\ \min\{d_{-}(x,a)+d_{0}((-1,a),y)\mid a\in B\},&x\in C_{-}\mbox{ and }y\in C_{0}.\\ \end{cases}$
We also set the measure $\mathfrak{n}_{1}:=(i_{-})_{\ast}(\mathfrak{m}_{-})+(i_{0})_{\ast}(\mathfrak{m}_{0})$, where $i_{-}\colon C_{-}\to X$ and $i_{0}\colon C_{0}\to X$ are the respective inclusions. Also observe that for $K_{1}=\min\{K_{0},0\}$ and $N_{1}=\max\{N^{F}_{-}+1+N^{P}_{\pm},\dim(B)+1\}$, we have that $C_{0}$, $C_{-}$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces. We consider $\delta_{1}<\min\{\varepsilon,1/A_{-}\}$ as well as the open subsets $U_{-}=C_{-}\cup([-1,-1+\delta)\times B\subset X_{1}$, and $U_{0}=(-1-\delta,1]\times B$, and observe that they are an open cover of $X_{1}$ due to the fact that the metrics $d_{-}$ and $d_{0}$ are the product metrics over the preimage of $[-1-(2/A_{-})-\delta,-1+\delta]\subset X_{1}^{\ast}$ under the quotient map $\pi\colon X_{1}\to X_{1}^{\ast}$. Moreover, since the metric $d_{1}$ and measure $\mathfrak{m}_{1}$ are the product over $[-1-\delta,-1+\delta]$, we have that the spaces $(\bar{U}_{-},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{-}})$, $(\bar{U}_{0},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{0}})$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces by construction. Thus by Theorem 2.11, we conclude that $X_{1}$ is an $\mathsf{RCD}(K_{1},N_{1})$-space.
In analogous fashion we can prove that the spaces
$X=X_{1}\bigcup_{\mathrm{Id}_{B}}C_{+}$
is an $\mathsf{RCD}(K,N)$-space where $K=K_{1}\leq 0$ and $N=\max\{N_{1},N^{P}_{+}+1+N^{P}_{+}\}$. ∎
We now prove Theorem B.
Proof of Theorem B.
We set $F_{\pm}$ to be the homogeneous Riemannian manifolds $G_{\pm}/H$ with the $G$-invariant Riemannian metrics of positive Ricci curvature. Then they are $\mathsf{RCD}((n_{\pm}-1)K_{\pm},n_{\pm})$-spaces, for $n_{\pm}=\dim(F_{\pm})$ and some $K_{\pm}>0$ fixed. Observe that in the case when either $n_{+}=1$ or $n_{-}=1$, then by taking $\lambda_{\pm}\geq\operatorname{diam}(F_{\pm})/(\pi/\sqrt{K_{\pm}})$, we have that by rescaling $F_{\pm}$ by $\lambda_{\pm}$, the rescaled space $(F_{\pm},\lambda_{\pm}d_{F_{\pm}},(\lambda_{\pm})^{n_{\pm}}\mathfrak{m}_{F_{\pm}})$ is an $\mathsf{RCD}((n_{\pm}-1)K_{\pm}/\lambda_{\pm},n_{\pm})$-space which satisfies $\operatorname{diam}(F_{\pm},\lambda_{\pm}d_{F_{\pm}})\leq\pi\lambda_{\pm}^{2}/\sqrt{K_{\pm}}$. Thus without loss of generality, we may assume that if either $n_{+}=1$ or $n_{-}=1$, we have $\operatorname{diam}(F_{\pm})\leq\pi/\sqrt{K_{\pm}}$.
We now consider $A_{\pm}>0$ such that $K_{\pm}>A_{\pm}^{2}$. We now consider the following two $\mathsf{RCD}(0,n_{\pm}+1+D)$-spaces:
$\Big{(}(G\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/G_{\pm},(d_{G}\otimes d_{A_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes d\mathfrak{m}_{A_{\pm}})_{\ast}\Big{)},$
[…]
We showed in Proposition 5.11 that over $(1/A_{\pm},2/A_{\pm}]\times F_{\pm}\subset[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}$ we have $d_{A_{\pm}}=d_{\mathbb{R}}\oplus d_{F_{\pm}}$. Since the action of $G_{\pm}$ on the $[0,2/A\pm]$ factor is free, we see that the image of $G\times[1/A_{\pm},2/A_{\pm}]\times F_{\pm}$ under the quotient map $G\times([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})\to(G\times([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}))/\Delta$ is isometric to
$\Big{(}[1/A_{\pm},2/A_{\pm}]\times\big{(}(G\times F_{\pm})/\Delta\big{)},d_{\mathbb{R}}\oplus(d_{G}\otimes d_{F_{\pm}})^{\ast},f_{\pm}^{n_{\pm}}d\mathcal{L}\otimes(\mathfrak{m}_{G}\otimes d\mathfrak{m}_{F_{\pm}})_{\ast}\Big{)}.$
We observe that the boundaries of the cone bundles $G\times_{G_{\pm}}([0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm}\to G/G_{\pm}$ are $G/H\cong G\times_{G_{+}}(G_{+}/H)=(G\times F_{+})/G_{+}=\{2/A_{+}\}\times((G\times F_{+})/G_{+})$ and $G/H\cong G\times_{G_{-}}(G_{-}/H)=(G\times F_{-})/G_{-}=\{2/A_{-}\}\times((G\times F_{-})/G_{-})$. Thus they are homeomorphic to $G/H$. From now on we set $B=G/H$.
We also recall that by construction the space $B$ with the induced metric is geodesic in $(G\times[0,2/A_{\pm}]\times_{f_{\pm}}F_{\pm})/G_{\pm}$ with respect to the the two $\mathsf{RCD}(0,n_{\pm}+1+D)$-structures $(d_{\pm},\mathfrak{m}_{\pm})$. Moreover, over the open collars $(G((1/A_{\pm},2/A_{\pm}]\times F_{\pm}))/G_{\pm}$ the metric structures $(d_{\pm},\mathfrak{m}_{\pm})$ are the product structures $(d_{\mathbb{R}}\oplus d_{G}\oplus d_{F_{\pm}},f^{n_{\pm}}_{\pm}d\mathcal{L}\otimes d\mathfrak{m}_{G}\otimes d\mathfrak{m}_{F_{\pm}})$. From this we observe that induced metric structures on $B$ are of the form $((d_{G}\otimes d_{F_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes\mathfrak{m}_{F_{\pm}})_{\ast})$ and are given by a Riemannian metric $g_{\pm}$ on $B$: since $G\times F_{\pm}$ with $g_{G}\oplus g_{F_{\pm}}$ is a Riemannian manifold inducing $(d_{G}\oplus d_{F_{\pm}},\mathfrak{m}_{G}\oplus\mathfrak{m}_{F_{\pm}})$, and the action of $G_{\pm}$ over $G\times F_{\pm}$ is free, we obtain quotient Riemannian metrics $g_{\pm}$ on $B$ inducing $((d_{G}\oplus d_{F_{\pm}})^{\ast},(\mathfrak{m}_{G}\otimes\mathfrak{m}_{F})_{\ast})$.
We consider $[-1,1]\times B$, and consider $g(t)$ a path in the set of Riemannian metrics on $B$ with $g(t)=g_{-}$ over $[-1,-1+\varepsilon]$ and $g(t)=g_{+}$ over $[1-\varepsilon,1]$. We set $C_{0}:=[-1,1]\times B$ and equip it with the metric $d_{0}$ and measure $\mathfrak{m}_{0}$ induced by the Riemannian metric $h=dt^{2}\oplus g(t)$. We observe that since $([-1,1]\times B,h)$ is a compact Riemannian manifold, then there exists $K_{0}\in\mathbb{R}$, $|K_{0}|\leq\infty$ such that $\operatorname{Ric}(h)\geq K_{0}$. This implies that $(C_{0},d_{0},\mathfrak{m}_{0})$ is an $\mathsf{RCD}(K_{0},\dim(B)+1)$-space.
We identify $[0,2/A_{-}]\times_{f_{-}}F_{-}$ with $C_{-}:=(G\times([-1-2/A_{-},-1]\times F/(-1-2/A_{-},p)\sim(-1-2/A_{-},q)))/G_{-}$ and push the $\mathsf{RCD}(0,n_{-}+1+D)$-structure $(d_{-},\mathfrak{m}_{-})$ of $(G\times[0,2/A_{-}]\times_{f_{-}}F_{-})/G_{-}$ to get an $\mathsf{RCD}(0,n_{-}+1+D)$-structure on $C_{-}$, which we also denote by $(d_{-},\mathfrak{m}_{-})$ abusing notation. We also identify $G\times([0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ with $C_{+}:=(G\times([1,1+2/A_{+}]\times F/(1+2/A_{+},p)\sim(1+2/A_{+},q)))/G_{+}$ and push the $\mathsf{RCD}(0,n_{+}1+D)$-structure $(d_{+},\mathfrak{m}_{+})$ of $(G\times[0,2/A_{+}]\times_{f_{+}}F_{+})/G_{+}$ to get an $\mathsf{RCD}(0,n_{+}1+D)$-structure on $C_{+}$, which we also denote by $(d_{+},\mathfrak{m}_{+})$.
We consider the glued space
$X_{1}:=C_{-}\bigcup_{\mathrm{Id}_{B}}C_{0}.$
We observe that it admits an action by $G$, and moreover by construction we have $X_{1}^{\ast}=X/G\cong[-1-2/A_{-},1]$. On $X_{1}$ we define a metric $d_{1}$ given by
$d_{1}(x,y)=\begin{cases}d_{-}(x,y),&x,y\in C_{-}\setminus\partial C_{-}=C_{-}\setminus B,\\ d_{0}(x,y),&x,y\in C_{0}\setminus(\{-1\}\times B),\\ \min\{d_{-}(x,a)+d_{0}((-1,a),y)\mid a\in B\},&x\in C_{-}\mbox{ and }y\in C_{0}.\\ \end{cases}$
We also set the measure $\mathfrak{n}_{1}:=(i_{-})_{\ast}(\mathfrak{m}_{-})+(i_{0})_{\ast}(\mathfrak{m}_{0})$, where $i_{-}\colon C_{-}\to X$ and $i_{0}\colon C_{0}\to X$ are the respective inclusions. Also observe that for $K_{1}=\min\{K_{0},0\}$ and $N_{1}=\max\{n_{-}+1+D,\dim(B)+1\}$, we have that $C_{0}$, $C_{-}$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces. We consider $\delta_{1}<\min\{\varepsilon,1/A_{-}\}$ as well as the open subsets $U_{-}=C_{-}\cup([-1,-1+\delta)\times B\subset X_{1}$, and $U_{0}=(-1-\delta,1]\times B$, and observe that they are an open cover of $X_{1}$ due to the fact that the metrics $d_{-}$ and $d_{0}$ are the product metrics over the preimage of $[-1-(2/A_{-})-\delta,-1+\delta]\subset X_{1}^{\ast}$ under the quotient map $\pi\colon X_{1}\to X_{1}^{\ast}$. Moreover, since the metric $d_{1}$ and measure $\mathfrak{m}_{1}$ are the product over $[-1-\delta,-1+\delta]$, we have that the spaces $(\bar{U}_{-},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{-}})$, $(\bar{U}_{0},d_{1},\mathfrak{m}_{1}\llcorner\bar{U_{0}})$ are $\mathsf{RCD}(K_{1},N_{1})$-spaces by construction. Thus by Theorem 2.11, we conclude that $X_{1}$ is an $\mathsf{RCD}(K_{1},N_{1})$-space. Observe that since the action of $G$ is transitive on $B$, we have that
(5) $\min\{d_{-}(x,a)$ $+d_{0}((-1,a),y)\mid a\in B\}$
(6) $=\min\{d_{-}(x,g^{-1}a)+d_{0}((-1,g^{-1}a),y)\mid g^{-1}a\in B\}$
(7) $=\min\{d_{-}(x,\tilde{a}])+d_{0}((-1,\tilde{a}),y)\mid\tilde{w}\in B\}.$
Thus $G$-acts by isometries on $X_{1}$. Consider $Z\subset X_{1}$ measurable. Take $Z_{-}=\mathrm{cl}(\mathrm{cl}(Z)\cap C_{-})$ and set $Z_{0}=\mathrm{cl}(Z)\setminus Z_{-}$. Then $Z_{-}\cup Z_{0}=\operatorname{Cl}(Z)$, and they are disjoint. Since $G$ preserves the measures $(i_{-})_{\ast}(\mathfrak{m}_{-})$ and $(i_{0})_{\ast}(\mathfrak{m}_{0})$ we get that $G$ acts by measure preserving isometries.
In analogous fashion we can prove that the spaces
$X=X_{1}\bigcup_{\mathrm{Id}_{B}}C_{+}$
is an $\mathsf{RCD}(K,N)$-space where $K=K_{1}\leq 0$ and $N=\max\{N_{1},n_{+}+1+d\}$, and $G$ acts by measure preserving isometries on $X$. Moreover, we have that $X/G$ is isometric to $([-1-2/A_{-},1+2/A_{+}],dt^{2},f\mathcal{L}^{1}$), where
$f(t)=\begin{cases}(f_{A_{-}}(t))^{n_{-}},&t\in[-1-2/A_{-},-1],\\ 1&t\in[-1,1],\\ (f_{A_{+}}(t))^{n_{+}},&t\in[1,1+2/A_{+}].\end{cases}$
Thus $X$ is an $\mathsf{RCD}(K,N)$-space of cohomogeneity one with group diagram $(G,H,G_{-},G_{+})$. ∎
6. Classification of non-collapsed cohomogeneity one $\mathsf{RCD}$-spaces with $\mathrm{dim}_{\mathrm{ess}}\leq 4$
In this section we obtain the topological classification of non-collapsed $\mathsf{RCD}$-spaces admitting effective actions by a compact Lie group $G$ via measure-preserving isometries of with cohomogeneity one, and which have essential dimension less than or equal to $4$. The classification for cohomogeneity one Alexandrov spaces up to dimension $4$ was obtained by Galaz-García and Searle [28], which in turn extends the extensive work done in the smooth manifold case by Mostert [61], [60], Neumann [64], Parker [67], and Hoelscher, [46], and in the topological category by Galaz-García and Zarei [29] (see also the thesis of Muzzy [62] for a very nice account of the classification of low dimensional cohomogeneity one manifolds in the simply-connected case).
In this section we obtain that non-collapsed $\mathsf{RCD}$-spaces admitting effective actions by a compact Lie group $G$ via measure-preserving isometries of with cohomogeneity one with $\mathrm{dim}_{\mathrm{ess}}\leq 4$ are equivariantly homeomorphic to Alexandrov spaces. Thus the topological classification follows from the one in [28].
6.1. Essential dimensions $1$ and $2$.
Without any symmetry assumptions, the work of Kitabeppu-Lakzian [53] and of Lytchak-Stadler [57] shows that non-collapsed $\mathsf{RCD}(K,1)$- and $\mathsf{RCD}(K,2)$-spaces are homeomorphic to Alexandrov spaces.
6.2. Essential dimensions $3$ and $4$.
We now analyze the case of non-collapsed spaces with essential dimensions $3$ and $4$.
Let $x_{0}\in X$ be contained in a principal orbit, and take $k=\mathrm{dim}(G(x_{0}))$. Recall that the orbit $G(x_{0})$ is homeomorphic to the homogeneous space $G/G_{x_{0}}$. By [14, Chapter III, Section IV 54], there exists a $G$-invariant Riemannian metric $g_{0}$ on $G/G_{x_{0}}$ such that $G\subset\mathrm{Isom}(G/G_{x_{0}},g_{0})$. Hence, due to [63] and [55, Chapter II, Section 3, Theorem 3.1], we have the following dimension estimate
(8) $k\leq\mathrm{dim}(G)\leq\mathrm{dim}(\mathrm{Isom}(G/G_{x_{0}},g_{0}))\leq\frac{k(k+1)}{2}.$
We use this dimension estimate to limit the possible groups that can act by cohomogeneity one on a non-collapsed $\mathsf{RCD}$-space of low essential-dimension.
Theorem J.
Let $G$ be a compact Lie group acting effectively by measure preserving isometries and cohomogeneity one on a closed non-collapsed $\mathsf{RCD}(K,3)$-space $(X,d,\mathfrak{m})$. Then $G$ is one of $\mathrm{SO}(3)$ or $T^{2}$ and $X$ is equivariantly homeomorphic to an Alexandrov $3$-space.
Proof.
We begin by pointing out that, if $X^{*}=\mathbf{S}^{1}$, then $X$ is homeomorphic to a $3$-manifold by item (c) of Theorem A, and therefore the statement holds true. Thus, in the following we assume that $X^{*}=[-1,1]$.
Consider $x_{0}\in X$ fixed, contained in a principal orbit. As the action is of cohomogeneity one, we have that $\mathrm{dim}(G(x_{0}))=2$. Then the dimension estimate (8), implies that $2\leq\mathrm{dim}(G)\leq 3$. It now follows from the classification of low dimensional compact Lie groups that $G$ is either $T^{2}$, $T^{3}$ or $\mathrm{SO}(3)$.
Assume first that $G=T^{3}$, and observe that all principal isotropy groups are the same for this action, since they are conjugated to each other (see [27, Theorem 4.7 and proof of Corollary 4.9]) and $G$ is abelian. Moreover, as $G$ acts by cohomogeneity one on $X$ we have a group diagram of the form:
[…]
Thus we have that $G_{x_{0}}\subset G_{+}$ and $G_{x_{0}}\subset G_{-}$. This implies that the ineffective kernel of the action $\cap_{y\in X}G_{y}$ contains $G_{x_{0}}$. But since the action is effective, we conclude that $G_{x_{0}}=\{e\}$. This implies that $G(x_{0})$ is homeomorphic to $T^{3}$ which is a contradiction. Thus $G$ is either $T^{2}$ or $\mathrm{SO}(3)$, proving the first statement.
To proceed we observe on one hand that the possible isotropy groups when $G=T^{2}$ are $\{e\}$, $\mathbb{Z}_{k}$, $\mathbf{S}^{1}$, $\mathbb{Z}_{k}\times\mathbf{S}^{1}$, and $T^{2}$. When $G=\mathrm{SO}(3)$, since the orbits have dimension $2$, we conclude that $\mathrm{SO}(2)$ is a subgroup of any isotropy group, and thus from any isotropy group. Thus the possible isotropy groups are $\mathrm{SO}(2)$ $\mathrm{O}(2)$, $\mathrm{SO}(3)$.
We recall from [27, Corollary 10] that $\pi^{-1}(-1,1)$ is homeomorphic to $(-1,1)\times G_{x}$, where $x\in\pi^{-1}(-1,1)$.
Now from Theorem 2.22 we have that $X$ is non-collapsed, and we consider $x_{\pm}\in X$ such that $x_{\pm}^{\ast}=\pm 1$. We study the tubular neighborhoods of the orbit $G(x_{\pm})$. We denote by $m_{\pm}=\mathrm{dim}(G(x_{\pm}))$, and observe that by our hypothesis $N=3$. Due to Theorem 2.29, we have that a tubular neighborhood of the orbit $G(x_{\pm})$ is homeomorphic to $G\times_{G_{\pm}}S_{\pm}$, where $S_{\pm}$ is the slice through $x_{\pm}$.
By Theorem D and Lemma 4.8, the slice $S_{\pm}$ is homeomorphic to $\mathrm{Cone}_{0}^{N}(F)$ where $(F,d_{F},\mathfrak{m}_{F})\in\mathsf{RCD}(3-m_{\pm}-2,3-m_{\pm}-1)=\mathsf{RCD}(1-m_{\pm},2-m_{\pm})$, and $(F,d_{F},\mathfrak{m}_{F})$ is a smooth homogeneous space $K_{\pm}/H_{\pm}$. We analyze different cases based on the dimension $m_{\pm}$.
When $m_{\pm}=2$, then $F$ is 00-dimensional, and $Z=\mathrm{Cone}_{0}^{N}(F)$ is an $\mathsf{RCD}(0,1)$-space. By [53], it follows that $Z$ is homeomorphic to one of $\mathbb{R}$, $[0,\infty)$, $[0,1]$, or $\mathbf{S}^{1}$. More over by Lemma 4.8, we have that the tangent cone at $x_{\pm}$ is isometric to $Y\times Z=\mathbb{R}^{2}\times Z$. Since $X$ is non-collapsed we have that the tangent cone at $x_{\pm}$ is isometric to $\mathbb{R}^{3}$. Thus we conclude that $Z$ is isometric to $\mathbb{R}$, and thus the slice $S_{\pm}$ is homeomorphic to $\mathbb{R}$. Moreover, in this case the dimension of $G_{\pm}$ is 00, and thus $G_{\pm}$ is a discrete group acting effectively on $\mathbb{R}$ by homeomorphisms, with $S_{\pm}/G_{\pm}=[0,\varepsilon)$. Thus $G_{\pm}=\mathbb{Z}_{2}$. This implies that $G=T^{2}$, and thus we have that a tubular neighborhood of $G(x_{\pm})$ is homeomorphic to $T^{2}\times_{\mathbb{Z}_{2}}\mathbb{R}$. From the classification in [28] we see that this space is homeomorphic to an Alexandrov space.
When $m_{\pm}=1$, then $F$ is an $\mathsf{RCD}(0,1)$-space, homogeneous, and of dimension $1$. From the characterization of low dimensional $\mathsf{RCD}$-spaces [53] we have that $(F,d_{F},\mathfrak{m}_{F})$ is one of $(\mathbb{R},d_{\mathrm{E}},\mathcal{L}^{1})$, $([0,\infty),d_{\mathrm{E}},h\mathcal{L}^{1})$, $([0,1],d_{\mathrm{E}},h\mathcal{L}^{1})$, $(\mathbf{S}^{1},d_{\mathbf{S}^{1}},h\mathcal{L}^{1})$, where $h$ is a $(0,1)$-convex function. We recall that the due to our non-collapse assumption, the tangent cone at $x_{\pm}$ is isometric to $\mathbb{R}^{3}$. Moreover, due to Lemma 4.4 and Lemma 4.8 we conclude that the slice is homeomorphic to $\mathbb{R}^{2}$, and a cone over $F$. Thus the cases cases when $F$ is homeomorphic to $[0,\infty)$ and $[0,1]$ are excluded. Moreover, the isotropy $G_{\pm}$ acts smoothly by cohomogeneity $1$ (see [29, Corollary E]). Thus we have that $G_{\pm}$ is either $\mathbf{S}^{1}$ when $G=T^{2}$, or $G_{\pm}=\mathrm{SO}(2)$ or $\mathrm{O}(2)$ when $G=\mathrm{SO}(3)$. Comparing to [28], we conclude that the tubular neighborhoods are Alexandrov spaces.
In the case that $m_{\pm}=0$, we have that $F$ is an $\mathsf{RCD}(1,2)$-space. Moreover, since it is a homogeneous metric space, the measure $m_{F}$ is up to a constant the Haar measure. This implies that $F$ admits a new $\mathsf{RCD}(1,2)$-structure with respect to the $2$-dimensional Hausdorff measure. By [57] it follows that $F$ is homeomorphic to an Alexandrov surface. Thus we have that $S_{\pm}$ is homeomorphic to the tubular neighborhood of $G(x_{\pm})=\{x_{\pm}\}$, and to the cone over $F$, which admits an Alexandrov space structure.
Thus by [28], our conclusion follows, since $X$ is equivariantly indistinguishable from a cohomogeneity one Alexandrov space. ∎
As pointed out in [46, Section 1.6], every compact connected Lie group $G$ has a finite cover of the form $G_{ss}\times T^{k}$, where $G_{ss}$ is semisimple and simply connected. If $G$ acts on $(X,d,m)$ by measure-preserving isometries then every finite cover $\tilde{G}$ of $G$ also acts on $(X,d,m)$ by measure-preserving isometries in the obvious way, but the action of $\tilde{G}$ has non-trivial (but finite) ineffective kernel. Therefore, if we allow for a finite ineffective kernel we can assume $G$ to be of the form $G_{ss}\times T^{k}$.
The classification of semisimple and simply connected compact Lie groups in low dimensions is well known. For the readers convenience we reproduce the part of [46, Table 1.29] that we need below, namely, the list of these groups up to dimension $3$ (there are no semisimple and simply connected compact groups on dimensions $>3$ up to dimension $8$).
Group Dimension
$\mathbf{S}^{1}\cong\mathrm{U}(1)\cong\mathrm{SO}(2)$ $1$
$\mathbf{S}^{3}\cong\mathrm{SU}(2)\cong\mathrm{Sp}(1)\cong\mathrm{spin}(3)$ $3$
Let us further observe that although our main topological recognition tool, Theorem D is formulated for effective actions, one obtains the same result in the case of almost effective actions. This follows immediately by observing that from an almost effective action of a Lie group $G$ one gets an effective action of $G/K$ where $K$ is the ineffective kernel of the $G$-action by letting $(gK)x:=gx$ for each $x\in X$. The $G/K$ action then has the same orbits as the $G$-action; consequently the slices of both actions coincide and, since $K$ is finite, the dimensions of $G$ and $G/K$ agree.
In the following theorem we assume as is usual that the action is non-reducible (cf. for example with [45, Definition 2.1.19]) and that the group is acting almost effectively so that we can use the characterization of groups up to cover in the table above. It is worth pointing out that in the list of groups appearing in the corresponding theorem for Alexandrov spaces [28, Theorem C], the identification $\mathrm{SO}(4)\cong\mathbf{S}^{3}\times\mathbf{S}^{3}$ is used. As we are primarily interested in the topological recognition of non-collapsed $\mathsf{RCD}(K,4)$-spaces $X$ we only assume that the action is almost effective, to deal with less possible groups acting on $X$. Note however that, once the classification has been established in this case, one can recover the full classification.
Theorem K.
Let $G$ be a compact Lie group acting almost effectively by measure preserving isometries and cohomogeneity one on a closed non-collapsed $\mathsf{RCD}(K,4)$-space $(X,d,\mathfrak{m})$. Then $G$ is, up to finite cover, one of $\mathbf{S}^{3}$, $\mathbf{S}^{3}\times\mathbf{S}^{3}$, $\mathbf{S}^{1}\times\mathbf{S}^{3}$, $T^{3}$, and $X$ is homeomorphic to an Alexandrov space.
Proof.
As in the three-dimensional case, the topological and equivariant type of the space is that of a cohomogeneity one manifold if $X^{*}=\mathbf{S}^{1}$, and the theorem holds by [67]. Therefore we only need to analyze the case in which $X^{*}=[-1,1]$.
We fix a point $x_{0}$ on a principal orbit and note that $\mathrm{dim}(G(x_{0}))=3$. Now the bound of the dimension estimates (8) grant that $3\leq\mathrm{dim}(G)\leq 6$. As we are allowing for a finite ineffective kernel, the classification of low dimensional compact Lie groups grants that $G$ is, up to finite cover, one of $\mathbf{S}^{3}$, $\mathbf{S}^{3}\times\mathbf{S}^{3}$, $\mathbf{S}^{1}\times\mathbf{S}^{3}$, $T^{2}\times\mathbf{S}^{3}$, $T^{3}\times\mathbf{S}^{3}$, $T^{3}$, $T^{4}$, $T^{5}$ and $T^{6}$. We can immediately rule out the tori $T^{k}$ for $k\geq 4$, as in these cases the principal orbits are homeomorphic to $T^{k}/H$ where $H$ is finite, which is impossible as the space $X$ is $4$-dimensional. Moreover, as we have assumed the action to be non-reducible we can also rule out $T^{2}\times\mathbf{S}^{3}$ by noting that in such a case, we have a $2$-dimensional subgroup $H$ of $\mathbf{S}^{3}\times\{1\}$, which is a contradiction since $\mathbf{S}^{3}$ has no $2$-dimensional subgroups. Finally $T^{3}\times\mathbf{S}^{3}$ is ruled out because the principal isotropy being $3$-dimensional, contains a torus as a proper normal subgroup acting trivially on any point of the space. This contradicts the almost effectiveness of the action. Whence, the first claim is settled.
We consider now $x_{\pm}\in X$ with $x_{\pm}^{\ast}=\pm 1$ and denote by $m_{\pm}=\mathrm{dim}(G(x_{\pm}))$ as in the previous theorem. Observe now that by our hypothesis $N=4$, and that by the Slice Theorem 2.29 we have that small tubular neighborhoods of the orbits $G(x_{\pm})$ are homeomorphic to $G\times_{G_{\pm}}S_{\pm}$, where $S_{\pm}$ is the slice through $x_{\pm}$. Therefore, Theorem D implies that the slice $S_{\pm}$ is homeomorphic to $\mathrm{Cone}_{0}^{N}(F)$ where $(F,d_{F},\mathfrak{m}_{F})\in\mathsf{RCD}(4-m_{\pm}-2,4-m_{\pm}-1)=\mathsf{RCD}(2-m_{\pm},3-m_{\pm})$, and $(F,d_{F},\mathfrak{m}_{F})$ is a smooth homogeneous space $K_{\pm}/H_{\pm}$. We now analyze different cases depending on the value of $m_{\pm}$.
In the case that $m_{\pm}=3$, then $F$ is 00-dimensional. Hence, $Z=\mathrm{Cone}_{0}^{N}(F)$ is an $\mathsf{RCD}(0,1)$-space and, as in the previous theorem, due to [53], is one of $\mathbb{R}$, $[0,\infty)$, $[0,1]$ or $\mathbf{S}^{1}$. Similarly to the proof of the previous theorem, using Lemma 4.8, we get that the tangent cone at $x_{\pm}$ is homeomorphic to $Y\times Z=\mathbb{R}^{3}\times Z$. By the non-collapsed assumption on $X$ we have that the tangent cone at $x_{\pm}$ is homeomorphic to $\mathbb{R}^{4}$. Whence, the slice $S_{\pm}$ is homeomorphic to $\mathbb{R}$. Now again, in this case the dimension of $G_{\pm}$ is zero and therefore $G_{\pm}$ is a discrete group acting effectively on $\mathbb{R}$ by homeomorphisms, with $S_{\pm}/G_{\pm}=[0,\varepsilon)$. This is only possible if $G_{\pm}=\mathbb{Z}_{2}$, in turn forcing $G$ to be either $\mathbf{S}^{3}$ or $T^{3}$. Thus, the small tubular neighborhoods of $G(x_{\pm})$ are equivariantly homeomorphic to either $\mathbf{S}^{3}\times_{\mathbb{Z}_{2}}\mathbb{R}$ or $T^{3}\times_{\mathbb{Z}_{2}}\mathbb{R}$, and it follows from the classification [28] that $X$ is then homeomorphic to an Alexandrov space.
Now we turn to the case in which $m_{\pm}\leq 2$. Under this assumption $2-m_{\pm}\geq 0$ and by [77, Theorem 1], $F$ has a Riemannian metric of positive Ricci curvature if $m_{\pm}\leq 1$ and of non-negative Ricci curvature if $m_{\pm}=2$. Hence, by using [41, Main Theorem], $F$ admits a metric of positive sectional curvature if $m_{\pm}\leq 1$ and of nonnegative sectional curvature if $m_{\pm}=2$. To conclude we only need to observe then that the possible choices of $F$ already appear in the classification [28], and therefore $X$ is homeomorphic to an Alexandrov space. ∎
7. New Examples of $\mathsf{RCD}$-spaces of cohomogeneity one
In this section we use the characterization of homogeneous spaces with positive Ricci curvature in [6] to present examples of non-collapsed $\mathsf{RCD}$-spaces with an action of cohomogeneity one by measure preserving isometries. We also give an explicit example of a non-collapsed $\mathsf{RCD}$ spaces of cohomogeneity one with essential dimension equal to $5$ that do not admit a metric making them an Alexandrov space, showing that the results in Section 6 are optimal.
Theorem 7.1.
Let $F=G/H$ be a homogeneous space of dimension $n\geq 2$ with finite fundamental group. Then there exists a $G$-invariant Riemannian metric $g$ with $\operatorname{Ric}(g)\geq(n-1)>0$ and $\operatorname{diam}(g)\leq\pi$. Moreover, for any $K\geq 0$, the $(K,n)$-cone $\mathrm{Con}^{n}_{K}(F)$ is a non-collapsed $\mathsf{RCD}(Kn,n+1)$ with an action by measure preserving isometries of $G$ making it a cohomogeneity one $\mathsf{RCD}$-space. In the case of $K>0$, the space is homeomorphic to the suspension of $F$ and the orbit space is isometric to $[0,\pi/\sqrt{K}]$ with weighted measure $\sin_{K}(t)^{n}d\mathcal{L}^{1}$ and we have the group diagram $(G,H,G,G)$. In the case when $K=0$, then the space is homeomorphic to the cone over $F$ with the vertex the only fixed point, and the orbit space is isometric to $[0,\infty)$ with weighted measure $t^{n}d\mathcal{L}^{1}$.
Proof.
By [6, Theorem 1] $F$ admits a $G$-invariant Riemannian metric $\tilde{g}$ with $\operatorname{Ric}(\tilde{g})\geq a>0$. Consider $\lambda>0$ such that $a/(n-1)\geq\lambda^{2}$ and $\pi/\operatorname{diam}(\tilde{g})\geq\lambda$. Then with respect to the Riemannian metric $g=\lambda g$, we have that $(F,d_{g},\mathrm{vol}(g))=(F,\lambda d_{\tilde{g}},\lambda^{n}\mathrm{vol}(\tilde{g}))$ is an $\mathsf{RCD}(a\lambda^{-2},n)$-space. Since by construction $a\lambda^{-2}\geq(n-1)$, then $(F,d_{g},\mathrm{vol}(g))$ is an $\mathsf{RCD}(n-1,n)$-space. By [73, Theorem 2 (i)] it follows that $\operatorname{Ric}(g)\geq(n-1)$. Moreover, by construction we have that $\operatorname{diam}(g)=\lambda\operatorname{diam}(\tilde{g})\leq\pi$. Then by [51, Theorem 1.1] we obtain that for $K\geq 0$ the $(K,n)$-cone $\mathrm{Con}^{n}_{K}(F)$ is $\mathsf{RCD}(Kn,n+1)$. But by construction for $K>0$ the underlying topological space of $\mathrm{Con}^{n}_{K}(F)$ is $F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})$, and thus homeomrphic to the suspension $\mathrm{Susp}(F):=F\times[0,1]/(F\times\{0,\pi/\sqrt{K}\})$. And for $K=0$ the underling space of $\mathrm{Con}_{K}^{n}(F)$ is the topological cone $F\times[0,\infty)/(F\times\{0\})$.
Recall that in both cases we can define an action of $G$ on $\mathrm{Con}_{K}^{n}(F)$ by setting
$g[x,t]:=[gx,t].$
In the case when $K=0$ we refer to this action by the cone of the action of $G$ on $F$, and for $K>0$ we refer to this action by suspension of the action of $G$ on $F$.
Recall from Section 2.4 that for $K\geq 0$ the distance function $d_{\mathrm{Con_{K}^{n}}}$ of $\mathrm{Con}^{n}_{K}(F)$ is given by
$d_{\mathrm{Con_{K}^{n}}}$ $([x,s],[y,t])$
:=$:=\begin{cases}\cos_{K}^{-1}\Big{(}\cos_{K}(s)\cos_{K}(t)+K\sin_{K}(s)\sin_{K}(t)\cos\big{(}d_{F}(x,y)\wedge\pi\big{)}\Big{)}&\mbox{if }K>0,\\ \sqrt{s^{2}+t^{2}-st\cos\big{(}d_{F}(x,y)\wedge\pi\big{)}}&\mbox{if }=0.\end{cases}$
where $d_{F}(x,y)\wedge\pi:=\min\{d_{F}(x,y),\pi\}$, $\cos_{K}\colon[0,\pi/\sqrt{K}]\to[0,1]$ given by $\cos_{K}(t):=\cos(\sqrt{K}t)$, and $\sin_{K}\colon[0,\pi/\sqrt{K}]\to[0,\infty)$ given by
$\sin_{K}(t):=\begin{cases}\frac{1}{\sqrt{K}}\sin(\sqrt{K}t)&\mbox{if }K>0,\\ t&\mbox{if }K=0.\end{cases}$
From the description of $d_{\mathrm{Con}_{K}^{n}}$ it easy to see that $G$ acts by isometries on $\mathrm{Con}_{K}^{n}(F)$, with respect to the cone and suspension of the action of $G$ on $F$ depending on $K$.
The measure $\mathfrak{m}_{\mathrm{Con}_{K}^{n}}$ of $\mathrm{Con}_{K}^{n}(F)$ is defined as $\mathfrak{m}_{\mathrm{Con}_{K}^{n}}:=(\sin_{K}(t))^{n}dt\otimes\mathfrak{m}_{F}$. Thus for $g\in G$ fixed, denoting by $\alpha_{g}\colon F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})\to F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})$ the homeomorphism given by $\alpha_{g}[x,t]:=[gx,t]$, we have for $A=A_{1}\times A_{2}\subset F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})$ that
$(\alpha_{g})_{\#}(\mathfrak{m}_{\mathrm{Con}_{K}^{n}})(A)$ $=\mathfrak{m}_{\mathrm{Con}_{K}^{n}}(\alpha_{g}^{-1}(A))$
$=\mathfrak{m}_{\mathrm{Con}_{K}^{n}}(A_{1}\times\alpha_{g}^{-1}(A_{2}))$
$=\int_{A_{1}\times\alpha_{g}^{-1}(A_{2})}\sin_{K}(t))^{N}\,dt\otimes d\mathfrak{m}_{F}$
$=\int_{A_{1}}\sin_{K}(t))^{N}\int_{\alpha_{g}^{-1}(A_{2})}d\,\mathfrak{m}_{F}\,dt$
$=\int_{A_{1}}\sin_{K}(t))^{N}\int_{A_{2})}d\,(\alpha_{g})_{\#}\mathfrak{m}_{F}\,dt$
$=\int_{A_{1}}\sin_{K}(t))^{N}\int_{A_{2})}d\,\mathfrak{m}_{F}\,dt$
$=\int_{A_{1}\times A_{2}}\sin_{K}(t))^{N}\,dt\otimes d\mathfrak{m}_{F}$
$=\mathfrak{m}_{\mathrm{Con}_{K}^{n}}(A).$
Thus $G$ acts by measure preserving isometries on $\mathrm{Con}_{K}^{n}(F)$. Moreover, the vertexes of the suspension and the vertex of the cone are fixed points, and the group $H$ is the isotropy group of any point $[x,t]$ with $0<t<\pi/\sqrt{K}$ when $K>0$, or with $t>0$ when $K=0$.
Last we point out that the essential dimension of $\mathrm{Con}^{n}_{K}(F)$ is $n+1$. Thus by [23, Theorem 1.12] and [11, Theorem 1.3] we have that $\mathrm{Con}^{n}_{K}(F)$ is non-collapsed. ∎
Theorem F.
The suspension $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ of $\mathbf{S}^{2}\times\mathbf{S}^{2}$ admits a non-collapsed $\mathsf{RCD}(K,5)$-structure for any $K\geq 0$ such that the suspension of the $\mathrm{SO}(3)\times\mathrm{SO}(3)$-action on $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is by measure preserving isometries and by cohomogeneity one, with group diagram $(\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(2)\times\mathrm{SO}(2),\mathrm{SO}(3)\times\mathrm{SO}(3),\mathrm{SO}(3)\times\mathrm{SO}(3))$. Moreover, $\mathrm{Susp}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ with this action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ cannot be an Alexandrov space of cohomogeneity one. I.e., it does not admit a metric making it an Alexandrov space and such that the action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ is by isometries.
Proof.
We observe that $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is the homogeneous space $\mathrm{SO}(3)\times\mathrm{SO}(3)/(\mathrm{SO}(2)\times\mathrm{SO}(2))\cong(\mathrm{SO}(3)/\mathrm{SO}(2))\times(\mathrm{SO}(3)/\mathrm{SO}(2))$. Thus by Theorem 7.1, there exists a $\mathrm{SO}(3)\times\mathrm{SO}(3)$-invariant Riemannian metric $g$ on $\mathbf{S}^{2}\times\mathbf{S}^{2}$ such that for any $K\geq 0$ the $(K,5)$-cone $\mathrm{Con}^{5}_{K}(\mathbf{S}^{2}\times\mathbf{S}^{2})$ is a non-collapsed $\mathsf{RCD}(K4,5)$-space with a cohomogeneity one action by measure preserving isometries of $\mathrm{SO}(3)\times\mathrm{SO}(3)$. But for $K>0$ the $(K,5)$-cone is homeomorphic to the suspension of $\mathbf{S}^{2}\times\mathbf{S}^{2}$. This proves the first claim.
Assume that the suspension of $\mathbf{S}^{2}\times\mathbf{S}^{2}$ admits an Alexandrov metric with an action of cohomogeneity one by isometries of $\mathrm{SO}(3)\times\mathrm{SO}(3)$ fixing the vertices of the suspension and with principal isotropy $\mathrm{SO}(2)\times\mathrm{SO}(2)$. Consider $X$ the space of directions at one of the vertexes of the suspension. This space is homeomorphic to $\mathbf{S}^{2}\times\mathbf{S}^{2}$, $\mathrm{SO}(3)\times\mathrm{SO}(3)$ acts transitively on $X$, and it is a $4$-dimensional Alexandrov space of curvature at least one. Thus $X$ is a homogeneous space, and by [5, Theorem 7 I] $X$ is isometric to a homogeneous Finsler manifold. Since $X$ is an Alexandrov space of curvature at least $1$, then it is also an $\mathsf{RCD}(3,4)$-space, then $X$ is isometric to a homogenous Riemannian manifold $(M,g)$. Moreover we have that $\sec(g)\geq 1$. Since we have non-trivial Killing vector fields by the action of $\mathrm{SO}(3)\times\mathrm{SO}(3)$, by [50, Theorem 1] $M$ is homeomorphic to $\mathbf{S}^{2}$ or $\mathbb{C}P^{2}$, which is a contradiction. Thus $\mathbf{S}^{2}\times\mathbf{S}^{2}$ is not an $\mathrm{SO}(3)\times\mathrm{SO}(3)$-cohomogeneity one Alexandrov space. ∎
Remark 7.2.
When considering $G/H$ a homogeneous smooth manifold with finite fundamental group, using Theorem B for the group diagram $(G,H,G,G)$ we also get an $\mathsf{RCD}(K,N)$-space. From the proof if follows that this case $K=0$ and $N=\dim(G/H)+\dim(G)+1$, and the orbit space is isometric to $[-1-2/A,1+2/A]$ for an appropriate choice of $A$ and measure $f^{\dim(G/H)}\mathcal{L}$, where $f$ is a non-smooth function. Nonetheless, this space is equivariantly homeomorphic to the suspension of $G/H$ and thus to the space given by Theorem 7.1. That is, for a group diagram $(G,H,G,G)$ with $G/H$ having finite fundamental group, by Theorem B and Theorem 7.1 we obtain two equivariantly homeomorphic spaces, but each one with a different $\mathsf{RCD}$-structure.